### v2

# Intuitionism with Truth Tables: A Decision Procedure for IPL Based on RNmatrices

Abstract
This paper provides a new decision procedure for intuitionistic propositional logic (IPL) based on a 3-valued restricted non-deterministic matrix and proves its soundness and completeness.
1 Introduction
A famous result, proved first by K. Gödel in $1932$, says that it is impossible to characterize intuitionistic propositional logic (IPL) via a finite-valued (deterministic) matrix [1]. In other words, Gödel proved that there is no effective decision procedure based on finite-valued ordinary truth-tables for IPL.
Later, in $1940$, a maybe less famous but equally important result was proved by Dugundji, who adapted Gödel’s argument to establish that the same is also true about the family of modal logics between S1 and S5 [1].
To overcome Dugundji’s result, J. Kearns [6] introduced a $4$-valued Nmatrix (non-deterministic matrix) for modal logics KT, S4, and S5 and restricted the set of valuations to a subset called “level valuations”. He proved that this restricted Nmatrix (hereafter called RNmatrix [3]) constitutes sound and complete semantics for these modal logics.
Unfortunately, in principle, Kearns’s level valuations fail to provide an effective decision procedure for any of the modal logics. More than that, until recently, it was unknown whether obtaining such a decision procedure from level valuations would be possible.
(…) the interesting problem of determining whether or not a decision procedure can be obtained by means of level valuations remains open. [2]
L. Grätz recently solved this problem for some modal systems [4]. He refined Kearn’s original RNmatrix to obtain a decidable $3$-valued RNmatrix for modal logics KT and S4. He noticed that, with an appropriate notion of partial valuation for level semantics (i.e., preserving analyticity and co-analyticity), the truth tables could be obtained by removing any spurious row eventually produced in the valuation process.
In this paper, we will try to overcome the original problem posed by Gödel on the characterizability of IPL via finite-valued matrix. We will propose a decidable 3-valued RNmatrix for IPL combining Grätz’s RNmatrix for S4 and Gödel-McKinsey-Tarski’s box translation.
2 Nmatrices and RNmatrices
Along this paper, all the propositional languages will be defined over a given denumerable set $\mathcal{P}=\{p_{0},p_{1},p_{2},\ldots\}$ of propositional variables.
Definition 2.1 (Nmatrix).
Let $\Theta$ be a signature and $For(\Theta)$ the propositional language over signature $\Theta$ generated by the set $\mathcal{P}$ of propositional variables. An Nmatrix for $\mathcal{L}$ over $\Theta$ is a tuple $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ where
  1. 1.
    $\mathcal{V}$ is the (non-empty) set of truth values;
  2. 2.
    $\mathcal{D}\subset\mathcal{V}$ is the (non-empty) set of designated values;
  3. 3.
    $\mathcal{O}$ is a function assigning, to each $n$-ary connective ##\## of $\Theta$, a (non-deterministic) truth function $\mathcal{O}(\#):\mathcal{V}^{m}\rightarrow(\wp(\mathcal{V}))\backslash\emptyset$.
A valuation over an Nmatrix $\mathcal{M}$ is a function $v:For(\Theta)\to\mathcal{V}$ such that, for every $n$-ary connective #normal-#\## of $\Theta$ and for every $\varphi_{1},\ldots,\varphi_{n}\in For(\Theta)$, $v(\#(\varphi_{1},\ldots,\varphi_{n}))\in\mathcal{O}(\#)(v(\varphi_{1}),\ldots,v(\varphi_{n}))$. The set of valuations over $\mathcal{M}$ is denoted by $Val(\mathcal{M})$.
Definition 2.2 (RNmatrix).
Let $\Theta$ be a signature. An RNmatrix [3] over $\Theta$ is a pair $\mathcal{R}=\langle\mathcal{M},\mathcal{F}\rangle$ where $\mathcal{M}$ is an Nmatrix over $\Theta$ and $\mathcal{F}\subseteq Val(\mathcal{M})$.
Definition 2.3 (Consequence relation in an (R)Nmatrix).
Let $\mathcal{M}$ ($\mathcal{R}$, resp.) be an Nmatrix (an RNmatrix, resp.) over $\Theta$, and $\Gamma\cup\{\varphi\}\subseteq For(\Theta)$. We say that $\varphi$ is a semantical consequence of $\Gamma$ w.r.t. $\mathcal{M}$ (w.r.t. $\mathcal{R}$, resp.), denoted by $\Gamma\models_{\mathcal{M}}\varphi$ (denoted by $\Gamma\models_{\mathcal{R}}\varphi$, resp.) if, for every $v\in Val(\mathcal{M})$ (for every $v\in\mathcal{F}$, resp.): $v(\varphi)\in\mathcal{D}$ whenever $v(\gamma)\in\mathcal{D}$ for every $\gamma\in\Gamma$. We write $\models_{\mathcal{M}}\varphi$ and $\models_{\mathcal{R}}\varphi$ instead of $\emptyset\models_{\mathcal{M}}\varphi$ and $\emptyset\models_{\mathcal{R}}\varphi$.
$\neg^{\mathcal{A}_{S4}}$
0 $\{1,2\}$
1 $\{0\}$
2 $\{0\}$
$\Box^{\mathcal{A}_{S4}}$
0 $\{0\}$
1 $\{0\}$
2 $\{2\}$
$\rightarrow^{\mathcal{A}_{S4}}$ 0 1 2
0 $\{1,2\}$$\{1,2\}$$\{1,2\}$
1 $\{0\}$$\{1,2\}$$\{1,2\}$
2 $\{0\}$$\{1\}$$\{1,2\}$
$\lor^{\mathcal{A}_{S4}}$ 0 1 2
0 $\{0\}$$\{1,2\}$$\{1,2\}$
1 $\{1,2\}$$\{1,2\}$$\{1,2\}$
2 $\{1,2\}$$\{1,2\}$$\{1,2\}$
$\land^{\mathcal{A}_{S4}}$ 0 1 2
0 $\{0\}$$\{0\}$$\{0\}$
1 $\{0\}$$\{1,2\}$$\{1,2\}$
2 $\{0\}$$\{1,2\}$$\{1,2\}$
Table 1: Tables for $\mathcal{M}_{S4}$.
Example 2.4 ($\mathcal{M}_{S4}$).
Let $\Sigma_{b}=\{\{\neg,\Box\},\{\rightarrow\}\}$ and Σ={{¬,},{,,}}normal-$\Sigma=\{\{\neg,\Box\},\{\rightarrow,\lor,\land\}\}$ be, respectively, the basic signature and the signature for modal logic S4. Given the set $\mathcal{P}$ of propositional variables, let $For(\Sigma_{b})$ and $For(\Sigma)$ be the corresponding propositional languages. Grätz’s Nmatrix over $\Sigma_{b}$ is the tuple $\mathcal{M}_{S4}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$, where $\mathcal{V}=\{0,1,2\}$, $\mathcal{D}=\{1,2\}$, $\mathcal{O}=\{\neg^{\mathcal{A}_{S4}},\Box^{\mathcal{A}_{S4}},\rightarrow^{\mathcal{A}_{S4}}\}$ and each $c\in\mathcal{O}$ is defined in Table 1.
$p$$\Box p$$\Box\Box p$$\Box p\rightarrow\Box\Box p$000121001222212
(a) Valuations in $\mathcal{M}_{S4}$ for $\Box p\rightarrow\Box\Box p$
Figure 1: Each tree level corresponds to the set of values resulting from the multioperation applied to the respective subformula (at left) in that tree. At the top are the values that the atomic formula can receive, and in the leaves are the values of the evaluated formula.
By setting $\varphi\lor\psi:=\neg\varphi\to\psi$ and $\varphi\land\psi:=\neg(\varphi\to\neg\psi)$, respectively, the multioperators $\lor^{\mathcal{A}_{S4}}$ and $\land^{\mathcal{A}_{S4}}$ are defined as in Table 1.
Valuations in an Nmatrix are typically represented as tables. However, if one focuses more on the computational aspects of these valuations, adopting the more flexible data structure of binary trees can be advantageous. In Figure 1, we show how to compute the valuation of two examples.
3 Kearns’s level valuations
Regarding semantics for modal logic, Kripke-style semantics are the usual choice. However, there are other attractive alternatives. In the 1980s, Kearns proposed an RNmatrix-based semantics for modal logics KT, S4 and S5. He introduced the level valuations as the set $\mathcal{F}$ of valuations to be used.
[Kearn’s account] is simpler than [Kripke’s account] in virtue of having dispensed with possible worlds and their relations. I also think that my account is philosophically preferable to the standard account for having done this. For I do not think there are such things as possible worlds, or even that they constitute a useful fiction. [6]
We will define the level valuations for the modal logic S4.
Figure 2: Assume that $Val(\mathcal{M}_{S4})=\{v_{0},v_{1},\ldots,v_{m},v_{n},v_{k},\ldots\}$. We remove a valuation $v$ in a level $\mathcal{L}_{i+1}$ iff there is some wf $\alpha$ such that $v(\alpha)=1$ and $w(\alpha)\in\mathcal{D}$ for every $w\in\mathcal{L}_{i}$.
Definition 3.1.
(Level valuations for S4)
Assume that $Val(\mathcal{M}_{S4})$ is the set of valuations over $\mathcal{M}_{S4}$. We define the level $\mathcal{L}^{S4}_{k}$, where $k\in\mathbb{N}$, as follows:
  1. i)
    $\mathcal{L}_{0}^{S4}=Val(\mathcal{M}_{S4})$;
  2. ii)
    $\mathcal{L}^{S4}_{k+1}=\{v\in\mathcal{L}^{S4}_{k}\mid\forall\alpha\in For(\Sigma_{b}),\mathcal{L}^{S4}_{k}(\alpha)\subseteq\mathcal{D}\Rightarrow v(\alpha)=2\}$
where $\mathcal{L}^{S4}_{k}(\alpha)=\{v(\alpha)\mid v\in\mathcal{L}^{S4}_{k}\}$.
Definition 3.2.
The set of level valuations in $\mathcal{M}_{S4}$ is defined as the intersection of every level valuation:
$\mathcal{L}_{S4}=\cap^{\infty}_{k\geq 0}\mathcal{L}^{S4}_{k}$
Definition 3.3 (Grätz’s RNmatrix for S4).
Let $\mathcal{M}_{S4}$ be the nmatrix defined in Definition 2.4. The RNmatrix for S4 is the pair $\mathcal{R}(\mathcal{M}_{S4})=\langle\mathcal{M}_{S4},\mathcal{L}_{S4}\rangle$.
Let $\vdash_{S4}^{b}$ be the consequence relation of the standard Hilbert calculus $\mathcal{H}^{b}_{S4}$ for S4 over $\Sigma_{b}$. As usual, $\Gamma\vdash_{S4}^{b}\varphi$ iff either $\varphi$ is a theorem of $\mathcal{H}^{b}_{S4}$ or there exists a non-empty set $\{\beta_{1},\ldots,\beta_{n}\}\subseteq\Gamma$ such that $\beta_{1}\to(\beta_{2}\to(\ldots\to(\beta_{n}\to\varphi)\ldots))$ is a theorem of $\mathcal{H}^{b}_{S4}$.
Theorem 3.4 ([4]).
For every $\Gamma$ and $\varphi$: $\Gamma\vdash_{S4}^{b}\varphi$ iff $\Gamma\models_{\mathcal{R}(\mathcal{M}_{S4})}\varphi$.
3.1 Grätz’s decision procedure for S4
While the soundness and completeness of level valuations for S4 (with the rule of necessitation) were already known, it was only recently that an effective decision procedure was discovered.
Before that, J. V. Ivlev seems to have suggested the following approach [4]: remove every partial valuation $v_{p}$ whenever (1) $v_{p}(\Box A)\not\in D$ and (2) $w_{p}(A)\in D$ for every partial valuation $w_{p}$. The reason for this method is very intuitive if we consider the definition of level valuations, and indeed, it works for $2$-valued semantics, as well as for some examples in $\mathcal{M}_{S4}$, such as $\Box\alpha$ where $\alpha$ is any propositional tautology. However, this criterion is not sufficient for $\mathcal{M}_{S4}$ because it leaves behind invalid partial level valuations (consider the valid formula $\Box p\rightarrow\Box\neg\neg p$ in $\mathcal{M}_{S4}$).
In [4], Grätz identified the need to find a suitable method by refining the concept of partial valuation. To address this, he suggests a new definition for partial level valuations:
Definition 3.5 (Partial valuation in $\mathcal{M}_{S4}$).
A partial valuation in $\mathcal{M}_{S4}$ is a function $\tilde{v}_{p}:\Lambda\rightarrow\mathcal{V}$ such that $\Lambda\subseteq For(\Sigma_{b})$ is closed under subformulas [^4] That is: if $\alpha\in\Lambda$ and $\beta$ is a subformula of $\alpha$ then $\beta\in\Lambda$. and, for every $\alpha,\beta\in\Lambda$:
[…]
  • if $\alpha\to\beta\in\Lambda$ then $\tilde{v}_{p}(\alpha\to\beta)\in{\to}^{\mathcal{A}_{S4}}(\tilde{v}_{p}(\alpha),\tilde{v}_{p}(\beta))$.
Definition 3.6 (Grätz’s partial level valuation).
A partial valuation $v_{p}:\Lambda\rightarrow\{0,1,2\}$ in $\mathcal{M}_{S4}$ is a partial level valuation in $\mathcal{R}(\mathcal{M}_{S4})$ ($v_{p}\in PLV$) iff
  1. 1.
    $\forall\alpha\in\Lambda$ such that $v_{p}(\alpha)=1$, there exists $w_{p}\in PLV$ such that $w_{p}(\alpha)=0$ and, $\forall\beta\in\Lambda$, $w_{p}(\beta)=2$ whenever $v_{p}(\beta)=2$.
Remark 3.7.
In [4, p. 15] it is observed that, being $\Lambda$ finite, the set of $PLV$s with domain $\Lambda$ is well-defined (that is, it is not cyclic), given that the search space is finite: the set of all the partial valuations with domain $\Lambda$ is finite. Moreover, it is proved in [4] that this notion is analytic in $\mathcal{M}_{S4}$ (i.e, every partial level valuation in $\mathcal{M}_{S4}$ can be extended to a level valuation in $\mathcal{M}_{S4}$) and co-analytic (i.e, for any level valuation $v$ in $\mathcal{M}_{S4}$, the restriction of $v$ to a partial level valuation $v_{p}$ defined over a finite set $\Lambda$ closed under subformulas is a partial level valuation).
Theorem 3.8.
$\mathcal{R}(\mathcal{M}_{S4})$ is analytic and co-analytic with PLV.
Using this, one can easily prove that the truth table method is sound and complete, that is,
Theorem 3.9 ([4]).
For every $\varphi$: $\models_{\mathcal{R}(\mathcal{M}_{S4})}\varphi$ if, and only if,for every partial valuation $v_{p}$ in $\mathcal{M}_{S4}$ defined over the set $\Lambda$ of subformulas of $\varphi$: if $v_{p}\in PLV$, then $v_{p}(\varphi)\in\mathcal{D}$.
Corollary 3.9.1 ([4]).
For every $\varphi$: $\vdash_{S4}^{b}\varphi$ if, and only if, for every partial valuation $v_{p}$ in $\mathcal{M}_{S4}$ defined over the set $\Lambda$ of subformulas of $\varphi$: if $v_{p}\in PLV$ then $v_{p}(\alpha)\in\mathcal{D}$.
Proof.
It is an immediate consequence of Theorems 3.9 and 3.4. ∎
Therefore, to construct a truth table, compute every partial valuation and remove any non-PLV. And this algorithmic process is a sound and complete decision procedure for S4. Obviously the soundness and completeness Theorem 3.9 (and so, the decision procedure) can be extended to inferences in S4 from a finite set of premises, taking into account that $\vdash_{S4}^{b}$ (and so $\models_{\mathcal{R}(\mathcal{M}_{S4})}$) satisfies the deduction metatheorem.
3.2 Reducing Grätz’s RNmatrix and algorithm for S4
As suggested by Grätz ([4], p. 17), it is possible to apply the notion of partial level valuation to the tables for the connectives of S4 (in our case, this includes conjunction and disjunction). Then, by eliminating rows $v_{p}$ which are not supported by any other $w_{p}$ as prescribed in Definition 3.6, it is obtained a reduced truth-table for each multioperator of S4 over the full signature $\Sigma$. Along this section we will deal exclusively with signature $\Sigma$ for S4.
[…]
$\neg^{\mathcal{A}_{S4}}$
0 $\{1,2\}$
1 $\{0\}$
2 $\{0\}$
$\Box^{\mathcal{A}_{S4}}$
0 $\{0\}$
1 $\{0\}$
2 $\{2\}$
$\rightarrow^{\mathcal{A}^{\prime}_{S4}}$ 0 1 2
0 $\{1,2\}$$\{1,2\}$$\{2\}$
1 $\{0\}$$\{1,2\}$$\{2\}$
2 $\{0\}$$\{1\}$$\{2\}$
$\lor^{\mathcal{A}^{\prime}_{S4}}$ 0 1 2
0 $\{0\}$$\{1,2\}$$\{2\}$
1 $\{1,2\}$$\{1,2\}$$\{2\}$
2 $\{2\}$$\{2\}$$\{2\}$
$\land^{\mathcal{A}^{\prime}_{S4}}$ 0 1 2
0 $\{0\}$$\{0\}$$\{0\}$
1 $\{0\}$$\{1\}$$\{1\}$
2 $\{0\}$$\{1\}$$\{2\}$
Table 2: Tables for $\mathcal{M}^{\prime}_{S4}$.
Of course this move requires adjusting the soundness and completeness proofs of Grätz to the new RNmatrix defined over signature $\Sigma$. Let $\vdash_{S4}$ be the consequence relation of the standard Hilbert calculus $\mathcal{H}_{S4}$ for S4 over $\Sigma$. The notion $\Gamma\vdash_{S4}\varphi$ of derivations from premises in $\mathcal{H}_{S4}$ is defined, as in the case of $\mathcal{H}^{b}_{S4}$, in terms of theoremhood.
Let $Val(\mathcal{M}^{\prime}_{S4})$ is the set of valuations over $\mathcal{M}^{\prime}_{S4}$. Consider the family of level $\mathcal{L}^{\prime S4}_{k}$ ($k\geq 0$) as in Definition 3.1, but now starting from ${\mathcal{L}^{\prime}}_{0}^{S4}=Val(\mathcal{M}^{\prime}_{S4})$ in item i), and considering formulas $\alpha\in For(\Sigma)$ in item ii). Let $\mathcal{L}^{\prime}_{S4}=\cap^{\infty}_{k\geq 0}\mathcal{L}^{\prime S4}_{k}$ and $\mathcal{R}(\mathcal{M}^{\prime}_{S4})=\langle\mathcal{M}^{\prime}_{S4},\mathcal{L}^{\prime}_{S4}\rangle$.
Remark 3.10.
Let $\mathcal{M}$ be an Nmatrix. By the very definitions it is immediate to see that, if $\mathcal{F}\subseteq\mathcal{F}^{\prime}\subseteq Val(\mathcal{M})$, then the following holds: $\Gamma\models_{\langle\mathcal{M},\mathcal{F}^{\prime}\rangle}\varphi$ implies that $\Gamma\models_{\langle\mathcal{M},\mathcal{F}\rangle}\varphi$.
Theorem 3.11 (Soundness of $\mathcal{H_{S4}}$ w.r.t. $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$).
For every $\Gamma$ and $\varphi$: $\Gamma\vdash_{S4}\varphi$ implies that $\Gamma\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\varphi$.
Proof.
It coincides with the proof for $\mathcal{R}(\mathcal{M}_{S4})$ found in [4, Theorems 3.4 and 3.5]. The only detail to be checked is that any axiom of $\mathcal{H_{S4}}$ is still valid in the reduced Nmatrix $\mathcal{M}^{\prime}_{S4}$. Details are left to the reader. ∎
Corollary 3.11.1.
If $\vdash_{S4}\varphi$ then $v(\varphi)=2$ for every $v\in\mathcal{L}^{\prime}_{S4}$.
Proof.
Assume that $\vdash_{S4}\varphi$, and let $v\in\mathcal{L}^{\prime}_{S4}$. By Theorem 3.11, $v(\varphi)\in\{1,2\}$. Since $\vdash_{S4}\Box\varphi$ (by necessitation rule) then $v(\Box\varphi)\in\tilde{\Box}v(\varphi)$ such that $v(\Box\varphi)\in\{1,2\}$, by Theorem 3.11. Hence, $v(\varphi)=2$. ∎
For completeness, recall that, given a Tarskian and finitary logic L and a formula $\varphi$, a set $\Delta$ of formulas is said to be $\varphi$-saturated in L if $\Delta\nvdash_{\bf L}\varphi$, but $\Delta,\alpha\vdash_{\bf L}\varphi$, whenever $\alpha\notin\Delta$. It is a well-known result that, if $\Gamma\nvdash_{\bf L}\varphi$, there exists a $\varphi$-saturated set $\Delta$ in L such that $\Gamma\subseteq\Delta$. If $\Delta$ is $\varphi$-saturated in L then it is closed, that is: $\Delta\vdash_{\bf L}\alpha$ iff $\alpha\in\Delta$. In particular, this holds in the logic S4, which is generated by $\mathcal{H_{S4}}$ over $\Sigma$. Since it contains classical logic, any $\varphi$-saturated set $\Delta$ satisfies: $\alpha\vee\beta\in\Delta$ iff $\alpha\in\Delta$ or $\beta\in\Delta$; $\alpha\land\beta\in\Delta$ iff $\alpha\in\Delta$ and $\beta\in\Delta$; and $\alpha\to\beta\in\Delta$ iff either $\alpha\notin\Delta$ or $\beta\in\Delta$. The following result is an adaptation of Lemma 3.7 in [4] to the reduced Nmatrix $\mathcal{M}^{\prime}_{S4}$:
Lemma 3.12.
Let $\Delta\subseteq For(\Sigma)$ be a $\varphi$-saturated set in the logic generated by $\mathcal{H_{S4}}$ (that is, S4). Let $v_{\Delta}:For(\Sigma)\to\mathcal{V}$ be the function defined as follows:
[…]
Then, $v_{\Delta}$ is a valuation over the Nmatrix $\mathcal{M}^{\prime}_{S4}$.
Proof.
[…]
Disjunction: Suppose that $v_{\Delta}(\alpha\vee\beta)=2$. Then, $\Box(\alpha\vee\beta)\in\Delta$ and so $\alpha\vee\beta\in\Delta$, by axiom T: $\Box\psi\to\psi$ and the fact that $\Delta$ is closed under logical inferences in $\mathcal{H_{S4}}$. This implies that $\alpha\in\Delta$ or $\beta\in\Delta$, hence $v_{\Delta}(\alpha)\in\mathcal{D}$ or $v_{\Delta}(\beta)\in\mathcal{D}$. This implies that $v_{\Delta}(\alpha\vee\beta)=2\in\vee^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Suppose now that $v_{\Delta}(\alpha\vee\beta)=1$. Then, $\Box(\alpha\vee\beta)\notin\Delta$ but $\alpha\vee\beta\in\Delta$ and so $\alpha\in\Delta$ or $\beta\in\Delta$. Suppose that $\Box\alpha\in\Delta$ or $\Box\beta\in\Delta$. Then, $\Box\alpha\vee\Box\beta\in\Delta$, and so $\Box(\alpha\vee\beta)\in\Delta$, given that $(\Box\alpha\vee\Box\beta)\to\Box(\alpha\vee\beta)$ is a theorem in $\mathcal{H_{S4}}$. But this is a contradiction, therefore $\Box\alpha\notin\Delta$ and $\Box\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)=1$ and $v_{\Delta}(\beta)\in\{0,1\}$, or vice versa. Then, $v_{\Delta}(\alpha\vee\beta)=1\in\{1,2\}=\vee^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Finally, if $v_{\Delta}(\alpha\vee\beta)=0$ then $\alpha\vee\beta\notin\Delta$, hence $\alpha\notin\Delta$ and $\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\vee\beta)=0\in\{0\}=\vee^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
Conjunction: Suppose that $v_{\Delta}(\alpha\land\beta)=2$. Then, $\Box(\alpha\land\beta)\in\Delta$ and so $\Box\alpha\land\Box\beta\in\Delta$, since $\Box(\alpha\land\beta)\to(\Box\alpha\land\Box\beta)$ is a theorem in $\mathcal{H_{S4}}$. This implies that $\Box\alpha\in\Delta$ and $\Box\beta\in\Delta$, hence $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=2$. Thus, $v_{\Delta}(\alpha\land\beta)=2\in\{2\}=\land^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Now, assume that $v_{\Delta}(\alpha\land\beta)=1$. Then, $\Box(\alpha\land\beta)\notin\Delta$ but $\alpha\land\beta\in\Delta$ and so $\alpha\in\Delta$ and $\beta\in\Delta$ but either $\Box\alpha\notin\Delta$ or $\Box\beta\notin\Delta$ (since $(\Box\alpha\land\Box\beta)\to\Box(\alpha\land\beta)$ is a theorem in $\mathcal{H_{S4}}$). That is, $v_{\Delta}(\alpha),v_{\Delta}(\beta)\in\{1,2\}$ and either $v_{\Delta}(\alpha)=1$ or $v_{\Delta}(\beta)=1$. Then, $v_{\Delta}(\alpha\land\beta)=1\in\{1\}=\land^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Finally, if $v_{\Delta}(\alpha\land\beta)=0$ then $\alpha\land\beta\notin\Delta$, hence $\alpha\notin\Delta$ or $\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)=0$ or $v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\land\beta)=0\in\{0\}=\land^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
Implication: Suppose that $v_{\Delta}(\alpha\to\beta)=2$. Then, $\Box(\alpha\to\beta)\in\Delta$ and so $\alpha\to\beta\in\Delta$, by axiom T. This implies that either $\alpha\notin\Delta$ or $\beta\in\Delta$. If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)=0$. Thus, $v_{\Delta}(\alpha\to\beta)=2\in{\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. If $\beta\in\Delta$, suppose first that $\Box\alpha\in\Delta$. Then, by axiom K it follows that $\Box\beta\in\Delta$. Hence, $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=2$ and so $v_{\Delta}(\alpha\to\beta)=2\in\{2\}={\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Otherwise, if $\Box\alpha\notin\Delta$ then $v_{\Delta}(\alpha)\neq 2$ and $v_{\Delta}(\beta)\neq 0$. From this, $v_{\Delta}(\alpha\to\beta)=2\in{\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Suppose now that $v_{\Delta}(\alpha\to\beta)=1$. Then, $\Box(\alpha\to\beta)\notin\Delta$ but $\alpha\to\beta\in\Delta$ and so $\alpha\notin\Delta$ or $\beta\in\Delta$. In addition, $\Box\beta\notin\Delta$, since $\Box\beta\to\Box(\alpha\to\beta)$ is a theorem in $\mathcal{H_{S4}}$. If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)=0$ and $v_{\Delta}(\beta)\in\{0,1\}$. Thus, $v_{\Delta}(\alpha\to\beta)=1\in\{1,2\}={\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Otherwise, if $\beta\in\Delta$ then $v_{\Delta}(\beta)=1$ and so $v_{\Delta}(\alpha\to\beta)=1\in{\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Finally, if $v_{\Delta}(\alpha\to\beta)=0$ then $\alpha\to\beta\notin\Delta$, hence $\alpha\in\Delta$ and $\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)\in\{1,2\}$ and $v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\to\beta)=0\in\{0\}={\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
This completes the proof. ∎
Theorem 3.13 (Completeness of $\mathcal{H_{S4}}$ w.r.t. $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$).
For every $\Gamma$ and $\varphi$: $\Gamma\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\varphi$ implies that $\Gamma\vdash_{S4}\varphi$.
Proof.
It is analogous to the proof for $\mathcal{R}(\mathcal{M}_{S4})$ found in [4, Lemma 3.9 and Theorem 3.10], but now by using our Lemma 3.12 instead of [4, Lemma 3.7]. ∎
Now, Definitions 3.5 and 3.6 can be easily adapted to $\mathcal{M}^{\prime}_{S4}$, obtaining so the set $PLV^{\prime}$ of partial level valuations in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over finite sets of formulas. However, it will be convenient to give a bit more detailed presentation of these sets, by specifying explicitly the domain of each valuation.
Definition 3.14.
Consider the sets
[…]
Definition 3.15 (Partial valuation in $\mathcal{M}^{\prime}_{S4}$).
Let $\Lambda\in CS(\Sigma)$. A partial valuation in $\mathcal{M}^{\prime}_{S4}$ is a function $\tilde{v}_{p}:\Lambda\rightarrow\mathcal{V}$ such that, for every $\alpha,\beta\in\Lambda$:
[…]
Let $PV(\Lambda)$ be the set of partial valuations in $\mathcal{M}^{\prime}_{S4}$ with domain $\Lambda$.
Definition 3.16 (Partial level valuation over $\Lambda$).
Let $\Lambda\in FCS(\Sigma)$. A partial valuation $\tilde{v}_{p}\in PV(\Lambda)$ is a partial level valuation in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$ iff
  1. $\forall\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=1$, there exists a partial level valuation $\tilde{w}_{p}$ in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$ such that $\tilde{w}_{p}(\alpha)=0$ and, $\forall\beta\in\Lambda$, $\tilde{w}_{p}(\beta)=2$ whenever $\tilde{v}_{p}(\beta)=2$.
The set of partial level valuations in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$ will be denoted by $PLV(\Lambda)$.
Remark 3.17.
Given $v,w$ and $\alpha$ let $\mathsf{P}_{\Lambda}(v,w,\alpha)$ iff $w(\alpha)=0$ and, $\forall\beta\in\Lambda$, $w(\beta)=2$ whenever $v(\beta)=2$. If $\Lambda\in FCS(\Sigma)$ then
$\begin{array}[]{lll}PLV(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in PLV(\Lambda)$}\big{)}\}.\end{array}$
As observed in Remark 3.7 for $\mathcal{M}_{S4}$, since $\Lambda$ is finite then $PV(\Lambda)$ is finite, hence the definition above is not cyclic.
The next step is to show analyticity and co-analyticity w.r.t. level valuations in $\mathcal{M}^{\prime}_{S4}$. In order to do this it is considered, adapting [4, Definition 3.12], the intermediary notion of partial’ level valuation in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$.
Definition 3.18 (Partial’ level valuation over $\Lambda$).
Let $\Lambda\in CS(\Sigma)$. A partial valuation $\tilde{v}_{p}\in PV(\Lambda)$ is a partial’ level valuation in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$ iff
  1. $\forall\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=1$, there exists a level valuation $w$ in $\mathcal{L}^{\prime}_{S4}$ such that $w(\alpha)=0$ and, $\forall\beta\in\Lambda$, $w(\beta)=2$ whenever $\tilde{v}_{p}(\beta)=2$.
The set of partial’ level valuations in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$ will be denoted by $PLV^{\prime}(\Lambda)$.
Remark 3.19.
Consider the predicate $\mathsf{P}_{\Lambda}(v,w,\alpha)$ introduced in Remark 3.17. Clearly,
$\begin{array}[]{lll}PLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $w\in\mathcal{L}^{\prime}_{S4}$}\big{)}\}.\end{array}$
Lemma 3.20 (Co-analyticity lemma for $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$).
Let $v\in\mathcal{L}^{\prime}_{S4}$ and $\Lambda\in CS(\Sigma)$. Then, the restriction $\tilde{v}_{p}:=v_{|\Lambda}$ of $v$ to the domain $\Lambda$ belongs to $PLV^{\prime}(\Lambda)$.
Proof.
It is an adaptation of the proof of [4, Lemma 3.13]. Suppose, by contradiction, that $\tilde{v}_{p}\notin PLV^{\prime}(\Lambda)$. Let $\Gamma:=\{\beta\in\Lambda\mid\tilde{v}_{p}(\beta)=2\}$. Then, there exists $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=1$ and, for every $w\in\mathcal{L}^{\prime}_{S4}$, $w(\alpha)=0$ implies that $w(\beta)\neq 2$ for some $\beta\in\Gamma$. Hence, for every $w\in\mathcal{L}^{\prime}_{S4}$, $w(\alpha)\not\in\mathcal{D}$ implies that $w(\Box\beta)\notin\mathcal{D}$ for some $\beta\in\Gamma$. By contraposition, this means that $\{\Box\beta\mid\beta\in\Gamma\}\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\alpha$. By Theorem 3.13, $\{\Box\beta\mid\beta\in\Gamma\}\vdash_{S4}\alpha$. Observe that $\nvdash_{S4}\alpha$: otherwise, $v(\alpha)=2$, by Corollary  3.11.1. Hence, there exist $\beta_{1},\ldots,\beta_{n}\in\Gamma$ such that $\vdash_{S4}\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots))$. By necessitation rule, $\vdash_{S4}\Box(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots)))$. By Theorem 3.11, $\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\Box(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots)))$. But $v(\Box\beta_{i})=2$ for $1\leq i\leq n$ and $v(\alpha)=1$, hence $v(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots)))=1$. This means that $v(\Box(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots))))=0$, a contradiction. From this we infer that $\tilde{v}_{p}\in PLV^{\prime}(\Lambda)$. ∎
Definition 3.21.
[…]
Remark 3.22.
Observe that it is possible to define an enumeration $\alpha_{1},\alpha_{2},\ldots\alpha_{m},\ldots$ (for $m\in\omega^{2}$) of $For(\Sigma)$ such that $\textsf{c}_{1}(\alpha_{i})\leq\textsf{c}_{1}(\alpha_{j})$ if $i\leq j$ and, for every $i$ such that $\textsf{c}_{1}(\alpha_{i})>0$, if $\beta$ is a strict subformula of $\alpha_{i}$ then $\beta=\alpha_{j}$ for some $j<i$. Such an enumeration of $For(\Sigma)$ can be defined as follows: every formula have and index in $\omega^{2}$, which is a denumerable ordinal, such that all the formulas with complexity 00 (which form a denumerable set) are placed first, with indexes in $I_{0}:=\omega$; after this, all the formulas with complexity $1$ (which form a denumerable set) are placed with an index in $I_{1}:=\omega\cdot 2\setminus\omega=\{\omega,\omega+1,\ldots\}$; in general, the formulas with complexity $n$ (which form a denumerable set) have an index in $I_{n}:=\omega\cdot(n+1)\setminus\omega\cdot n=\{\omega\cdot n,\omega\cdot n+1,\ldots\}$.
Lemma 3.23 (Analyticity lemma for $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$).
Let $\Lambda\in CS(\Sigma)$ and $\tilde{v}_{p}\in PLV^{\prime}(\Lambda)$. Then, there exists $v\in\mathcal{L}^{\prime}_{S4}$ such that the restriction $v_{|\Lambda}$ of $v$ to the domain $\Lambda$ coincides with $\tilde{v}_{p}$.
Proof.
It is an adaptation of the proof of [4, Lemma 3.14]. In this case, more modifications are required because of the conjunction and disjunction connectives, and by the fact of considering a fixed set $\Lambda$.
Suppose first that $\Lambda=For(\Sigma)$. By induction on $n\in\omega$, it will be shown that $\tilde{v}_{p}\in\mathcal{L}^{\prime S4}_{n}$ for every $n$. By Definition 3.15 of partial valuation, $\tilde{v}_{p}\in Val(\mathcal{M}^{\prime}_{S4})={\mathcal{L}^{\prime}}_{0}^{S4}$. Suppose that $\tilde{v}_{p}\in\mathcal{L}^{\prime S4}_{n}$ for a given $n\geq 0$, and let $\alpha$ such that $\mathcal{L}^{\prime S4}_{n}(\alpha)\subseteq\mathcal{D}$. In particular, $\tilde{v}_{p}(\alpha)\in\mathcal{D}$. Suppose that $\tilde{v}_{p}(\alpha)=1$. By Definition 3.18 of partial’ level valuation, there exists a level valuation $w$ in $\mathcal{L}^{\prime}_{S4}$ such that $w(\alpha)=0$. But this is a contradiction, since $w\in\mathcal{L}^{\prime S4}_{n}$. Then $\tilde{v}_{p}(\alpha)=2$ and so $\tilde{v}_{p}\in\mathcal{L}^{\prime S4}_{n+1}$. From this, $\tilde{v}_{p}\in\mathcal{L}^{\prime}_{S4}$ and the result clearly holds.
Suppose now that $\Lambda\neq\emptyset$ is a proper subset of $For(\Sigma)$. Consider an enumeration of $For(\Sigma)$ as in Remark 3.22. The valuation $v$ will be defined by induction on the complexity $n$ of the formulas. Observe that $\Lambda\cup\mathcal{P}\in CS(\Sigma)$ and, for every $n\in\omega$, $\Lambda_{n}\in CS(\Sigma)$ for $\Lambda_{n}:=\Lambda\cup For(\Sigma)_{n}$. In addition, for every $m\in\omega^{2}$ and $n\geq 0$, if $\alpha_{m}$ is such that $\textsf{c}_{1}(\alpha_{m})=n+1$, then $\Lambda_{n}^{m}\in CS(\Sigma)$, where $\Lambda_{n}^{m}=\Lambda_{n}\cup\{\alpha_{m}\}$. Clearly $\Lambda_{n+1}=\cup_{k\in\omega}\Lambda_{n}^{\omega\cdot(n+1)+k}$.
Base $n=0$: Define $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v(p)=0$ if $p$ is a propositional variable which does not belong to $\Lambda$. Clearly, $v\in PLV^{\prime}(\Lambda\cup\mathcal{P})$ up to this stage.
Inductive step: Assume that $v(\alpha)$ was defined for every $\alpha\in\Lambda_{n}$ such that $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v\in PLV^{\prime}(\Lambda_{n})$ up to this stage, for a given $n\geq 0$ (Induction Hypothesis, IH). Let $\alpha=\alpha_{m}$ (for $m\in I_{n+1}$) such that $\alpha\not\in\Lambda$ and $\textsf{c}_{1}(\alpha)=n+1$. Observe that $v(\beta)$ was already defined, for every strict subformula $\beta$ of $\alpha$. Since $v\in PLV^{\prime}(\Lambda_{n})$ then, for every $\delta\in\Lambda_{n}$ such that $v(\delta)=1$ there exists $w\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}_{\Lambda_{n}}(v,w,\delta)$. Such a $w$ will be denoted by $w^{\delta}$ (observe that it is possible to have more that one $w^{\delta}$ for each $\delta$, and it is possible to have $w^{\delta}=w^{\gamma}$ for $\delta\neq\gamma$). Recall that $\Lambda_{n}^{m}=\Lambda_{n}\cup\{\alpha_{m}\}$ is closed under subformulas. There are four cases to analyze:
[…]
(2.1) There is $\delta\in\Lambda_{n}$ with $v(\delta)=1$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=0$. Define $v(\alpha)=1$. Hence $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\delta}$.
(2.2) There is $\delta\in\Lambda_{n}$ with $v(\delta)=1$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=1$. By Lemma 3.20, $v^{\prime}=w^{\delta}_{|\Lambda_{n}^{m}}\in PLV^{\prime}(\Lambda_{n}^{m})$, hence there exists $w^{\prime\prime}\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}_{\Lambda_{n}^{m}}(v^{\prime},w^{\prime\prime},\alpha)$. Define $v(\alpha)=1$. Hence, $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\prime\prime}$.
(2.3) For every $\delta\in\Lambda_{n}$ such that $v(\delta)=1$, $w^{\delta}(\alpha)=2$ for every $w^{\delta}$. Define $v(\alpha)=2$. Clearly, $v\in PLV^{\prime}(\Lambda_{n}^{m})$.
[…]
(3.1) $\alpha=\beta\to\gamma$. Then $v(\beta)=2$ and $v(\gamma)=1$. Define $v(\alpha)=1$. Then $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\gamma}$. Indeed, since $w^{\gamma}(\beta)=2$ and $w^{\gamma}(\gamma)=0$ then $w^{\gamma}(\alpha)=0$.
(3.2) $\alpha=\beta\land\gamma$. Then $v(\beta)=1$ and $v(\gamma)\in\mathcal{D}$ or vice versa. Then, there exists $w\in\mathcal{L}^{\prime}_{S4}$ (where $w=w^{\beta}$ or $w=w^{\gamma}$) such that $w(\beta)=0$ or $w(\gamma)=0$ and so $w(\alpha)=0$. Define $v(\alpha)=1$. Then $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\beta}$ or $w^{\alpha}=w^{\gamma}$.
[…]
We have shown in cases (1)-(4) how to extend the domain of $v$ to the additional formula $\alpha_{m}$, for every $\alpha_{m}$ such that $\textsf{c}_{1}(\alpha_{m})=n+1$, showing that $v\in PLV^{\prime}(\Lambda_{n}^{m})$. Consider now the simultaneous extension of $v$ with the values assigned to every $\alpha_{m}$ as prescribed above. We obtain so an extension of $v$ to $\Lambda_{n+1}=\cup_{m}\Lambda_{n}^{m}$. By the very definition of partial’ level valuations it follows that $v\in PLV^{\prime}(\Lambda_{n+1})$, and $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$. This completes the inductive step. ∎
Proposition 3.24.
Let $\Lambda\in FCS(\Sigma)$. Then $PLV(\Lambda)=PLV^{\prime}(\Lambda)$.
Proof.
Let $\tilde{v}_{p}\in PLV^{\prime}(\Lambda)$. Then, $\tilde{v}_{p}\in PV(\Lambda)$. Suppose that $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=1$. Then, there exists $w\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)$. Let $\tilde{w}_{p}=w_{|\Lambda}$. By Lemma 3.20, $\tilde{w}_{p}\in PLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)$. Now, suppose that there exists $\tilde{w}^{\prime}_{p}\in PLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}^{\prime}_{p},\alpha)$. By Lemma 3.23, there exists $v^{\prime}\in\mathcal{L}^{\prime}_{S4}$ such that the restriction of $v^{\prime}$ to $\Lambda$ coincides with $\tilde{w}^{\prime}_{p}$. Hence, $\mathsf{P}_{\Lambda}(\tilde{v}_{p},v^{\prime},\alpha)$. This shows that
$\begin{array}[]{lll}PLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in PLV^{\prime}(\Lambda)$}\big{)}\}.\end{array}$ From this, being $PV(\Lambda)$ finite, it follows that $PLV(\Lambda)=PLV^{\prime}(\Lambda)$. ∎
Theorem 3.25.
Let $\varphi\in For(\Sigma)$, and let $\Lambda$ be the set of subformulas of $\varphi$. Then: $\vdash_{S4}\varphi$ if, and only if, for every partial valuation $\tilde{v}_{p}\in PLV(\Lambda)$ in $\mathcal{M}^{\prime}_{S4}$, $\tilde{v}_{p}(\varphi)=2$.
Proof.
Let $\Lambda$ be the set of subformulas of $\varphi$. Then, $\Lambda\in FCS(\Sigma)$. Suppose that $\vdash_{S4}\varphi$. By Theorem 3.11, $\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\varphi$. Let $\tilde{v}_{p}\in PLV(\Lambda)$. By Proposition 3.24, $\tilde{v}_{p}\in PLV^{\prime}(\Lambda)$. By Lemma 3.23, there exists $v\in\mathcal{L}^{\prime}_{S4}$ such that the restriction $v_{|\Lambda}$ of $v$ to $\Lambda$ coincides with $\tilde{v}_{p}$. Then, $\tilde{v}_{p}(\varphi)=v(\varphi)=2$, by Corollary 3.11.1.
Conversely, suppose that $\tilde{v}_{p}(\varphi)=2$ for every $\tilde{v}_{p}\in PLV(\Lambda)$. By Proposition 3.24, $\tilde{v}_{p}(\varphi)=2$ for every $\tilde{v}_{p}\in PLV^{\prime}(\Lambda)$. Let $v\in\mathcal{L}^{\prime}_{S4}$. By Lemma 3.20, the restriction $\tilde{v}_{p}:=v_{|\Lambda}$ of $v$ to $\Lambda$ belongs to $PLV^{\prime}(\Lambda)$. Then, $v(\varphi)=\tilde{v}_{p}(\varphi)=2$ and then $\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\varphi$. By Theorem 3.13, $\vdash_{S4}\varphi$. ∎
This means that Grätz’s algorithm can be adapted to the reduced (and expanded with disjunction and conjunction) RNmatrix $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$.
4 An RNmatrix for IPL
Definition 4.1.
Consider the signature Ω={{¬},{,,}}normal-$\Omega=\{\{\neg\},\{\land,\lor,\rightarrow\}\}$ for intuitionistic propositional logic (IPL). The set of intuitionistic formulas will be denoted by as $For(\Omega)$.
4.1 Gödel-McKinsey-Tarski Box Translation
Definition 4.2.
(Box translation function) Let $f_{\Box}:For(\Omega)\rightarrow For(\Sigma)$ be a function such that:
  1. i)
    $f_{\Box}(p)=\Box p$, if $p$ is a propositional variable;
  2. ii)
    $f_{\Box}(\neg\alpha)=\Box\neg f_{\Box}(\alpha)$;
  3. iii)
[…]
It is widely known that this translation is faithful [9].
4.2 Truth tables for intuitionistic logic
We introduce the truth tables for intuitionistic logic based on the presented translation.
4.2.1 Intuitionistic non-deterministic valuations
Remark 4.3.
If $v$ is a valuation in $\mathcal{M}_{S4}$, then, for every $\alpha\in For(\Sigma)$, $v(\Box\alpha)=\Box^{\mathcal{A}_{S4}}(v(\alpha))$. This follows from the fact that the multioperation $\Box^{\mathcal{A}_{S4}}$ is deterministic.
Definition 4.4.
We define the function $h:For(\Omega)\rightarrow For(\Sigma)$ as follows:
  1. i)
    $h(p)=p$, if $p$ is a propositional variable;
  2. ii)
    $h(\neg\alpha)=\neg f_{\Box}(\alpha)$;
  3. iii)
[…]
We call $h$ as the semi-translation function.
Lemma 4.5.
For every $\alpha\in For(\Omega)$, $f_{\Box}(\alpha)=\Box(h(\alpha))$.
Proof.
Let $\alpha$ be an intuitionistic formula. We proceed by case analysis.
Case $\alpha=p_{n}$
In this case, $f_{\Box}(\alpha)=f_{\Box}(p_{n})=\Box p_{m}$. Also, $\Box(h(\alpha))=\Box(h(p_{n}))=\Box p_{m}$. Note that $n=m$ in both cases. Hence, $f_{\Box}(p_{n})=\Box(h(p_{n}))$.
Case $\alpha=\neg\beta$
By Definition 4.4, $\Box(h(\neg\beta))=\Box(\neg f_{\Box}(\beta))$. Moreover, by Definition 4.2, $f_{\Box}(\neg\beta)=\Box(\neg f_{\Box}(\beta))$. Hence, $f_{\Box}(\neg\beta)=\Box(h(\neg\beta))$.
[…]
By Definition 4.4, we have that for each ##\##, $\Box(h(\beta\#\gamma))=\Box(f_{\Box}(\beta)\#f_{\Box}(\gamma))$. Moreover, by Definition 4.2, $f_{\Box}(\beta\#\gamma)=\Box(f_{\Box}(\beta)\#f_{\Box}(\gamma))$. Therefore, $f_{\Box}(\beta\#\gamma)=\Box(h(\beta\#\gamma))$.
Remark 4.6.
By Lemma 4.5 and Remark 4.3, $v(f_{\Box}(\alpha))=\Box^{\mathcal{A}_{S4}}(v(h(\alpha)))$, for every $\alpha\in For(\Omega)$.
Definition 4.7.
(Truth function) Let $v$ be a valuation function in $\mathcal{M}_{S4}$ and $\alpha\in For(\Omega)$. We define the function $\tilde{v}:For(\Omega)\rightarrow V\times V$ as follows:
$\tilde{v}(\alpha)\triangleq\langle v(h(\alpha));v(f_{\Box}(\alpha))\rangle$
We will say that $\tilde{v}$ is induced by $v$ and will call $\tilde{v}$ as the truth function.
By Remark 4.6, we get $\tilde{v}(\alpha)=\langle v(h(\alpha));\Box^{\mathcal{A}_{S4}}(v(h(\alpha)))\rangle$. This motivates the following:
Definition 4.8 (Intuitionistic truth values).
The set of intuitionistic truth values is given by
$V_{IPL}=\left\{\langle c;\Box^{\mathcal{A}_{S4}}(c)\rangle\mid c\in\mathcal{V}\right\}=\{\mathbf{F},\mathbf{U},\mathbf{T}\}$
where $\mathbf{F}=\langle 0;0\rangle$, $\mathbf{U}=\langle 1;0\rangle$ and $\mathbf{T}=\langle 2;2\rangle$.
Therefore, the truth function 4.7 is a function $\tilde{v}:For(\Omega)\rightarrow\{\mathbf{F},\mathbf{U},\mathbf{T}\}$.
[…]
Definition 4.9 (Intuitionistic multioperations).
Let $\langle a,b\rangle,\langle c,d\rangle\in V_{IPL}$. Define the intuitionistic multioperations as follows:
  1. 1.
    $\neg^{\prime\mathcal{A}_{IPL}}(\langle a;b\rangle)=\left\{\langle c;\Box^{\mathcal{A}_{S4}}(c)\rangle\mid c\in\neg^{\mathcal{A}_{S4}}(b)\right\}$
  2. 2.
    #$\#^{\prime\mathcal{A}_{IPL}}(\langle a;b\rangle,\langle c;d\rangle)=\left\{\langle e;\Box^{\mathcal{A}_{S4}}(e)\rangle\mid e\in\#^{\mathcal{A}^{\prime}_{S4}}(b,d)\right\}$
[…]
Definition 4.10.
The intuitionistic Nmatrix over signature $\Omega$ is given by $\mathcal{M}_{IPL}=\langle V_{IPL},\mathcal{D}_{IPL},\mathcal{O}_{IPL}\rangle$, where $\mathcal{D}_{IPL}=\{\mathbf{T}\}$, $\mathcal{O}_{IPL}=\{\neg^{\mathcal{A}_{IPL}},\rightarrow^{\mathcal{A}_{IPL}},\lor^{\mathcal{A}_{IPL}},\land^{\mathcal{A}_{IPL}}\}$ and each $o\in\mathcal{O}_{IPL}$ is defined in Table 3.
$\neg^{\mathcal{A}_{IPL}}$
[…]
$\rightarrow^{\mathcal{A}_{IPL}}$ F U T
[…]
$\lor^{\mathcal{A}_{IPL}}$ F U T
[…]
$\land^{\mathcal{A}_{IPL}}$ F U T
[…]
Table 3: Tables for #$\#^{\mathcal{A}_{IPL}}$.
Lemma 4.11.
Let #$\#\in\Omega$. Then, #$\#^{\mathcal{A}_{IPL}}=\#^{\prime\mathcal{A}_{IPL}}$. That is, the multioperators in Definition 4.9 coincide with the ones in Definition 4.10.
Proof.
It is straightforward to verify that, for every $z,w\in V_{IPL}$:
  1. 1.
    $\neg^{\mathcal{A}_{IPL}}(z)=\neg^{\prime\mathcal{A}_{IPL}}(z)$; and
  2. 2.
[…]
Definition 4.12.
[…]
Definition 4.13.
Given $n\geq 0$, the set of formulas over $\Omega$ with complexity less or equal than $n$ will be denoted by $For(\Omega)_{n}$. That is: $For(\Omega)_{n}=\{\alpha\in For(\Omega)\mid\textsf{c}(\alpha)\leq n\}$.
Definition 4.14.
Let $Val(\mathcal{M}_{IPL})$ be the set of valuations over $\mathcal{M}_{IPL}$, and let $Val_{+}(\mathcal{M}_{IPL})$ be the set of valuations $v$ in $Val(\mathcal{M}_{IPL})$ such that $v(p)\in\{\mathbf{T},\mathbf{F}\}$ for every propositional variable $p$.
Definition 4.15.
(Level valuations for IPL) Let $Val(\mathcal{M}_{IPL})$ be the set of valuations over $\mathcal{M}_{IPL}$. We define the level $\mathcal{L}^{IPL}_{k}$, where $k\in\mathbb{N}$, as follows:
  1. i)
    $\mathcal{L}_{0}^{IPL}=Val_{+}(\mathcal{M}_{IPL})$;
  2. ii)
    $\mathcal{L}^{IPL}_{k+1}=\{v\in\mathcal{L}^{IPL}_{k}\mid\forall\alpha\in For(\Omega)_{k+1}$, if $v(\alpha)=\mathbf{U}$ then there exists $w\in\mathcal{L}^{IPL}_{k}$ such that $w(\alpha)=\mathbf{F}$ and: $v(\beta)=\mathbf{T}$ implies $w(\beta)=\mathbf{T}$ for every $\beta\in For(\Omega)\}$.
Definition 4.16.
The set of level valuations in $\mathcal{M}_{IPL}$ is defined as the intersection of every level valuation:
$\mathcal{L}_{IPL}=\cap^{\infty}_{k\geq 0}\mathcal{L}^{IPL}_{k}$
Definition 4.17 (RNmatrix for IPL).
Let $\mathcal{M}_{IPL}$ be the Nmatrix defined in Definition 4.10. The RNmatrix for IPL is the pair $\mathcal{R}(\mathcal{M}_{IPL})=\langle\mathcal{M}_{IPL},\mathcal{L}_{IPL}\rangle$.
From now on, $H_{IPL}$ will denote the standard Hilbert calculus for IPL over $\Omega$, given by the following axioms, and with Modus Ponens (MP) as the only inference rule:
$\begin{array}[]{ll}\textsf{(Ax1)}&\alpha\to(\beta\to\alpha)\\[5.69054pt] \textsf{(Ax2)}&(\alpha\to(\beta\to\gamma))\to((\alpha\to\beta)\to(\alpha\to\gamma))\\[5.69054pt] \textsf{(Ax3)}&\alpha\to(\beta\to(\alpha\land\beta))\\[5.69054pt] \textsf{(Ax4)}&(\alpha\land\beta)\to\alpha\\[5.69054pt] \textsf{(Ax5)}&(\alpha\land\beta)\to\beta\\[5.69054pt] \textsf{(Ax6)}&\alpha\to(\alpha\lor\beta)\\[5.69054pt] \textsf{(Ax7)}&\beta\to(\alpha\lor\beta)\\[5.69054pt] \textsf{(Ax8)}&(\alpha\to\gamma)\to((\beta\to\gamma)\to((\alpha\lor\beta)\to\gamma))\\[5.69054pt] \textsf{(Ax9)}&(\beta\to\alpha)\to((\beta\to\neg\alpha)\to\neg\beta)\\[5.69054pt] \textsf{(Ax10)}&\alpha\to(\neg\alpha\to\beta)\\[5.69054pt] \end{array}$
The consequence relation of $H_{IPL}$ will be denoted by $\vdash_{IPL}$. Observe that derivations in $H_{IPL}$ from premises are defined as usual in Hilbert calculi. That is, premises can be used in the derivations in the same way as axioms.
The proof of soundness of $H_{IPL}$ w.r.t. the RNmatrix $\mathcal{R}(\mathcal{M}_{IPL})$ will be postponed to Subsection 4.2.2. We will concentrate now on the completeness proof.
Remark 4.18.
Recall from Subsection 3.2 the notion of $\varphi$-saturated set in L. If $\Delta$ is a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$ (that is, IPL) then: $\alpha\in\Delta$ implies that $\neg\alpha\notin\Delta$; $\alpha\vee\beta\in\Delta$ iff $\alpha\in\Delta$ or $\beta\in\Delta$; $\alpha\land\beta\in\Delta$ iff $\alpha\in\Delta$ and $\beta\in\Delta$; $\alpha\to\beta\in\Delta$ and $\alpha\in\Delta$ implies that $\beta\in\Delta$; and $\beta\in\Delta$ implies that $\alpha\to\beta\in\Delta$.
Definition 4.19 (Valuation associated to a $\varphi$-saturated set in IPL).
Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$ (that is, IPL). The valuation associated to $\Delta$ in IPL is the function $v_{\Delta}:For(\Omega)\to V_{IPL}$ defined inductively as follows (here, $p$ is a propositional variable):
[…]
Remark 4.20.
Observe that $v_{\Delta}(\neg\alpha)=\mathbf{U}$ whenever $\alpha$ is undetermined in $\Delta$, that is: $\neg\alpha\notin\Delta$ and $\alpha\notin\Delta$. Hence, the failure of excluded middle (which is a principle valid in classical logic but not in IPL) occurs only for formulas $\alpha$ with such undetermined status in $\Delta$. Analogously, $v_{\Delta}(\alpha\to\beta)=\mathbf{U}$ whenever $\alpha$ and $\alpha\to\beta$ are jointly undetermined in $\Delta$, that is: neither $\alpha$ nor $\alpha\to\beta$ belong to $\Delta$. Hence, the failure of $\alpha\vee(\alpha\to\beta)$ (which is a principle valid in classical logic but not in IPL) occurs only for formulas $\alpha$ and $\alpha\to\beta$ with such joint undetermined status in $\Delta$.
Proposition 4.21.
Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$. Then, the valuation $v_{\Delta}$ associated to $\Delta$ in IPL is a valuation in the Nmatrix $\mathcal{M}_{IPL}$ which belongs to $Val_{+}(\mathcal{M}_{IPL})$ (recall Definition 4.14) such that, for every $\alpha$: $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$.
Proof.
Observe that, by definition and by Remark 4.18, $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$, for every formula $\alpha$. Moreover, $v_{\Delta}(p)\in\{\mathbf{T},\mathbf{F}\}$ for every propositional variable $p$.
Negation: Let us prove first that $v_{\Delta}(\neg\alpha)\in\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$ for every $\alpha$. If $v_{\Delta}(\neg\alpha)=\mathbf{T}$ then $\neg\alpha\in\Delta$, by definition. Hence, $\alpha\notin\Delta$ and so $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$. From this, $v_{\Delta}(\neg\alpha)=\mathbf{T}\in\{\mathbf{U},\mathbf{T}\}=\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$. Now, if $v_{\Delta}(\neg\alpha)=\mathbf{U}$ then $\alpha\notin\Delta$, by definition. Hence $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and so $v_{\Delta}(\neg\alpha)=\mathbf{U}\in\{\mathbf{U},\mathbf{T}\}=\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$. Finally, if $v_{\Delta}(\neg\alpha)=\mathbf{F}$ then $\alpha\in\Delta$, by definition. Hence $v_{\Delta}(\alpha)=T$ and so $v_{\Delta}(\neg\alpha)=\mathbf{F}\in\{\mathbf{F}\}=\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$.
Implication: There are 3 cases:
(I1) Suppose that $v_{\Delta}(\alpha\to\beta)=\mathbf{T}$. Then $\alpha\to\beta\in\Delta$, by definition.
– If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{T}\in{\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
– If $\alpha\in\Delta$ then $\beta\in\Delta$, by MP. Hence $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=\mathbf{T}$ and so $v_{\Delta}(\alpha\to\beta)=\mathbf{T}\in\{\mathbf{T}\}={\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(I2) Suppose that $v_{\Delta}(\alpha\to\beta)=\mathbf{U}$. Then $\alpha\to\beta\notin\Delta$ and $\alpha\notin\Delta$, by definition. Moreover, $\beta\notin\Delta$, by Remark 4.18. Then $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{U}\in\{\mathbf{U},\mathbf{T}\}={\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(I3) Suppose that $v_{\Delta}(\alpha\to\beta)=\mathbf{F}$. Then $\alpha\to\beta\notin\Delta$ and $\alpha\in\Delta$, by definition. Moreover, $\beta\notin\Delta$, by Remark 4.18. Then $v_{\Delta}(\alpha)=\mathbf{T}$ and $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{F}\in\{\mathbf{F}\}={\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
Conjunction: There are 2 cases:
(C1) Suppose that $v_{\Delta}(\alpha\land\beta)=\mathbf{T}$. Then $\alpha\in\Delta$ and $\beta\in\Delta$, by definition. From this, $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=\mathbf{T}$. Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{T}\in\{\mathbf{T}\}=\land^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(C2) Suppose that $v_{\Delta}(\alpha\land\beta)=\mathbf{F}$. Then, either $\alpha\notin\Delta$ or $\beta\notin\Delta$, by definition. From this, $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ or $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{F}\in\{\mathbf{F}\}=\land^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
Disjunction: There are 2 cases:
(D1) Suppose that $v_{\Delta}(\alpha\vee\beta)=\mathbf{T}$. Then, either $\alpha\in\Delta$ or $\beta\in\Delta$, by definition. From this, $v_{\Delta}(\alpha)=\mathbf{T}$ or $v_{\Delta}(\beta)=\mathbf{T}$. Hence, $v_{\Delta}(\alpha\vee\beta)=\mathbf{T}\in\{\mathbf{T}\}=\vee^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(D2) Suppose that $v_{\Delta}(\alpha\vee\beta)=\mathbf{F}$. Then $\alpha\notin\Delta$ and $\beta\notin\Delta$, by definition. From this, $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{F}\in\{\mathbf{F}\}=\vee^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
This concludes the proof. ∎
Proposition 4.22.
Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$. Then, the valuation $v_{\Delta}$ associated to $\Delta$ in IPL is a level valuation in $\mathcal{M}_{IPL}$ such that, for every $\alpha$: $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$.
Proof.
By Proposition 4.21, $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$ for every $\alpha$. It will be shown, by induction on $n$, that $v_{\Delta}\in\mathcal{L}^{IPL}_{n}$ for every $n\geq 0$.
Base $n=0$: $v_{\Delta}\in\mathcal{L}^{IPL}_{0}=Val_{+}(\mathcal{M}_{IPL})$, by Proposition 4.21.
[…]
(1) $\alpha=\neg\beta$ for some $\beta$. By Definition 4.19, $\neg\beta\not\in\Delta$ and $\beta\not\in\Delta$. Suppose that $\Delta,\beta\vdash_{IPL}\neg\beta$. By the deduction metatheorem, $\Delta\vdash_{IPL}\beta\to\neg\beta$. By (Ax9) and by MP, $\Delta,\beta\to\beta,\beta\to\neg\beta\vdash_{IPL}\neg\beta$. But $\beta\to\beta\in\Delta$ and $\beta\to\neg\beta\in\Delta$, hence $\Delta\vdash_{IPL}\neg\beta$, that is, $\neg\beta\in\Delta$, a contradiction. From this, $\Delta,\beta\nvdash_{IPL}\neg\beta$. Then, there exists a $\neg\beta$-saturated set $\Delta^{\prime}$ in IPL such that $\Delta\cup\{\beta\}\subseteq\Delta^{\prime}$. By IH, $v_{\Delta^{\prime}}\in\mathcal{L}^{IPL}_{n}$ such that, by Definition 4.19, $v_{\Delta^{\prime}}(\neg\beta)=\mathbf{F}$. Moreover, for every $\delta\in For(\Omega)$, if $v_{\Delta}(\delta)=\mathbf{T}$ then $v_{\Delta^{\prime}}(\delta)=\mathbf{T}$, since $\Delta\subseteq\Delta^{\prime}$.
(2) $\alpha=\beta\to\gamma$ for some $\beta$ and $\gamma$. By Definition 4.19, $\beta\to\gamma\not\in\Delta$ and $\beta\not\in\Delta$. Suppose that $\Delta,\beta\vdash_{IPL}\beta\to\gamma$. By MP, $\Delta,\beta\vdash_{IPL}\gamma$ and so, by the deduction metatheorem, $\Delta\vdash_{IPL}\beta\to\gamma$, a contradiction. From this, $\Delta,\beta\nvdash_{IPL}\beta\to\gamma$. Then, there exists a $(\beta\to\gamma)$-saturated set $\Delta^{\prime\prime}$ in IPL such that $\Delta\cup\{\beta\}\subseteq\Delta^{\prime\prime}$. By IH, $v_{\Delta^{\prime\prime}}\in\mathcal{L}^{IPL}_{n}$ such that, by Definition 4.19, $v_{\Delta^{\prime\prime}}(\beta\to\gamma)=\mathbf{F}$. Moreover, for every $\delta\in For(\Omega)$, if $v_{\Delta}(\delta)=\mathbf{T}$ then $v_{\Delta^{\prime\prime}}(\delta)=\mathbf{T}$, since $\Delta\subseteq\Delta^{\prime\prime}$.
From the previous analysis we conclude that $v_{\Delta}\in\mathcal{L}^{IPL}_{n+1}$. This shows that $v_{\Delta}\in\mathcal{L}^{IPL}_{n}$ for every $n\geq 0$ and so $v_{\Delta}\in\mathcal{L}_{IPL}$. ∎
Theorem 4.23 (Completeness of $\mathcal{H}_{IPL}$ w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$).
For every $\Gamma$ and $\varphi$: $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$ implies that $\Gamma\vdash_{IPL}\varphi$.
Proof.
Suppose that $\Gamma\nvdash_{IPL}\varphi$. Then, there exists a $\varphi$-saturated set $\Delta$ in IPL such that $\Gamma\subseteq\Delta$. By Proposition 4.22, $v_{\Delta}$ is a level valuation in $\mathcal{M}_{IPL}$ such that $v_{\Delta}(\alpha)=\mathbf{T}$ for every $\alpha\in\Gamma$, but $v_{\Delta}(\varphi)\neq\mathbf{T}$. This shows that $\Gamma\not\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$. ∎
4.2.2 Intuitionistic partial (level) valuations and intuitionistic truth tables
Definition 4.24.
Consider the sets
[…]
Definition 4.25 (Partial valuation in $\mathcal{M}_{IPL}$).
Let $\Lambda\in CS(\Omega)$. A partial valuation in $\mathcal{M}_{IPL}$ is a function $\tilde{v}_{p}:\Lambda\rightarrow V_{IPL}$ such that, for every $\alpha,\beta\in\Lambda$:
  • if $\alpha\in\mathcal{P}\cap\Lambda$ then $\tilde{v}_{p}(\alpha)\in\{\mathbf{T},\mathbf{F}\}$;
  • if $\neg\alpha\in\Lambda$ then $\tilde{v}_{p}(\neg\alpha)\in\neg^{\mathcal{A}_{IPL}}(\tilde{v}_{p}(\alpha))$;
[…]
Let $iPV(\Lambda)$ be the set of partial valuations in $\mathcal{M}_{IPL}$ with domain $\Lambda$.
Definition 4.26 (Intuitionistic partial level valuation over $\Lambda$).
Let $\Lambda\in FCS(\Omega)$. A partial valuation $\tilde{v}_{p}\in iPV(\Lambda)$ is an intuitionistic partial level valuation in $\mathcal{M}_{IPL}$ over $\Lambda$ iff
  1. $\forall\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$, there exists an intuitionistic partial level valuation $\tilde{w}_{p}$ in $\mathcal{M}_{IPL}$ over $\Lambda$ such that $\tilde{w}_{p}(\alpha)=\mathbf{F}$ and, $\forall\beta\in\Lambda$, $\tilde{w}_{p}(\beta)=\mathbf{T}$ whenever $\tilde{v}_{p}(\beta)=bT$.
The set of intuitionistic partial level valuations in $\mathcal{M}_{IPL}$ over $\Lambda$ will be denoted by $iPLV(\Lambda)$.
Remark 4.27.
Given $v,w$ and $\alpha$ let $\mathsf{P}_{\Lambda}^{i}(v,w,\alpha)$ iff $w(\alpha)=\mathbf{F}$ and, $\forall\beta\in\Lambda$, $w(\beta)=\mathbf{T}$ whenever $v(\beta)=\mathbf{T}$. If $\Lambda\in FCS(\Omega)$ then
$\begin{array}[]{lll}iPLV(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in iPLV(\Lambda)$}\big{)}\}.\end{array}$
As observed in Remark 3.7 for $\mathcal{M}_{S4}$, since $\Lambda$ is finite then $iPV(\Lambda)$ is finite, hence the definition of $iPLV(\Lambda)$ is not cyclic.
Definition 4.28.
An intuitionistic truth table in $\mathcal{M}_{IPL}$ is an exhaustive list of all intuitionistic partial level valuations for a given finite domain closed under subformulas.
Remark 4.29.
Each of these partial valuations corresponds to a row of the table.
The following is an example of the algorithm for searching the (intuitionistic) partial level valuations. [^5] This example, as the ones presented in Appendix A, were automatically generated by a Coq module for RNmatrices written by R. Leme, see  https://github.com/renatoleme/Forest We start with the full table for $\neg\neg(p\lor\neg p)$.
Table 4: Initial table.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(1) F U F U U
(2) F U F U T
(3) F U F T F
(4) F T T F U
(5) F T T F T
(6) T F T F U
(7) T F T F T
Table 5: First cycle.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(1) $\times$ 6, 7 $\times$ 4, 5, 6, 7 3
(2) $\times$ 7 $\times$ 5, 7 $\times$
(3) $\times$ $\emptyset$ $\times$ $\times$ $\times$
(4) $\times$ $\times$ $\times$ $\times$ $\emptyset$
(5) $\times$ $\times$ $\times$ $\times$ $\times$
(6) $\times$ $\times$ $\times$ $\times$ $\emptyset$
(7) $\times$ $\times$ $\times$ $\times$ $\times$
The cells marked with the multiplication symbol ($\times$) in a collumn $\alpha$ of a partial valuation $v$ indicates that $v(\alpha)\neq\mathbf{U}$, so we can ignore that cell. The cells marked with the empty set symbol ($\emptyset$) in a collumn $\alpha$ of a partial valuation $v$ indicates that $v(\alpha)=\mathbf{U}$ and there is no other line $w$ to validate $v$, so that line must be removed.
Table 6: Intermediate table.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(1) F U F U U
(2) F U F U T
(5) F T T F T
(7) T F T F T
Table 7: Second cycle.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(1) $\times$ 7 $\times$ 5, 7 $\emptyset$
(2) $\times$ 7 $\times$ 5, 7 $\times$
(5) $\times$ $\times$ $\times$ $\times$ $\times$
(7) $\times$ $\times$ $\times$ $\times$ $\times$
Table 8: Final table.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(2) F U F U T
(5) F T T F T
(7) T F T F T
Table 9: Validators.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(2) $\times$ 7 $\times$ 5, 7 $\times$
(5) $\times$ $\times$ $\times$ $\times$ $\times$
(7) $\times$ $\times$ $\times$ $\times$ $\times$
Note that, in the validators table, there is no occurrence of $\emptyset$. So, there is no further row to remove and we can stop the searching algorithm.
4.2.3 Soundness and completeness of the intuitionistic truth tables
Now, it will be shown that the generalized notion of truth tables for IPL proposed in Subsection 4.2.2 is sound and complete. As in the case of S4, this will be shown by proving first analyticity and co-analyticity w.r.t. level valuations in $\mathcal{M}_{IPL}$. In order to do this it is considered, adapting [4, Definition 3.12], the intermediary notion of intuitionistic partial’ level valuation in $\mathcal{R}(\mathcal{M}_{IPL})$ over $\Lambda$.
Definition 4.30 (Intuitionistic partial’ level valuation over $\Lambda$).
Let $\Lambda\in CS(\Omega)$. An intuitionistic partial valuation $\tilde{v}_{p}\in iPV(\Lambda)$ is an intuitionistic partial’ level valuation in $\mathcal{R}(\mathcal{M}_{IPL})$ over $\Lambda$ iff
  1. $\forall\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$, there exists a level valuation $w$ in $\mathcal{L}_{IPL}$ such that $w(\alpha)=\mathbf{F}$ and, $\forall\beta\in\Lambda$, $w(\beta)=\mathbf{T}$ whenever $\tilde{v}_{p}(\beta)=\mathbf{T}$.
The set of intuitionistic partial’ level valuations in $\mathcal{R}(\mathcal{M}_{IPL})$ over $\Lambda$ will be denoted by $iPLV^{\prime}(\Lambda)$.
Remark 4.31.
Recall the predicate $\mathsf{P}_{\Lambda}^{i}(v,w,\alpha)$ introduced in Remark 4.27. Clearly,
$\begin{array}[]{lll}iPLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},w,\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $w\in\mathcal{L}_{IPL}$}\big{)}\}.\end{array}$
Lemma 4.32 (Co-analyticity lemma for $\mathcal{R}(\mathcal{M}_{IPL})$).
Let $v\in\mathcal{L}_{IPL}$ and $\Lambda\in CS(\Omega)$. Then, the restriction $\tilde{v}_{p}:=v_{|\Lambda}$ of $v$ to the domain $\Lambda$ belongs to $iPLV^{\prime}(\Lambda)$.
Proof.
It is an adaptation of the proof of [4, Lemma 3.13]. Suppose, by contradiction, that $\tilde{v}_{p}\notin iPLV^{\prime}(\Lambda)$. Let $\Gamma:=\{\beta\in\Lambda\mid\tilde{v}_{p}(\beta)=\mathbf{T}\}$. Then, there exists $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$ and, for every $w\in\mathcal{L}_{IPL}$, $w(\alpha)=\mathbf{F}$ implies that $w(\beta)\neq\mathbf{T}$ for some $\beta\in\Gamma$. By contraposition, this means that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha$. But $v(\beta)=\mathbf{T}$ for every $\beta\in\Gamma$ and $v(\alpha)=\mathbf{U}$, a contradiction. From this we infer that $\tilde{v}_{p}\in iPLV^{\prime}(\Lambda)$. ∎
Remark 4.33.
Observe that, as it was done in Remark 3.22 for S4, it is possible to define an enumeration $\alpha_{1},\alpha_{2},\ldots\alpha_{m}\ldots$ (for $m\in\omega^{2}$) of $For(\Omega)$ such that $\textsf{c}(\alpha_{i})\leq\textsf{c}(\alpha_{j})$ if $i\leq j$ and, for every $i$ such that $\textsf{c}(\alpha_{i})>0$, if $\beta$ is a strict subformula of $\alpha_{i}$ then $\beta=\alpha_{j}$ for some $j<i$. Such an enumeration of $For(\Omega)$ can be defined as follows: every formula have and index in $\omega^{2}$, which is a denumerable ordinal, such that all the formulas with complexity 00 (which form a denumerable set) are placed first, with indexes in $I_{0}:=\omega$; after this, all the formulas with complexity $1$ (which form a denumerable set) are placed with an index in $I_{1}:=\omega\cdot 2\setminus\omega=\{\omega,\omega+1,\ldots\}$; in general, the formulas with complexity $n$ (which form a denumerable set) have an index in $I_{n}:=\omega\cdot(n+1)\setminus\omega\cdot n=\{\omega\cdot n,\omega\cdot n+1,\ldots\}$.
Lemma 4.34 (Analyticity lemma for $\mathcal{R}(\mathcal{M}_{IPL})$).
Let $\Lambda\in CS(\Omega)$ and $\tilde{v}_{p}\in iPLV^{\prime}(\Lambda)$. Then, there exists $v\in\mathcal{L}_{IPL}$ such that the restriction $v_{|\Lambda}$ of $v$ to the domain $\Lambda$ coincides with $\tilde{v}_{p}$.
Proof.
It is an adaptation of the proof of [4, Lemma 3.14] to $\mathcal{R}(\mathcal{M}_{IPL})$.
Suppose first that $\Lambda=For(\Omega)$. By induction on $n\in\omega$, it will be shown that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n}$ for every $n$. By Definition 4.25 of partial valuation in $\mathcal{M}_{IPL}$, $\tilde{v}_{p}\in Val_{+}(\mathcal{M}_{IPL})={\mathcal{L}}_{0}^{IPL}$. Suppose that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n}$ for a given $n\geq 0$, and let $\alpha\in For(\Omega)_{n+1}$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$. By Definition 4.30 of intuitionistic partial’ level valuation, there exists a level valuation $w$ in $\mathcal{L}_{IPL}$ (and so, $w\in\mathcal{L}^{IPL}_{n}$) such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},w,\alpha)$. Since $\Lambda=For(\Omega)$, this means that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n+1}$. From this, $\tilde{v}_{p}\in\mathcal{L}_{IPL}$ and the result clearly holds.
Suppose now that $\Lambda\neq\emptyset$ is a proper subset of $For(\Omega)$. Consider an enumeration of $For(\Omega)$ as in Remark 4.33. The valuation $v$ will be defined by induction on the complexity $n$ of the formulas. Observe that $\Lambda\cup\mathcal{P}\in CS(\Omega)$ and, for every $n\in\omega$, $\bar{\Lambda}_{n}\in CS(\Omega)$ for $\bar{\Lambda}_{n}:=\Lambda\cup For(\Omega)_{n}$. In addition, for every $m\in\omega^{2}$ and $n\geq 0$, if $\alpha_{m}$ is such that $\textsf{c}(\alpha_{m})=n+1$, then $\bar{\Lambda}_{n}^{m}\in CS(\Sigma)$, where $\bar{\Lambda}_{n}^{m}=\bar{\Lambda}_{n}\cup\{\alpha_{m}\}$. Clearly $\bar{\Lambda}_{n+1}=\cup_{k\in\omega}\bar{\Lambda}_{n}^{\omega\cdot(n+1)+k}$.
Base $n=0$: Define $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v(p)=\mathbf{F}$ if $p$ is a propositional variable which does not belong to $\Lambda$. Clearly, $v\in PLV^{\prime}(\Lambda\cup\mathcal{P})$ up to this stage.
Inductive step: Assume that $v(\alpha)$ was defined for every $\alpha\in\bar{\Lambda}_{n}$ such that $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v\in iPLV^{\prime}(\bar{\Lambda}_{n})$ up to this stage, for a given $n\geq 0$ (Induction Hypothesis, IH). Let $\alpha=\alpha_{m}$ (for $m\in I_{n+1}$) such that $\alpha\not\in\Lambda$ and $\textsf{c}(\alpha)=n+1$. Observe that $v(\beta)$ was already defined, for every strict subformula $\beta$ of $\alpha$. Since $v\in iPLV^{\prime}(\bar{\Lambda}_{n})$ then, for every $\delta\in\bar{\Lambda}_{n}$ such that $v(\delta)=\mathbf{U}$ there exists $w\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\bar{\Lambda}_{n}}^{i}(v,w,\delta)$. Such a $w$ will be denoted by $w^{\delta}$ (observe that it is possible to have more that one $w^{\delta}$ for each $\delta$, and it is possible to have $w^{\delta}=w^{\gamma}$ for $\delta\neq\gamma$). Recall that $\bar{\Lambda}_{n}^{m}=\bar{\Lambda}_{n}\cup\{\alpha_{m}\}$ is closed under subformulas. There are three cases to analyze:
(1) $\alpha=\neg\beta$ such that $\neg^{\mathcal{A}_{IPL}}v(\beta)=\{\mathbf{U},\mathbf{T}\}$, or $\alpha=\beta\to\gamma$ such that $\to^{\mathcal{A}_{IPL}}(v(\beta),v(\gamma))=\{\mathbf{U},\mathbf{T}\}$. There are three subcases to analyze:
(1.1) There is $\delta\in\bar{\Lambda}_{n}$ with $v(\delta)=\mathbf{U}$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=\mathbf{F}$. Define $v(\alpha)=\mathbf{U}$. Hence $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$ such that $w^{\alpha}=w^{\delta}$.
(1.2) There is $\delta\in\bar{\Lambda}_{n}$ with $v(\delta)=\mathbf{U}$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=\mathbf{U}$. By Lemma 4.32, $v^{\prime}=w^{\delta}_{|\bar{\Lambda}_{n}^{m}}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$, hence there exists $w^{\prime\prime}\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\bar{\Lambda}_{n}^{m}}^{i}(v^{\prime},w^{\prime\prime},\alpha)$. Define $v(\alpha)=\mathbf{U}$. Hence, $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$ such that $w^{\alpha}=w^{\prime\prime}$.
(1.3) For every $\delta\in\bar{\Lambda}_{n}$ such that $v(\delta)=\mathbf{U}$, $w^{\delta}(\alpha)=\mathbf{T}$ for every $w^{\delta}$. Define $v(\alpha)=\mathbf{T}$. Clearly, $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$.
[…]
We have shown in cases (1)-(3) how to extend the domain of $v$ to the additional formula $\alpha_{m}$, for every $\alpha_{m}$ such that $\textsf{c}(\alpha_{m})=n+1$, showing that $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$. Consider now the simultaneous extension of $v$ with the values assigned to every $\alpha_{m}$ as prescribed above. We obtain so an extension of $v$ to $\bar{\Lambda}_{n+1}=\cup_{m}\bar{\Lambda}_{n}^{m}$. By the very definition of intuitionistic partial’ level valuations it follows that $v\in iPLV^{\prime}(\bar{\Lambda}_{n+1})$, and $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$. This completes the inductive step. ∎
Proposition 4.35.
Let $\Lambda\in FCS(\Omega)$. Then $iPLV(\Lambda)=iPLV^{\prime}(\Lambda)$.
Proof.
Let $\tilde{v}_{p}\in iPLV^{\prime}(\Lambda)$. Then, $\tilde{v}_{p}\in iPV(\Lambda)$. Suppose that $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$. Then, there exists $w\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},w,\alpha)$. Let $\tilde{w}_{p}=w_{|\Lambda}$. By Lemma 4.32, $\tilde{w}_{p}\in iPLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}_{p},\alpha)$. Now, suppose that there exists $\tilde{w}^{\prime}_{p}\in iPLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}^{\prime}_{p},\alpha)$. By Lemma 4.34, there exists $v^{\prime}\in\mathcal{L}_{IPL}$ such that the restriction of $v^{\prime}$ to $\Lambda$ coincides with $\tilde{w}^{\prime}_{p}$. Hence, $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},v^{\prime},\alpha)$. This shows that
$\begin{array}[]{lll}iPLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in iPLV^{\prime}(\Lambda)$}\big{)}\}.\end{array}$ From this, being $iPV(\Lambda)$ finite, it follows that $iPLV(\Lambda)=iPLV^{\prime}(\Lambda)$. ∎
Definition 4.36.
Let $\Gamma\cup\{\varphi\}$ be a finite subset of $For(\Omega)$. Let $\Lambda$ be the (finite) set of subformulas of $\Gamma\cup\{\varphi\}$. We say that $\varphi$ is a consequence of $\Gamma$ w.r.t. intuitionistic truth-tables, denoted by $\Gamma\models_{iPLV}\varphi$, if, for every intuitionistic partial level valuation $\tilde{v}_{p}\in iPLV(\Lambda)$, if $\tilde{v}_{p}(\beta)=\mathbf{T}$ for every $\beta\in\Gamma$ then $\tilde{v}_{p}(\varphi)=\mathbf{T}$. In particular, $\emptyset\models_{iPLV}\varphi$, denoted simply by $\models_{iPLV}\varphi$, if $\tilde{v}_{p}(\varphi)=\mathbf{T}$ for every $\tilde{v}_{p}\in iPLV(\Lambda)$, where $\Lambda$ is the set of subformulas of $\varphi$.
Theorem 4.37 (Soundness and completeness of intuitionistic truth-tables w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$).
Let $\Gamma\cup\{\varphi\}$ be a finite subset of $For(\Omega)$. Then: $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$ if, and only if, $\Gamma\models_{iPLV}\varphi$.
Proof.
Let $\Lambda$ be the set of subformulas of $\Gamma\cup\{\varphi\}$. Observe that $\Lambda\in FCS(\Omega)$. Suppose that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$. Let $\tilde{v}_{p}\in iPLV(\Lambda)$ such that $\tilde{v}_{p}(\beta)=\mathbf{T}$ for every $\beta\in\Gamma$. By Proposition 4.35, $\tilde{v}_{p}\in iPLV^{\prime}(\Lambda)$. By Lemma 4.34, there exists $v\in\mathcal{L}_{IPL}$ such that the restriction $v_{|\Lambda}$ of $v$ to $\Lambda$ coincides with $\tilde{v}_{p}$. Then, $v(\beta)=\mathbf{T}$ for every $\beta\in\Gamma$ and so, by hypothesis, $v(\varphi)=\mathbf{T}$. That is, $\tilde{v}_{p}(\varphi)=\mathbf{T}$. This shows that $\Gamma\models_{iPLV}\varphi$.
Conversely, suppose that $\Gamma\models_{iPLV}\varphi$. Let $v\in\mathcal{L}_{IPL}$ such that $v(\beta)=\mathbf{T}$ for every $\beta\in\Gamma$. By Lemma 4.32, the restriction $\tilde{v}_{p}:=v_{|\Lambda}$ of $v$ to $\Lambda$ belongs to $iPLV^{\prime}(\Lambda)$. By Proposition 4.35, $\tilde{v}_{p}\in iPLV(\Lambda)$ such that $\tilde{v}_{p}(\beta)=\mathbf{T}$ for every $\beta\in\Gamma$. By hypothesis, $\tilde{v}_{p}(\varphi)=\mathbf{T}$. Then, $v(\varphi)=\mathbf{T}$ and so $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$. ∎
Proposition 4.38.
Let $\Gamma\cup\{\alpha,\beta\}\subseteq For(\Omega)$. If $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$ then $\Gamma,\alpha\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$.
Proof.
Assume that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$, and let $v\in\mathcal{L}_{IPL}$ such that $v(\gamma)=v(\alpha)=\mathbf{T}$ for every $\gamma\in\Gamma$. By hypothesis, $v(\alpha\to\beta)=\mathbf{T}$. By definition of $\to^{\mathcal{A}_{IPL}}$, and given that $v\in Val(\mathcal{M}_{IPL})$, it follows from this that $v(\beta)=\mathbf{T}$. This shows that $\Gamma,\alpha\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$. ∎
Theorem 4.39 (Soundness of $\mathcal{H}_{IPL}$ w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$).
For every $\Gamma$ and $\varphi$: $\Gamma\vdash_{IPL}\varphi$ implies that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$.
Proof.
Since the consequence relation $\mathcal{H}_{IPL}$ of the Hilbert calculus for IPL is finitary and satisfies the deduction metatheorem, and taking into account Proposition 4.38, it is enough to prove the result for $\Gamma=\emptyset$. We prove first the following two facts:
Fact 1: If $\varphi$ is an instance of an axiom in $\mathcal{H}_{IPL}$ then $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$. Indeed, let $\varphi$ be an instance of an axiom in $\mathcal{H}_{IPL}$, and let $\Lambda$ be the set of subformulas of $\varphi$. Then, $\Lambda\in FCS(\Omega)$. It is routine to check that the truth-table for $\varphi$ over $\Lambda$ validates $\varphi$ (see Appendix A). That is: for every intuitionistic partial level valuation $\tilde{v}_{p}\in iPLV(\Lambda)$, $\tilde{v}_{p}(\varphi)=\mathbf{T}$. By Theorem 4.37 it follows that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$.
Fact 2: If $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha$ and $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$ then $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$. Indeed, assume that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha$ and $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$. If $v\in\mathcal{L}_{IPL}$ then $v(\alpha)=v(\alpha\to\beta)=\mathbf{T}$. By definition of $\to^{\mathcal{A}_{IPL}}$, and by the fact that $v\in Val(\mathcal{M}_{IPL})$, it follows that $v(\beta)=\mathbf{T}$. This means that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$.
Now, suppose that $\vdash_{IPL}\varphi$. By induction on the length of a derivation of $\varphi$ in $\mathcal{H}_{IPL}$, and by Fact 1 and Fact 2 it is immediate to prove that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$. ∎
Corollary 4.39.1 (Finitariness for $\mathcal{R}(\mathcal{M}_{IPL})$).
Let $\Gamma\cup\{\varphi\}\subseteq For(\Omega)$. If $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$ then $\Gamma_{0}\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$ for some finite subset $\Gamma_{0}$ of $\Gamma$.
Proof.
It follows from Theorems 4.23, 4.39 and the fact that $\mathcal{H}_{IPL}$ is finitary. ∎
Corollary 4.39.2 (Deduction metatheorem for $\mathcal{R}(\mathcal{M}_{IPL})$).
Let $\Gamma\cup\{\alpha,\beta\}\subseteq For(\Omega)$. If $\Gamma,\alpha\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$ then $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$.
Proof.
It follows from Theorems 4.23, 4.39 and the fact that $\mathcal{H}_{IPL}$ satisfies the deduction metatheorem. ∎
Corollary 4.39.3 (Deduction-detachment metatheorem for intuitionistic truth-tables).
Let $\Gamma\cup\{\alpha,\beta\}$ be a finite subset of $For(\Omega)$. Then, $\Gamma,\alpha\models_{iPLV}\beta$ if, and only if, $\Gamma\models_{iPLV}\alpha\to\beta$.
Proof.
It follows from Theorem 4.37, Proposition 4.38 and Corollary 4.39.2. ∎
It is worth noting that the latter result is useful to reduce the complexity of the truth-tables: instead of generating the (reduced) truth table for a formula $\delta=\beta_{1}\to(\beta_{2}\to(\ldots\to(\beta_{n}\to\varphi)\ldots))$ to be tested as a tautology, it suffices generating the (reduced) truth-table for $\{\beta_{1},\ldots,\beta_{n},\varphi\}$ and then determine if, in every row such that $\beta_{1},\ldots,\beta_{n}$ receive simultaneously the value $\mathbf{T}$, $\varphi$ also receives the value $\mathbf{T}$. This reduces the number of rows required to test $\delta$.
Corollary 4.39.4 (Soundness and completeness of $\mathcal{H}_{IPL}$ w.r.t. intuitionistic truth-tables).
Let $\Gamma\cup\{\varphi\}$ be a finite subset of $For(\Omega)$. Then: $\Gamma\vdash_{IPL}\varphi$ if, and only if, $\Gamma\models_{iPLV}\varphi$.
Proof.
It follows from Theorems 4.39, 4.23 and 4.37. ∎
The table is never empty.
Lemma 4.40.
Let $\Delta\in CS(\Sigma)$ and $\beta\in For(\Sigma)$. If $v_{p}:\Delta\to\mathcal{V}$ is $iPLV$ and $sub(\beta)\subseteq\Delta$, then $v_{p}$ has at least one extension $v^{\prime}_{p}:\Delta\cup\{\beta\}\to\mathcal{V}$ such that $v^{\prime}_{p}$ is $iPLV$.
Proof.
By hypothesis, $v_{p}$ is iPLV. Then, by analiticity of $\mathcal{R}(\mathcal{M}_{IPL})$, there is a level valuation $v:For(\Sigma)\to\mathcal{V}$ such that $v_{p}=v_{\mid\Delta}$, i.e, for every $\alpha\in\Delta$, $v(\alpha)=v_{p}(\alpha)$. More then that, $v$ is defined for $\beta$. Now, given that $sub(\beta)\subseteq\Delta$ and $\Delta\in CS(\Sigma)$, then $\Delta\cup\{\beta\}\in CS(\Sigma)$ and, by co-analiticity of $\mathcal{R}(\mathcal{M}_{IPL})$, $v^{\prime}_{p}=v_{\mid\Delta\cup\{\beta\}}$ is iPLV. ∎
5 Complexity
Given a non-deterministic matrix $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ defined over a language $\mathcal{L}$ and a formula $\alpha\in\mathcal{L}$, we can estimate the total number of rows of the final truth table for $\alpha$ via the following upper bound.
Let $sub(\alpha)\subseteq\mathcal{L}$ be the set of all sub-formulas of $\alpha$, $atoms(\alpha)\subseteq sub(\alpha)$ the set of sub-formulas whose complexity is 00, and $\kappa\geq 1$ the branching factor. Then, the maximum number of rows of the truth table for $\alpha$ in $\mathcal{M}$ is given by
Definition 5.1.
$ub_{\mathcal{M}}(\alpha)=\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}\times|\mathcal{V}|^{|atoms(\alpha)|}$ (1)
In 5.1, we assume that the branching factor is the maximum degree of non-determinism. More formally,
Definition 5.2.
Let $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ be some nmatrix. The branching factor of $\mathcal{M}$ is the highest $\kappa$ such that, for every n-ary ~$\tilde{\circ}\in\mathcal{O}$, every formula $\alpha_{1},\ldots,\alpha_{n}$ and every valuation $v\in Val(\mathcal{M})$, |~$|\tilde{\circ}(v(\alpha_{1}),\ldots,v(\alpha_{n}))|\leq\kappa$.
For example, in the case of $\mathcal{M}_{IPL}$, we have that disjunction and conjunction have branching factor $1$, while implication and negation have branching factor $2$. Therefore, in this case, $\kappa=2$.
Remark 5.3.
If we consider the valuations in the nmatrix as branches in trees, we can say that $|\mathcal{V}|^{|atoms(\alpha)|}$ gives the number of trees, while $\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}$ gives the (maximum) number of branches of each tree.
We can also obtain a lower bound by assuming $\kappa=1$, which is the branching factor of deterministic matrices. This gives us
Definition 5.4.
$lb_{\mathcal{M}}(\alpha)=|\mathcal{V}|^{|atoms(\alpha)|}$ (2)
From (5.1) and (5.4), it follows that
$ub_{\mathcal{M}}(\alpha)=\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}\times lb_{\mathcal{M}}(\alpha)$ (3)
$p$ $p$
$\mathcal{V}_{1}$ $\mathcal{V}_{1}$
$\vdots$ $\vdots$
$\mathcal{V}_{n}$ $\mathcal{V}_{n}$
Table 10: Base case.
Now we shall prove that these are faithful lower and upper bounds for the number of rows.
Theorem 5.5.
Let $\mathcal{M}$ be a nmatrix and $rows(\alpha)$ the set of every row of the truth table for some formula $\alpha$. Then
$lb_{\mathcal{M}}(\alpha)\leq|rows(\alpha)|\leq ub_{\mathcal{M}}(\alpha)$
Proof.
Let $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ be the nmatrix for the language of $\alpha$, where $\mathcal{V}=\{\mathcal{V}_{1},\ldots,\mathcal{V}_{n}\}$. Induction on the complexity of $\alpha$.
Base case If $\alpha=p$, the truth table is depicted in Table 10, which has $|\mathcal{V}|=n$ rows. Now, $|sub(\alpha)|=|atoms(\alpha)|=1$. For this reason, $lb_{\mathcal{M}}(\alpha)=|\mathcal{V}|^{1}=|\mathcal{V}|$ and $ub(\alpha)=\kappa^{0}\times|\mathcal{V}|^{1}=|\mathcal{V}|$. Therefore, $lb_{\mathcal{M}}(p)=|rows(p)|=ub_{\mathcal{M}}(p)$.
IH If $c(\alpha)<c(\alpha^{\prime})$, then the theorem holds for $\alpha$.
$\alpha=\neg\beta$
Let $rows(\beta)=\{v_{0},\ldots,v_{n}\}$ be the list of rows of the nmatrix for $\beta$. Because $atoms(\beta)=atoms(\neg\beta)$, we conclude that $lb_{\mathcal{M}}(\neg\beta)=lb_{\mathcal{M}}(\beta)$. Now, observe that $|sub(\neg\beta)|=|sub(\beta)|+1$. From this, we get $ub_{\mathcal{M}}(\neg\beta)=\kappa^{|sub(\beta)|-|atoms(\neg\beta)|+1}\times lb_{\mathcal{M}}(\neg\beta)=ub_{\mathcal{M}}(\beta)\times\kappa$. By the definition of branching factor, for every $v_{n}\in rows(\beta)$, |¬~$|\tilde{\neg}(v_{n}(\beta))|\leq\kappa$. Therefore, $|rows(\neg\beta)|\leq|rows(\beta)|\times\kappa$. Now, by (IH), $lb_{\mathcal{M}}(\beta)\leq|rows(\beta)|\leq ub_{\mathcal{M}}(\beta)$. Therefore, $lb_{\mathcal{M}}(\neg\beta)\leq|rows(\neg\beta)|\leq ub_{\mathcal{M}}(\neg\beta)$.
[…]
Let $rows(\beta)=\{v_{0},\ldots,v_{n}\}$ be the nmatrix of $\beta$ and let $rows(\gamma)=\{w_{0},\ldots,w_{n}\}$ be the nmatrix of $\gamma$. By (IH), we get the following
$lb_{\mathcal{M}}(\beta)\leq|rows(\beta)|\leq ub_{\mathcal{M}}(\gamma)$ (4)
$lb_{\mathcal{M}}(\gamma)\leq|rows(\gamma)|\leq ub_{\mathcal{M}}(\gamma)$ (5)
Using 4 and 5 and some laws of algebra, we get
$lb_{\mathcal{M}}(\beta)\times lb_{\mathcal{M}}(\gamma)\leq rows(\beta)\times rows(\gamma)\leq ub_{\mathcal{M}}(\beta)\times ub_{\mathcal{M}}(\gamma)$
Now, to construct the truth table for $\alpha$, we combine $rows(\beta)$ with $rows(\gamma)$. In the worst case, $sub(\beta)\cap sub(\gamma)=\emptyset$. In this case, we append, to each row of $rows(\beta)$, a row of $rows(\gamma)$. Then, each row creates $\kappa-1$ fresh rows. Therefore,
$|rows(\alpha)|\leq|rows(\beta)|\times|rows(\gamma)|\times\kappa$ (6)
Moreover,
$|sub(\alpha)|$ $\leq|sub(\beta)|+|sub(\gamma)|+1$ (7)
$|atoms(\alpha)|$ $\leq|atoms(\beta)|+|atoms(\gamma)|$ (8)
Therefore,
$|\mathcal{V}|^{|atoms(\alpha)|}$ $\leq|\mathcal{V}|^{|atoms(\beta)|+|atoms(\gamma)|}$
$=|\mathcal{V}|^{|atoms(\beta)|}\times|\mathcal{V}|^{|atoms(\gamma)|}$
$=lb_{\mathcal{M}}(\beta)\times lb_{\mathcal{M}}(\gamma)$
and
$\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}\times|\mathcal{V}|^{|atoms(\alpha)|}$
$\leq\kappa^{(|sub(\beta)|+|sub(\gamma)|+1)-(|atoms(\beta)|+|atoms(\gamma)|)}\times|\mathcal{V}|^{|atoms(\alpha)|}$
$=\kappa^{(|sub(\beta)|-|atoms(\beta)|)+(|sub(\gamma)|-|atoms(\gamma)|)+1}\times\left(|\mathcal{V}|^{|atoms(\beta)|}\times|\mathcal{V}|^{|atoms(\gamma)|}\right)$
$=\left(\kappa^{(|sub(\beta)|-|atoms(\beta)|)}\times|\mathcal{V}|^{|atoms(\beta)|}\right)\times\left(\kappa^{(|sub(\gamma)|-|atoms(\gamma)|)}\times|\mathcal{V}|^{|atoms(\gamma)|}\right)\times\kappa$
$=ub_{\mathcal{M}}(\beta)\times ub_{\mathcal{M}}(\gamma)\times\kappa$
Hence,
$lb_{\mathcal{M}}(\beta\#\gamma)\leq|rows(\beta\#\gamma)|\leq ub_{\mathcal{M}}(\beta\#\gamma)$
5.1 Algorithms
Given a formula $\varphi$, to decide if $\varphi$ is valid we have to proceed in two steps: (1) construct the initial matrix by applying the definitions of the multi-operations, and, then, (2) refine the matrix by removing every spurious row. In the previous section, we have proposed a way to measure the space complexity of (1). Now, we will find the time complexity of (2).
5.1.1 Pseudo-codes
In what follows, given an ordered pair $\langle a;b\rangle$,
$\pi_{1}(\langle a;b\rangle)=a$
and
$\pi_{2}(\langle a;b\rangle)=b$
Moreover, $\mathcal{L}$ denotes the set of every intuitionistic formula and $\mathcal{V}=\{\textbf{F},\textbf{U},\textbf{T}\}$.
Algorithm 1 Compatibility procedure. This algorithm receives two valuations, $v$ and $w$, and check if, for every evaluated formula $\alpha$, if $v(\alpha)=\textbf{T}$ then $w(\alpha)=\textbf{T}$ ($v$ is compatible with $w$). We assume that every valuation list is equally sorted.
procedure isCompatible($v,w:list(\mathcal{L}\times\mathcal{V})$)
     $i\leftarrow 0$
     while $i<length(v)$ do
         $a\leftarrow\pi_{2}(v[i])$triangleright Get the value in $v$ on index $i$.
         if $a=$ T then
              $b\leftarrow\pi_{2}(w[i])$triangleright Get the value in $w$ on the same index.
              if $b=$ T then
                  continue
              else
                  return $false$
              end if
         else
              continue
         end if
         $i\leftarrow i+1$
     end while
     return $true$
end procedure
Algorithm 2 Search algorithm. This algorithm receives a formula $A\in\mathcal{L}$, a (partial) valuation $v:\mathcal{L}\rightarrow\mathcal{V}$ and a matrix. It runs over the entire matrix looking for a valuation $w$ that validates $v$, that is, a $w$ such that $v$ is compatible with and such that $w(A)=\textbf{F}$.
procedure validate($A:\mathcal{L},v:list(\mathcal{L}\times\mathcal{V}),matrix:list(list(\mathcal{L}\times\mathcal{V}))$)
     $i\leftarrow 0$
     while $i<length(matrix)$ do
         $w\leftarrow matrix[i]$triangleright Get a valuation function.
         $j\leftarrow 0$
         while $j<length(w)$ do
              $A^{\prime}\leftarrow\pi_{1}(w[j])$triangleright Get a valuated formula.
              if $A^{\prime}=A$ then
                  $wA\leftarrow\pi_{2}(w[j])$triangleright Get $w(A)$.
                  if $wA=$ F then
                       $compatible\leftarrow isCompatible(v,w)$
                       if $compatible$ then
                           return $true$triangleright In this case, $w$ validates $v$.
                       else
                           continue
                       end if
                  else
                       continue
                  end if
              else
                  continue
              end if
              $j\leftarrow j+1$
         end while
         $i\leftarrow i+1$
     end while
     return $false$triangleright In this case, no $w$ validates $v$.
end procedure
Algorithm 3 Refinement algorithm. In this pseudo-code, we assume that $remove$ is a function that receives a matrix and a row, returning the matrix minus the row.
procedure refineMatrix($matrix:list(list(\mathcal{L}\times\mathcal{V}))$)
     $i\leftarrow 0$
     $newmatrix\leftarrow matrix$
     while $i<length(matrix)$ do
         $v\leftarrow matrix[i]$triangleright Get a valuation function.
         $j\leftarrow 0$
         while $j<length(v)$ do
              $A\leftarrow\pi_{1}(v[j])$triangleright Get a valuated formula.
              $vA\leftarrow\pi_{2}(v[j])$triangleright Get $v(A)$.
              if $vA=$ U then
                  $isValid\leftarrow validate(A,v,newmatrix)$
                  if $isValid$ then
                       continue $\triangleright$ In this case, we keep the matrix unaltered.
                  else
                       $newmatrix\leftarrow remove(v,newmatrix)$
                       break
                  end if
              else
                  continue
              end if
              $j\leftarrow j+1$
         end while
         $i\leftarrow i+1$
     end while
     return $newmatrix$triangleright Return the refined matrix.
end procedure
Algorithm 4 In the best case, the refinement algorithm does not remove any row. However, if it does drag at least one row, we have to rerun it, given that it could have been the case that the removed row was used to validate some other rows. For this reason, in the worst case, the refinement algorithm removes every row and one row at once. This auxiliar procedure should address this possibility.
procedure refineAux($matrix:list(list(\mathcal{L}\times\mathcal{V}))$)
     $i\leftarrow 0$
     $nrows\leftarrow length(matrix)$
     while $i<nrows$ do
         $matrix\leftarrow refineMatrix(matrix)$
         $i\leftarrow i+1$
     end while
     return $matrix$
end procedure
Theorem 5.6.
Given as input a matrix with $n$ rows and $m$ columns, the time complexity of refineAux is $O(n^{3}*m^{3})$.
Proof.
Given a list $v$ with length equal to $m$, in the worst case, the loop in $isCompatible$ runs $m$ times. Therefore, the complexity of this function is $O(m)$. Now, while, in the worst case, the external loop in $validate$ runs $n$ times, where $n$ is the number of rows in the matrix, the internal loop runs $m$ times. Because this functions calls $n*m$ times the function $isCompatible$, then the complexity of $validate$ is $O(n*m)*O(m)=O(n*m^{2})$. In the case of $refineMatrix$, the external loop runs $n$ times and the internal loop runs $m$ times. Therefore, it calls $validate$ $n*m$ times, which gives us $O(n*m)*O(n*m^{2})=O(n^{2}*m^{3})$. Finally, the function $refineAux$ calls $refineMatrix$ $n$ times. Therefore, the complexity of $refineAux$ is $O(n)*O(n^{2}*m^{3})=O(n^{3}*m^{3})$. ∎
5.2 Complexity of $\mathcal{R}(\mathcal{M}_{IPL})$
In the valuation of a formula in $\mathcal{M}_{IPL}$, we consider only $\mathbf{F}$ and $\mathbf{T}$ in the atomic cases. Given that, in this case, $\kappa=2$, we can simplify the upper bound to:
$ub_{\mathcal{M}_{IPL}}(\alpha)$ $=2^{(|sub(\alpha)|-|atoms(\alpha)|)}\times 2^{|atoms(\alpha)|}$ (9)
$=2^{|sub(\alpha)|}$ (10)
Now, given a formula $\varphi\in For(\Omega)$, assume that $|sub(\varphi)|=m$ and that $ub_{\mathcal{M}_{IPL}}(\varphi)=n$. Then
$n^{3}$ $=(2^{m})^{3}=8^{m}$
Therefore, given a formula $\varphi$, the time complexity to compute the truth table (in the worst case) in $\mathcal{R}(\mathcal{M}_{IPL})$ is given by $O(8^{m}*m^{3})$, which is in EXPTIME.
6 Related work and concluding remarks
In this paper, we proposed a new decision procedure for intuitionistic propositional logic IPL based on a 3-valued restricted Nmatrix semantics. Starting from the nice decision procedure for S4 obtained by Grätz in [4] based on a 3-valued RNmatrix, and taking advantage of the conservative translation of Gödel from IPL to S4, it is immediate to obtain an algorithm to decide validity in IPL. By means of an intermediate translation mapping from the language of IPL into the language of S4 we define a 3-valued Nmatrix which is the basis for an RNmatrix semantics for IPL. Moreover, by adapting Grätz’s algorithm for S4, a very simple algorithm for deleting spurious rows in the non-deterministic truth-table generated by the Nmatrix of IPL for a given formula is also provided. The sound and completeness proofs for the RNmatrix and for the obtained truth-tables algorithm for IPL, are presented with full technical details. These proofs were adapted from new ones for S4 we obtained in Subsection 3.2. On the one hand, we reduced the original Nmatrix for S4 presented by Grätz. On the other, we expanded it by adding the truth-tables for conjunction and disjunction. This required the extension and modification of the proofs originally given in [4]. It is worth noting that the Nmatrix for IPL we introduced in Definition 4.10 has a clear and intuitive interpretation in terms of constructibility, by observing the definition of the canonical valuation associated to a $\varphi$-saturated set in IPL given in Definition 4.19, as well as the observations made in Remark 4.20.
The idea of using finite non-deterministic truth-tables with restricted valuations for characterizing intuitionistic logic IPL is not new. Andréa Loparic already obtained in 1977, announced in [7] and finally published in [8], a semantical characterization of Johannson-Kolmogorov minimal logic and intuitionistic propositional logic IPL in terms of a $2$-valued, non truth-functional valuation semantics which induces a decision procedure for these logics. Similar to Grätz method (and to the adaptation to IPL proposed here), Loparic’s approach is based on a criterion for distinguish between that she calls semivaluations and valuations, the latter being the useful ones. This would correspond to level valuations, since they are exactly the semivaluations associate to $\varphi$-saturated sets, and the ones which characterizes the logic. The associated algorithm, as in the cases of S4 and IPL presented in this paper, generates a non-deterministic truth-table for a given formula in which, by using some formal criteria, some rows are deleted, and so the formula is tautology of the given logic if and only if it gets the value 1 in all the undeleted rows.
More recently, Alejandro Solares-Rojas proposed a 3-valued Nmatrix semantics for IPL based on intuitionistic Kripke models and the novel notion of depth-bounded approximations (see [11, Chapter 5]). Within this semantic approach to IPL, a 3-valued Nmatrix – whose truth-values are intuitionistic truth (1), falsity (0) and indeterminacy ($u$) – gives the meaning of the connectives in terms of information that a given agent holds. The valuations over the Nmatrix intend to model actual information states, which are ordered by a certain refinement relation. Such refinement relation aims to model the way knowledge of the agent evolves when new information, provided by reliable external sources, is acquired. The semantics for IPL is given by intuitionistic Kripke frames expanded by a family of valuations indexed by the worlds and where the refinement order corresponds to the accessibility relation of the frame. This interesting 3-valued non-deterministic semantics for IPL departs from a different perspective than ours. Besides the 3-valued Nmatrices proposed in [11, Chapter 5] and the one we give in Definition 4.10 are quite different, it is not immediate whether the semantics proposed by Solares-Rojas could be seen as given by an RNmatrix. This is the subject of future investigation.
It is interesting to notice that, by considering finite and decidable RNmatrix semantics (as the ones for S4 and IPL analyzed here) as a generalization of the standard notion of finite-valued (deterministic) truth-tables, the results presented here for S4 and IPL are a way to overcome the limiting results obtained by Gödel and Dugundji for IPL and S4, respectively, mentioned in Section 1.
As a topic of future research we intend to explore the meaning of the partial valuations corresponding to the rows of the non-deterministic truth-tables in IPL and S4. In particular, we plan to investigate their possible connection with the mini canonical models considered by Hughes and Cresswell in [5, Chapter 8] to prove the finite model property for some modal systems, including S4.
Finally, the extension of this methodology to other logics such as S5 and Prawitz’s ecumenical systems (see [10]) will also be investigated.
Acknowledgments: The first and second author thanks the support of the São Paulo Research Foundation (FAPESP, Brazil) through the PhD scholarship grant #21/01025-3 and the Thematic Project RatioLog #20/16353-3. The second author also acknowledges support from the National Council for Scientific and Technological Development (CNPq, Brazil), through the individual research grant #309830/2023-0.
Appendix A
The following truth-tables show that each axiom of the Hilbert calculus $\mathcal{H}_{IPL}$ for IPL is valid in the intuitionistic truth-tables. These tables were automatically generated by a Coq module for RNmatrices written by R. Leme, see
$q$ $p$ $q\to p$ $p\to(q\to p)$
F F U T
F F T T
F T T T
T F F T
T T T T
$p$ $q$ $p\land q$ $q\to p\land q$ $p\to(q\to p\land q)$
F F F U T
F F F T T
F T F F T
T F F T T
T T T T T
$q$ $p$ $r$ $q\to r$ $p\to q$ $p\to r$ $p\to(q\to r)$ $(p\to q)\to(p\to r)$ $(p\to(q\to r))\to((p\to q)\to(p\to r))$
F F F U U U U U T
F F F U U U T T T
F F F U U T T T T
F F F U T U U F T
F F F U T T T T T
F F F T U U T T T
F F F T U T T T T
F F F T T T T T T
F F T T U T T T T
F F T T T T T T T
F T F U F F F U T
F T F T F F T T T
F T T T F T T T T
T F F F T U U F T
T F F F T T T T T
T F T T T T T T T
T T F F T F F F T
T T T T T T T T T
$q$ $p$ $p\land q$ $p\land q\to p$ $p\land q\to q$
F F F T T
F T F T T
T F F T T
T T T T T
$p$ $q$ $p\lor q$ $p\to p\lor q$ $q\to p\lor q$
F F F T T
F T T T T
T F T T T
T T T T T
$p$ $q$ $r$ $p\to r$ $q\to r$ $p\lor q$ $p\lor q\to r$ $(q\to r)\to(p\lor q\to r)$ $(p\to r)\to((q\to r)\to((p\lor q)\to r))$
F F F U U F U U T
F F F U U F U T T
F F F U T F U F T
F F F T U F U T T
F F F T T F T T T
F F T T T F T T T
F T F U F T F T T
F T F T F T F T T
F T T T T T T T T
T F F F U T F U T
T F F F U T F T T
T F F F T T F F T
T F T T T T T T T
T T F F F T F T T
T T T T T T T T T
$p$ $q$ $q\to p$$\neg p$$\neg q$ $q\to\neg p$ $(q\to\neg p)\to\neg q$ $(q\to p)\to((q\to\neg p)\to\neg q)$
F F U U U U U T
F F U U U U T T
F F U U U T F T
F F U T U T F T
F F T U U U T T
F F T U T T T T
F F T T T T T T
F T F U F F U T
F T F U F F T T
F T F T F T F T
T F T F U U T T
T F T F T T T T
T T T F F F T T
$p$ $q$$\neg p$$\neg p\to q$ $p\to(\neg p\to q)$
F F U U T
F F U T T
F F T F T
F T U T T
F T T T T
T F F T T
T T F T T

### v4

@@
-Intuitionism with Truth Tables: A Decision Procedure for IPL Based on RNmatrices
+A new decision method for Intuitionistic Logic by 3-valued non-deterministic truth-tables (pre-print version)
 Abstract
-This paper provides a new decision procedure for intuitionistic propositional logic (IPL) based on a 3-valued restricted non-deterministic matrix and proves its soundness and completeness.
+Kurt Gödel proved that it is not possible to characterize Intuitionistic Propositional Logic ($\mathbf{IPL}$) by means of finite and deterministic truth-tables.
+After extending the same result with respect to non-deterministic matrices, we provide a semantical characterization of $\mathbf{IPL}$ by means of a $3$-valued non-deterministic matrix with a restricted set of valuations.
+This structure allows to define an algorithm to delete unsound rows from the non-deterministic truth-tables generated for each formula, which constitutes a new and very simple decision procedure for $\mathbf{IPL}$.
+This method can be seen as truth-tables in a broader sense, and a way to overcome Gödel’s limiting result.
+MSC class: 03B20; 03B45; 03B25; 03B35; 68V15
 1 Introduction
-A famous result, proved first by K.
-Gödel in $1932$, says that it is impossible to characterize intuitionistic propositional logic (IPL) via a finite-valued (deterministic) matrix [1].
-In other words, Gödel proved that there is no effective decision procedure based on finite-valued ordinary truth-tables for IPL.
-Later, in $1940$, a maybe less famous but equally important result was proved by Dugundji, who adapted Gödel’s argument to establish that the same is also true about the family of modal logics between S1 and S5 [1].
-To overcome Dugundji’s result, J.
-Kearns [6] introduced a $4$-valued Nmatrix (non-deterministic matrix) for modal logics KT, S4, and S5 and restricted the set of valuations to a subset called “level valuations”.
-He proved that this restricted Nmatrix (hereafter called RNmatrix [3]) constitutes sound and complete semantics for these modal logics.
-Unfortunately, in principle, Kearns’s level valuations fail to provide an effective decision procedure for any of the modal logics.
-More than that, until recently, it was unknown whether obtaining such a decision procedure from level valuations would be possible.
+In 1932, Kurt Gödel established the uncharacterizability of intuitionistic propositional logic (IPL) with respect to finite-valued deterministic matrices [10].
+He demonstrated that any finite deterministic matrix, strongly enough to validate every intuitionistic tautology, will also validate at least one intuitionistically invalid proposition.
+Later, in $1940$, James Dugundji proved a similarly significant result by adapting Gödel’s argument to show that the same holds true for modal logic S5 and all of its subsystems [8].
+These limitations led to new interpretations for modal and intuitionistic systems, such as Saul Kripke’s relational semantics, where the notion of possible world plays a central role.
+The tradition of using possible worlds to explain modalities dates back to Leibniz and is at the core of philosophical debates: Are possible worlds real or just helpful fiction?
+John Kearns challenged this tradition by defending that they are neither one nor the other [19, p.
+86].
+To support his position, he proposed, in $1981$, a four-valued semantics for the modal systems KT, S4, and S5 where the truth-value of a formula depends solely on the truth assignments of its subformulas.
+In Kearns’ approach, the meaning of certain connectives is not entirely determined by their matrices, as some inputs can yield more than one output.
+This type of semantics was later formalized under the notion of non-deterministic matrices (Nmatrices, in short) [2].
+To address this indeterminacy, instead of possible worlds, Kearns introduced a method for restricting the set of valuations, which was later referred to more generally as level valuations (see Section 4).
+In this approach, not all valuations in the Nmatrix are permitted.
+Instead, only a subset that meets specific conditions to ensure the validity of the necessitation rule is allowed; this subset is precisely the set of level valuations.
+Nmatrices with a restricted set of valuations lead to the notion of Restricted non-deterministic matrices or Restricted Nmatrices (RNmatrices, in short), central in the present paper.
+Unlike Kripke semantics, however, Kearns’s level valuations do not draw a straightforward decision procedure for those logics.
+Until recently, it was unclear whether obtaining such a decision procedure from level valuations would be possible.
 (…) the interesting problem of determining whether or not a decision procedure can be obtained by means of level valuations remains open. [6]
-L.
-Grätz recently solved this problem for some modal systems [4].
-He refined Kearn’s original RNmatrix to obtain a decidable $3$-valued RNmatrix for modal logics KT and S4.
-He noticed that, with an appropriate notion of partial valuation for level semantics (i.e., preserving analyticity and co-analyticity), the truth tables could be obtained by removing any spurious row eventually produced in the valuation process.
-In this paper, we will try to overcome the original problem posed by Gödel on the characterizability of IPL via finite-valued matrix.
-We will propose a decidable 3-valued RNmatrix for IPL combining Grätz’s RNmatrix for S4 and Gödel-McKinsey-Tarski’s box translation.
-2 Nmatrices and RNmatrices
-Along this paper, all the propositional languages will be defined over a given denumerable set $\mathcal{P}=\{p_{0},p_{1},p_{2},\ldots\}$ of propositional variables.
-Definition 2.1 (Nmatrix).
-Let $\Theta$ be a signature and $For(\Theta)$ the propositional language over signature $\Theta$ generated by the set $\mathcal{P}$ of propositional variables.
-An Nmatrix for $\mathcal{L}$ over $\Theta$ is a tuple $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ where
+In $2021$, Lukas Grätz solved this problem for some modal systems (see [12]).
+He built upon Kearns’s original RNmatrix $3$-valued RNmatrices for KT and S4 with two decidable notions of partial level valuation, one for KT and other for S4 and show that both are analytic and co-analytic regarding the level valuations of the RNmatrix for the respective modal systems.
+To obtain the truth-table, one discards (by means of a simple algorithm) any valuation that does not fit the corresponding definition.
+Ultimately, the remaining partial valuations suffice to decide the formula’s validity.
+Analyticity and co-analyticity ensure that partial level valuation semantics (which constitute the rows of the truth-tables) coincide with the corresponding RNmatrix semantics, which characterizes each logic.
+Constructing non-deterministic (i.e., row-branching) truth-tables in a finite-valued Nmatrix can be seen as a generalization of the standard (deterministic) truth-tables method for deciding validity in a given logic.
+Indeed, non-deterministic matrices were formally introduced by A.
+Avron and I.
+Lev in [2] with the aim of overcoming Gödel-Dugundji’s uncharacterizability results for some logics.
+The main purpose of Nmatrices was to obtain a decision procedure, analogous to finite-valued truth-tables, for logics that do not admit such a characterization.
+It is worth noting that Nmatrices were already considered in the literature before these authors.
+For instance, N.
+Rescher already defined a non-deterministic matrix semantics in 1962, under the name of quasi-truth-functional systems (see [28]).
+Afterwards, and among other authors, this notion was independently proposed by Y.
+Ivlev for studying non-normal modal logics, starting in the 1970s (see for instance [14, 15, 16, 17]), as well as J.
+Kearns (with restricted valuations).
+In spite of its greater expressive power with respect to standard (deterministic) matrices, the generalization of the notion of truth-tables by means of Nmatrices has also limitations: some logics, such as da Costa’s paraconsistent logic $C_{1}$, cannot be characterized by a single finite Nmatrix, as proved by Avron in [1, Theorem 11].
+The same is true for KT and S4, as proved by Grätz in [12].
+Moreover, in Theorem 3.5 below we will show that $\mathbf{IPL}$ also cannot be characterized by a single finite Nmatrix, extending so Gödel’s theorem from $1932$.
+These examples show that Gödel-Dugundji’s results hold, for some logics, in a stronger sense, that is, w.r.t.
+Nmatrices.
+The key to overcoming this limitation is to consider some valuations in a suitable Nmatrix.
+This kind of semantical structures was formally introduced in [7], under the name of restricted Nmatrices, or RNmatrices, showing that each da Costa’s logics $C_{n}$ (for $n\geq 1$) can be characterized by a single (decidable) $(n+2)$-valued RNmatrix; in particular, validity in $C_{1}$ can be decided by a $3$-valued RNmatrix by means of branching truth-tables (plus a simple criterion to delete inadequate rows).
+Analogously, Grätz decidable $3$-valued RNmatrices produce a decision procedure for KT and S4 by means of $3$-valued branching truth-tables (together with a simple method to delete unsound rows).
+Finite-valued non-deterministic truth-tables, together with an algorithm to delete inadequate rows, is a decision procedure that can be seen as truth-tables in a broader sense, and is a way to overcome Gödel-Dugundji’s limiting results, even for more complicated logics such as KT, S4, $C_{1}$ and $\mathbf{IPL}$.
+The aim of this paper is, precisely, to obtain a new decision procedure for $\mathbf{IPL}$ based on this kind of generalized truth-tables.
+Recall that, in $1933$, Gödel proposed an interpretation of $\mathbf{IPL}$ into S4 by means of a translation mapping $t$ between their languages [11].
+In $1948$, Tarski and McKinsey prove that Gödel’s mapping $t$ constitutes a faithful interpretation, that is: $\alpha$ is a theorem of $\mathbf{IPL}$ iff its translation $t(\alpha)$ is a theorem of S4 [25].
+Since Gödel translation $t$ is computable, then, by composing it with any algorithm for deciding S4, a decision procedure for $\mathbf{IPL}$ is immediately obtained, defined by using a translation (from $\mathbf{IPL}$ to S4).
+In particular, by composing Gödel translation $t$ with Grätz’s algorithm for S4, a decision procedure is obtained for $\mathbf{IPL}$, based on an RNmatrix for S4 (as well as the translation $t$).
+The definition of a pure RNmatrix for $\mathbf{IPL}$ (i.e., without depending on Gödel and Grätz’s constructions), which is also decidable by truth-tables, as discussed above, is a challenge that naturally arises.
+A pure RNmatrix for $\mathbf{IPL}$ with such features is desirable for several reasons.
+First, it can lead to improvements in efficiency.
+As will become clear later, the computational cost of producing a truth-table is directly related to the number of subformulas, and the translated formulas are twice as large as the original intuitionistic formulas.
+The second reason is that a pure RNmatrix can explain the actual meaning of intuitionistic operators.
+As it will be argued below, the RNmatrix for $\mathbf{IPL}$ proposed here has significant differences when compared to the RNmatrix for S4, and those differences are motivated by philosophical considerations on the nature of $\mathbf{IPL}$ itself.
+In this paper, a pure (and decidable by non-deterministic truth-tables) RNmatrix for $\mathbf{IPL}$ will be constructed by abstracting the composition of Gödel’s translation algorithm and Grätz’s decision procedure for S4.
+This abstraction will provide a direct way to define the RNmatrix for $\mathbf{IPL}$.
+Hence, a $3$-valued Nmatrix, together with suitable notions of level valuations and partial level valuations (different, in nature, to the ones for modal logics) is introduced, and the soundness and completeness of the method are proved without relying on the results of Gödel [11] and Grätz [12].
+The paper is organized as follows.
+Section 2 introduces the central notions that will be used, such as the notion of non-deterministic matrices, restricted non-deterministic matrices, and their valuation functions.
+Section 3 extends Gödel’s uncharacterizability theorem for $\mathbf{IPL}$ to the non-deterministic case.
+Section 4 presents Kearns’s Nmatrices, and the notion of level valuation for S4 is detailed.
+Section 5 presents Grätz’s RNmatrix for S4.
+Section 7 expands the RNmatrix for S4 with conjunction and disjunction and introduces Gödel’s box-translation.
+Section 8 introduces the RNmatrix for $\mathbf{IPL}$ and proves its soundness and completeness regarding a new notion of level valuation for $\mathbf{IPL}$.
+Section 9 presents the truth-table method based on the RNmatrix, and some examples are discussed.
+Section 10 briefly considers some aspects of the computational complexity of the method.
+Finally, Section 11 discusses some related works and possibilities of future research.
+2 Preliminaries: Logics, Nmatrices and RNmatrices
+Along this paper, a propositional signature is a denumerable family $\Theta=(\Theta_{n})_{n\geq 0}$ of pairwise disjoint sets.
+Elements in $\Theta_{n}$ are called $n$-ary connectives.
+An algebra over $\Theta$ is a pair $\mathcal{A}=\langle A,\mathcal{O}\rangle$ such that $A$ is a non-empty set called the universe or domain of $\mathcal{A}$ and, $\mathcal{O}$ is a function assigning, to each $n$-ary connective #\# of $\Theta$, a function $\mathcal{O}(\#):A^{n}\rightarrow A$.
+The (absolutely) free algebra over $\Theta$ generated by a denumerable set $\mathcal{P}=\{p_{0},p_{1},\ldots\}$ of propositional variables is called the algebra of formulas over $\Theta$, and will be denoted by $For(\Theta)$.
+A logic is a pair ${\bf L}=\langle For,\vdash_{\bf L}\rangle$ where $For$ is the set of formulas and $\vdash_{\bf L}$ is the consequence relation.
+A logic L is said to be Tarskian if it satisfies the following: (1) if $\varphi\in\Gamma$ then $\Gamma\vdash_{\bf L}\varphi$; (2) if $\Gamma\vdash_{\bf L}\varphi$ and $\Gamma\subseteq\Delta$ then $\Delta\vdash_{\bf L}\varphi$; and (3) if $\Delta\vdash_{\bf L}\varphi$ and $\Gamma\vdash_{\bf L}\psi$ for every $\psi\in\Delta$ then $\Gamma\vdash_{\bf L}\varphi$.
+A Tarskian logic L is finitary if it satisfies: (4) if $\Gamma\vdash_{\bf L}\varphi$ then there exists a finite subset $\Gamma_{0}$ of $\Gamma$ such that $\Gamma_{0}\vdash_{\bf L}\varphi$.
+A very natural way to characterize (Tarskian) logics is by means of logical matrices, which formalize the idea of truth-tables:
+Definition 2.1 (Logical matrices).
+Let $\Theta$ be a signature. A logical matrix over $\Theta$ is a triple $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ where
   1. 1.
-    $\mathcal{V}$ is the (non-empty) set of truth values;
+    $\mathcal{V}$ is a non-empty set (of truth-values);
   2. 2.
-    $\mathcal{D}\subset\mathcal{V}$ is the (non-empty) set of designated values;
+    $\mathcal{D}\subseteq\mathcal{V}$ is a non-empty set (of designated values);
   3. 3.
-    $\mathcal{O}$ is a function assigning, to each $n$-ary connective ##\## of $\Theta$, a (non-deterministic) truth function $\mathcal{O}(\#):\mathcal{V}^{m}\rightarrow(\wp(\mathcal{V}))\backslash\emptyset$.
-A valuation over an Nmatrix $\mathcal{M}$ is a function $v:For(\Theta)\to\mathcal{V}$ such that, for every $n$-ary connective #normal-#\## of $\Theta$ and for every $\varphi_{1},\ldots,\varphi_{n}\in For(\Theta)$, $v(\#(\varphi_{1},\ldots,\varphi_{n}))\in\mathcal{O}(\#)(v(\varphi_{1}),\ldots,v(\varphi_{n}))$.
+    $\mathcal{O}$ is a function assigning, to each $\in\Theta_{n}$, a truth function $\tilde{\#}:\mathcal{V}^{n}\rightarrow\mathcal{V}$ (that is, $\mathcal{A}_{\mathcal{M}}=\langle\mathcal{V},\mathcal{O}\rangle$ is an algebra over $\Theta$).
+Definition 2.2.
+A valuation over a matrix $\mathcal{M}$ is a homomorphism $v:For(\Theta)\to\mathcal{A}_{\mathcal{M}}$.
+That is, for every $\in\Theta_{n}$ and for every $\varphi_{1},\ldots,\varphi_{n}\in For(\Theta)$: $v(\#(\varphi_{1},\ldots,\varphi_{n}))=\mathcal{O}(\#)(v(\varphi_{1}),\ldots,v(\varphi_{n}))$.
 The set of valuations over $\mathcal{M}$ is denoted by $Val(\mathcal{M})$.
-Definition 2.2 (RNmatrix).
+We say that $\varphi$ is a semantical consequence of $\Gamma$ w.r.t.
+$\mathcal{M}$, denoted by $\Gamma\models_{\mathcal{M}}\varphi$ if, for every $v\in Val(\mathcal{M})$: $v(\varphi)\in\mathcal{D}$ whenever $v(\gamma)\in\mathcal{D}$ for every $\gamma\in\Gamma$.
+We write $\models_{\mathcal{M}}\varphi$ instead of $\emptyset\models_{\mathcal{M}}\varphi$.
+As mentioned in the Introduction, the formal notion of non-deterministic matrices (Nmatrices) was introduced by Avron and Lev in [2] as a way to overcome some Gödel-Dugundji’s uncharacterizability results.
+Finite-valued Nmatrices which are analytical (i.e., each partial valuation can be extended to a full valuation) provide a decision procedure for some logics which are uncharacterizable by a single finite-valued logic, such as some paraconsistent logics (see, for instance, [1]).
+Definition 2.3 (Nmatrices, [2]).
+Let $\Theta$ be a signature. A non-deterministic matrix (Nmatrix, for short) over $\Theta$ is a triple $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ where
+  1. 1.
+    $\mathcal{V}$ is a non-empty set (of truth-values);
+  2. 2.
+    $\mathcal{D}\subseteq\mathcal{V}$ is a non-empty set (of designated values);
+  3. 3.
+    $\mathcal{O}$ is a function assigning, to each $\in\Theta_{n}$, a (non-deterministic) truth function $\mathcal{O}(\#):\mathcal{V}^{n}\rightarrow(\wp(\mathcal{V}))\setminus\emptyset$ (that is, $\mathcal{A}_{\mathcal{M}}=\langle\mathcal{V},\mathcal{O}\rangle$ is a non-deterministic algebra over $\Theta$).
+Definition 2.4 (Valuations on Nmatrices, [2]).
+A valuation over an Nmatrix $\mathcal{M}$ is a function $v:For(\Theta)\to\mathcal{V}$ such that, for every $\in\Theta_{n}$ and for every $\varphi_{1},\ldots,\varphi_{n}\in For(\Theta)$, $v(\#(\varphi_{1},\ldots,\varphi_{n}))\in\mathcal{O}(\#)(v(\varphi_{1}),\ldots,v(\varphi_{n}))$.
+The set of valuations over $\mathcal{M}$ is denoted by $Val^{nd}(\mathcal{M})$.
+The notion of semantical consequence w.r.t.
+an Nmatrix $\mathcal{M}$, as well as the notation, will be analogous to the ones in Definition 2.2, but now w.r.t.
+valuations over the Nmatrix $\mathcal{M}$.
+However, as mentioned before, certain logics are still uncharacterizable by finite-valued Nmatrices, such as da Costa’s paraconsistent logic $C_{1}$ (see [1]), modal logics KT and S4 (see [12]), or $\mathbf{IPL}$ (see Theorem 3.5 below).
+In these cases, finite-valued and decidable RNmatrices, which are analytical and co-analytical, can overcome this limitation.
+Definition 2.5 (RNmatrices, [7]).
 Let $\Theta$ be a signature.
-An RNmatrix [3] over $\Theta$ is a pair $\mathcal{R}=\langle\mathcal{M},\mathcal{F}\rangle$ where $\mathcal{M}$ is an Nmatrix over $\Theta$ and $\mathcal{F}\subseteq Val(\mathcal{M})$.
-Definition 2.3 (Consequence relation in an (R)Nmatrix).
-Let $\mathcal{M}$ ($\mathcal{R}$, resp.) be an Nmatrix (an RNmatrix, resp.) over $\Theta$, and $\Gamma\cup\{\varphi\}\subseteq For(\Theta)$.
-We say that $\varphi$ is a semantical consequence of $\Gamma$ w.r.t.
-$\mathcal{M}$ (w.r.t.
-$\mathcal{R}$, resp.), denoted by $\Gamma\models_{\mathcal{M}}\varphi$ (denoted by $\Gamma\models_{\mathcal{R}}\varphi$, resp.) if, for every $v\in Val(\mathcal{M})$ (for every $v\in\mathcal{F}$, resp.): $v(\varphi)\in\mathcal{D}$ whenever $v(\gamma)\in\mathcal{D}$ for every $\gamma\in\Gamma$.
-We write $\models_{\mathcal{M}}\varphi$ and $\models_{\mathcal{R}}\varphi$ instead of $\emptyset\models_{\mathcal{M}}\varphi$ and $\emptyset\models_{\mathcal{R}}\varphi$.
-$\neg^{\mathcal{A}_{S4}}$
+An RNmatrix over $\Theta$ is a pair $\mathcal{R}=\langle\mathcal{M},\mathcal{F}\rangle$ where $\mathcal{M}$ is an Nmatrix over $\Theta$ and $\mathcal{F}\subseteq Val^{nd}(\mathcal{M})$.
+The notion of semantical consequence w.r.t.
+an RNmatrix $\mathcal{R}=\langle\mathcal{M},\mathcal{F}\rangle$, as well as the notation, will be analogous to the ones in Definition 2.2, but now w.r.t.
+valuations in $\mathcal{F}$.
+It is worth noting that the logic generated by an RNmatrix is indeed Tarskian.
+$\neg^{S4}$
 0 $\{1,2\}$
@@
 2 $\{0\}$
-$\Box^{\mathcal{A}_{S4}}$
+$\Box^{S4}$
 0 $\{0\}$
@@
 2 $\{2\}$
-$\rightarrow^{\mathcal{A}_{S4}}$ 0 1 2
+$\rightarrow^{S4}$ 0 1 2
 0 $\{1,2\}$ $\{1,2\}$ $\{1,2\}$
@@
 2 $\{0\}$ $\{1\}$ $\{1,2\}$
-$\lor^{\mathcal{A}_{S4}}$ 0 1 2
+$\lor^{S4}$ 0 1 2
 0 $\{0\}$ $\{1,2\}$ $\{1,2\}$
@@
 2 $\{1,2\}$ $\{1,2\}$ $\{1,2\}$
-$\land^{\mathcal{A}_{S4}}$ 0 1 2
+$\land^{S4}$ 0 1 2
 0 $\{0\}$ $\{0\}$ $\{0\}$
@@
 Table 1: Tables for $\mathcal{M}_{S4}$.
-Example 2.4 ($\mathcal{M}_{S4}$).
-Let $\Sigma_{b}=\{\{\neg,\Box\},\{\rightarrow\}\}$ and Σ={{¬,},{,,}}normal-$\Sigma=\{\{\neg,\Box\},\{\rightarrow,\lor,\land\}\}$ be, respectively, the basic signature and the signature for modal logic S4.
+Example 2.6 (Grätz’s Nmatrix for S4).
+Let $\Sigma_{b}=\{\{\neg,\Box\},\{\rightarrow\}\}$ and $\Sigma=\{\{\neg,\Box\},\{\rightarrow,\lor,\land\}\}$ be, respectively, the basic signature and the signature for modal logic S4.
 Given the set $\mathcal{P}$ of propositional variables, let $For(\Sigma_{b})$ and $For(\Sigma)$ be the corresponding propositional languages.
-Grätz’s Nmatrix over $\Sigma_{b}$ is the tuple $\mathcal{M}_{S4}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$, where $\mathcal{V}=\{0,1,2\}$, $\mathcal{D}=\{1,2\}$, $\mathcal{O}=\{\neg^{\mathcal{A}_{S4}},\Box^{\mathcal{A}_{S4}},\rightarrow^{\mathcal{A}_{S4}}\}$ and each $c\in\mathcal{O}$ is defined in Table 1.
-$p$$\Box p$$\Box\Box p$$\Box p\rightarrow\Box\Box p$000121001222212
-(a) Valuations in $\mathcal{M}_{S4}$ for $\Box p\rightarrow\Box\Box p$
-Figure 1: Each tree level corresponds to the set of values resulting from the multioperation applied to the respective subformula (at left) in that tree.
-At the top are the values that the atomic formula can receive, and in the leaves are the values of the evaluated formula.
-By setting $\varphi\lor\psi:=\neg\varphi\to\psi$ and $\varphi\land\psi:=\neg(\varphi\to\neg\psi)$, respectively, the multioperators $\lor^{\mathcal{A}_{S4}}$ and $\land^{\mathcal{A}_{S4}}$ are defined as in Table 1.
-Valuations in an Nmatrix are typically represented as tables.
-However, if one focuses more on the computational aspects of these valuations, adopting the more flexible data structure of binary trees can be advantageous.
-In Figure 1, we show how to compute the valuation of two examples.
-3 Kearns’s level valuations
-Regarding semantics for modal logic, Kripke-style semantics are the usual choice.
-However, there are other attractive alternatives.
-In the 1980s, Kearns proposed an RNmatrix-based semantics for modal logics KT, S4 and S5.
-He introduced the level valuations as the set $\mathcal{F}$ of valuations to be used.
-[Kearn’s account] is simpler than [Kripke’s account] in virtue of having dispensed with possible worlds and their relations.
+Grätz’s Nmatrix over $\Sigma_{b}$ is the tuple $\mathcal{M}_{S4}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$, where $\mathcal{V}=\{0,1,2\}$, $\mathcal{D}=\{1,2\}$, $\mathcal{O}=\{\neg^{S4},\Box^{S4},\rightarrow^{S4}\}$ and each $c\in\mathcal{O}$ is defined in Table 1.
+We will return to this Nmatrix in Section 5.
+By setting $\varphi\lor\psi:=\neg\varphi\to\psi$ and $\varphi\land\psi:=\neg(\varphi\to\neg\psi)$, respectively, the multioperators $\lor^{S4}$ and $\land^{S4}$ are defined as in Table 1.
+As proven in [12], no finite Nmatrix can characterize S4; in particular, $\mathcal{M}_{S4}$ cannot characterize this modal logic.
+In Section 5 we will analyze an RNmatrix based on $\mathcal{M}_{S4}$ which characterizes S4.
+3 Gödel’s uncharacterizability theorem updated
+As mentioned in the Introduction, Gödel proved in $1932$ that $\mathbf{IPL}$ cannot be characterized by a single finite-valued (deterministic) logical matrix.
+It is a natural question to determine whether $\mathbf{IPL}$ could be characterized by means of a single finite-valued non-deterministic matrix.
+As we shall show in the sequel, it is easy to adapt the proof by Gödel in [10] in order to see that the answer to the question above is negative.
+Definition 3.1.
+Consider the signature $\Omega=(\{\neg\},\{\land,\lor,\rightarrow\})$ for intuitionistic propositional logic. The set of intuitionistic formulas will be denoted by $For(\Omega)$.
+Definition 3.2 (Gödel [10]).
+For $n\geq 1$ let $G_{n}$ be the formula expressing that “there are at most $n$ truth-values”:
+$G_{n}:=\bigvee_{1\leq i<j\leq n+1}(p_{i}\to p_{j})\land(p_{j}\to p_{i}).$
+Lemma 3.3 (Gödel [10]).
+None of the formulas $G_{n}$ is provable in IPL.
+Lemma 3.4.
+If $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ is a finite Nmatrix such that $\mathcal{V}$ has exactly $n\geq 1$ elements and $\mathcal{M}$ models IPL, then the formula $G_{n}$ is valid in $\mathcal{M}$.
+Proof.
+For each connective $\in\Omega$ let us denote $\mathcal{O}(\#)$ by $\bar{\#}$.
+Since $\mathcal{M}$ models IPL then, for every $a,b\in\mathcal{V}$: (i) $a\,\bar{\vee}\,b\subseteq\mathcal{D}$ if either $a\in\mathcal{D}$ or $b\in\mathcal{D}$; (ii) $a\,\bar{\land}\,b\subseteq\mathcal{D}$ if both $a\in\mathcal{D}$ and $b\in\mathcal{D}$; and (iii) $a\,\bar{\to}\,a\subseteq\mathcal{D}$ (since $\varphi\to\varphi$ is a theorem of IPL).
+Given that $\mathcal{V}$ has exactly $n\geq 1$ elements, if $v$ is a valuation over $\mathcal{M}$ then, by the pigeonhole principle, there exists $1\leq i<j\leq n+1$ such that $v(p_{i})=v(p_{j})$.
+Let $a:=v(p_{i})$.
+By consideration (iii) above, $v(p_{i}\to p_{j})\in a\,\bar{\to}\,a\subseteq\mathcal{D}$, hence $v(p_{i}\to p_{j})\in\mathcal{D}$.
+Analogously, $v(p_{j}\to p_{i})\in\mathcal{D}$ and so, by (ii), $v((p_{i}\to p_{j})\land(p_{j}\to p_{i}))\in\mathcal{D}$.
+Thus, from the fact that $G_{n}$ is a disjunction where $(p_{i}\to p_{j})\land(p_{j}\to p_{i})$ is one of its disjuncts, repeated application of (i) implies $v(G_{n})\in\mathcal{D}$.
+∎
+Theorem 3.5.
+The logic IPL cannot be characterized by a single finite-valued Nmatrix.
+Proof.
+Suppose that $\mathcal{M}$ is a finite Nmatrix with exactly $n\geq 1$ elements such that $\mathcal{M}$ models IPL.
+Suppose, by absurdum, that $\mathcal{M}$ characterizes IPL.
+Hence, every formula valid in $\mathcal{M}$ can be proved in IPL.
+By Lemma 3.4, the formula $G_{n}$ is valid in $\mathcal{M}$.
+By hypothesis, $G_{n}$ must be provable in IPL.
+But this contradicts Lemma 3.3.
+This shows that no finite Nmatrix $\mathcal{M}$ can characterize IPL.
+∎
+4 Kearns’s characterization of S4 by RNmatrices
+Regarding semantics for normal modal logic, Kripke-style semantics are the usual choice.
+However, there are other attractive alternatives, such as algebraic semantics.
+In the 1980s, Kearns proposed in [19] an RNmatrix-based semantics for modal logics KT, S4 and S5.
+He introduced the level valuations as the set $\mathcal{F}$ of valuations to be used in his RNmatrix.
+According to him,
+[It] is simpler than the standard [Kripke’s] account in virtue of having dispensed with possible worlds and their relations.
 I also think that my account is philosophically preferable to the standard account for having done this.
 For I do not think there are such things as possible worlds, or even that they constitute a useful fiction.
-[6]
-We will define the level valuations for the modal logic S4.
-Figure 2: Assume that $Val(\mathcal{M}_{S4})=\{v_{0},v_{1},\ldots,v_{m},v_{n},v_{k},\ldots\}$.
-We remove a valuation $v$ in a level $\mathcal{L}_{i+1}$ iff there is some wf $\alpha$ such that $v(\alpha)=1$ and $w(\alpha)\in\mathcal{D}$ for every $w\in\mathcal{L}_{i}$.
+(Kearns, [19, p.
+86])
+Figure 1 reconstructs Kearns’s Nmatrix for S4 over the signature $\Sigma_{Ke}=(\{\Box,\neg\},\{\vee\})$, where the truth-values are: $T$ (necessarily true), $t$ (contingently true), $f$ (contingently false) and $F$ (necessarily false).
+In this Nmatrix, the set of designated values is given by $\mathcal{D}=\{T,t\}$.
+Figure 1: Tables for $\mathcal{M}_{Ke}$.
+Kearns proved that, in order to fully characterize S4 using $\mathcal{M}_{Ke}$, it is sufficient to introduce the notion of level valuations, which are defined as follows:
 Definition 4.1.
-(Level valuations for S4)
-Assume that $Val(\mathcal{M}_{S4})$ is the set of valuations over $\mathcal{M}_{S4}$. We define the level $\mathcal{L}^{S4}_{k}$, where $k\in\mathbb{N}$, as follows:
+(Kearns’s level valuations for S4)
+Let $Val^{nd}(\mathcal{M}_{Ke})$ be the set of valuations over $\mathcal{M}_{Ke}$. We define the set of $n$th-level valuations $\mathcal{L}^{Ke}_{k}$, where $n\in\omega$, as follows:
   1. i)
-    $\mathcal{L}_{0}^{S4}=Val(\mathcal{M}_{S4})$;
+    $\mathcal{L}_{0}^{Ke}=Val^{nd}(\mathcal{M}_{Ke})$;
   2. ii)
-    $\mathcal{L}^{S4}_{k+1}=\{v\in\mathcal{L}^{S4}_{k}\mid\forall\alpha\in For(\Sigma_{b}),\mathcal{L}^{S4}_{k}(\alpha)\subseteq\mathcal{D}\Rightarrow v(\alpha)=2\}$
-where $\mathcal{L}^{S4}_{k}(\alpha)=\{v(\alpha)\mid v\in\mathcal{L}^{S4}_{k}\}$.
+    $\mathcal{L}^{Ke}_{n+1}=\{v\in\mathcal{L}^{Ke}_{n}\mid\forall\alpha\in For(\Sigma_{Ke}),\mathcal{L}^{Ke}_{n}(\alpha)\subseteq\{T,t\}\Rightarrow v(\alpha)=T\}$
+where $\mathcal{L}^{Ke}_{n}(\alpha)=\{v(\alpha)\mid v\in\mathcal{L}^{Ke}_{n}\}$.
+The set of level valuations in $\mathcal{M}_{Ke}$ is the intersection of the sets $\mathcal{L}^{Ke}_{n}$, for $n\geq 0$:
+$\mathcal{L}_{Ke}=\bigcap^{\infty}_{n\geq 0}\mathcal{L}^{Ke}_{n}.$
+The way in which level valuations work can be appreciated in Figure 2.
+Figure 2: Assume that $Val^{nd}(\mathcal{M}_{Ke})=\{v_{0},v_{1},\ldots,v_{m},v_{n},v_{k},\ldots\}$.
+We remove a valuation $v$ in level $\mathcal{L}_{n+1}$ iff there is some $\alpha$ such that $v(\alpha)=t$ and $w(\alpha)\in\{T,t\}$ for every $w\in\mathcal{L}_{n}$.
+Definition 4.2 (Kearns’s RNmatrix for S4).
+The RNmatrix for S4 defined by Kearns is the pair $\mathcal{R}(\mathcal{M}_{Ke})=\langle\mathcal{M}_{Ke},\mathcal{L}_{Ke}\rangle$.
+Theorem 4.3 ([19]).
+For every $\Gamma$ and $\varphi$ over $\Sigma_{Ke}$: $\Gamma\vdash_{S4}\varphi$ iff $\Gamma\models_{\mathcal{R}(\mathcal{M}_{Ke})}\varphi$.
+5 Grätz’s decision procedure for S4
+Kearns proved that his Nmatrix for S4 with level valuations is sound and complete regarding the logic S4.
+However, as Section 4 explains, level valuations are defined as the intersection of an infinite family of subsets of valuation functions.
+How do we reproduce this procedure using finite steps?
+While it is true that level valuation semantics characterizes S4, it is not straightforward to decide the validity of a given formula using Kearns’s tables.
+In particular, given any theorem $\varphi$, the immediate table for $\Box\varphi$ will always present a partial valuation $v$ such that $v(\Box\varphi)$ is not a designated value, thus invalidating the rule of necessitation.
+Grätz noticed that the problem lies in the definition of partial valuation: such partial valuations are precisely the ones that cannot be extended to full level valuations; thus, they should be ruled out by some appropriate definition.
+Grätz’s Nmatrix for S4:
+Instead of $4$ truth-values, Grätz use a set with three values, $0,1$ and $2$, where each one has the following intended meaning:
+$v(\varphi)=0$
+means that $\varphi$ is false;
+$v(\varphi)=1$
+means that $\varphi$ is contingently true, i.e, $\varphi$ is true but possibly false;
+$v(\varphi)=2$
+means that $\varphi$ is necessarily true.
+From this interpretation, defining the subset $\mathcal{D}=\{2,1\}$ as the set of designated values is natural.
+As primitive operators, it is used negation ($\neg$), implication ($\to$), and box ($\Box$).
+The Nmatrix $\mathcal{M}_{S4}$ over $\Sigma_{b}=(\{\neg,\Box\},\{\rightarrow\})$ is defined as in Example 2.6 (see also Table 1).
+To understand the Nmatrix for the modal operator $\Box$, we must recall the axioms that define S4.
+In particular, the axioms $\Box\varphi\to\varphi$ (reflexivity) and $\Box\varphi\to\Box\Box\varphi$ (transitivity).
+Consider the contrapositive of the reflexivity axiom: $\neg\varphi\to\neg\Box\varphi$.
+This means that if $\varphi$ is false, then $\Box\varphi$ is also false.
+The Nmatrix for $\Box$ reflects this behavior when $v(\varphi)=0$.
+Next, when $v(\varphi)=2$ (indicating that $\varphi$ is necessary, or in other words, $\Box\varphi$ is true), we have $v(\Box\varphi)=\{2\}$.
+This signifies that $\Box\varphi$ is also necessary, which implies that $\Box\Box\varphi$ is true as well.
+Finally, when $v(\varphi)=1$, it indicates that $\varphi$ is contingent, meaning it is possibly false, hence $\Box\varphi$ is false.
 Definition 5.1.
-The set of level valuations in $\mathcal{M}_{S4}$ is defined as the intersection of every level valuation:
-$\mathcal{L}_{S4}=\cap^{\infty}_{k\geq 0}\mathcal{L}^{S4}_{k}$
+(Grätz’s level valuations for S4)
+Let $Val^{nd}(\mathcal{M}_{S4})$ be the set of valuations over $\mathcal{M}_{S4}$. We define the set of $n$th-level valuations $\mathcal{L}^{S4}_{k}$, where $n\in\omega$, as follows:
+  1. i)
+    $\mathcal{L}_{0}^{S4}=Val^{nd}(\mathcal{M}_{S4})$;
+  2. ii)
+    $\mathcal{L}^{S4}_{n+1}=\{v\in\mathcal{L}^{S4}_{n}\mid\forall\alpha\in For(\Sigma_{b}),\mathcal{L}^{S4}_{n}(\alpha)\subseteq\{1,2\}\Rightarrow v(\alpha)=2\}$
+where $\mathcal{L}^{S4}_{n}(\alpha)=\{v(\alpha)\mid v\in\mathcal{L}^{S4}_{n}\}$.
+The set of level valuations in $\mathcal{M}_{S4}$ is the intersection of the sets $\mathcal{L}^{S4}_{n}$, for $n\geq 0$:
+$\mathcal{L}_{S4}=\bigcap^{\infty}_{n\geq 0}\mathcal{L}^{S4}_{n}.$
 Definition 5.2 (Grätz’s RNmatrix for S4).
-Let $\mathcal{M}_{S4}$ be the nmatrix defined in Definition 2.4. The RNmatrix for S4 is the pair $\mathcal{R}(\mathcal{M}_{S4})=\langle\mathcal{M}_{S4},\mathcal{L}_{S4}\rangle$.
+The RNmatrix for S4 defined by Grätz is the pair $\mathcal{R}(\mathcal{M}_{S4})=\langle\mathcal{M}_{S4},\mathcal{L}_{S4}\rangle$.
 Let $\vdash_{S4}^{b}$ be the consequence relation of the standard Hilbert calculus $\mathcal{H}^{b}_{S4}$ for S4 over $\Sigma_{b}$.
 As usual, $\Gamma\vdash_{S4}^{b}\varphi$ iff either $\varphi$ is a theorem of $\mathcal{H}^{b}_{S4}$ or there exists a non-empty set $\{\beta_{1},\ldots,\beta_{n}\}\subseteq\Gamma$ such that $\beta_{1}\to(\beta_{2}\to(\ldots\to(\beta_{n}\to\varphi)\ldots))$ is a theorem of $\mathcal{H}^{b}_{S4}$.
-Theorem 3.4 ([4]).
-For every $\Gamma$ and $\varphi$: $\Gamma\vdash_{S4}^{b}\varphi$ iff $\Gamma\models_{\mathcal{R}(\mathcal{M}_{S4})}\varphi$.
-3.1 Grätz’s decision procedure for S4
-While the soundness and completeness of level valuations for S4 (with the rule of necessitation) were already known, it was only recently that an effective decision procedure was discovered.
-Before that, J.
-V.
-Ivlev seems to have suggested the following approach [4]: remove every partial valuation $v_{p}$ whenever (1) $v_{p}(\Box A)\not\in D$ and (2) $w_{p}(A)\in D$ for every partial valuation $w_{p}$.
-The reason for this method is very intuitive if we consider the definition of level valuations, and indeed, it works for $2$-valued semantics, as well as for some examples in $\mathcal{M}_{S4}$, such as $\Box\alpha$ where $\alpha$ is any propositional tautology.
-However, this criterion is not sufficient for $\mathcal{M}_{S4}$ because it leaves behind invalid partial level valuations (consider the valid formula $\Box p\rightarrow\Box\neg\neg p$ in $\mathcal{M}_{S4}$).
-In [4], Grätz identified the need to find a suitable method by refining the concept of partial valuation. To address this, he suggests a new definition for partial level valuations:
+Theorem 5.3.
+([12, Theorems 3.5 and 3.10]) For every $\Gamma$ and $\varphi$ over $\Sigma_{b}$: $\Gamma\vdash_{S4}^{b}\varphi$ iff $\Gamma\models_{\mathcal{R}(\mathcal{M}_{S4})}\varphi$.
+In order to obtain a decision procedure for S4 based on the $3$-valued RNmatrix $\mathcal{R}(\mathcal{M}_{S4})$, Grätz proposes the notion of partial level valuation.
+The idea is as follows: given a formula $\varphi$ to be checked as a tautology in S4, construct the (finite) table generated by the Nmatrix $\mathcal{M}_{S4}$ for $\varphi$ and all of its subformulas.
+Then, by means of a decision procedure, it is possible to determine which rows of the table are allowed (since they correspond to level valuations); the non-allowed ones correspond to valuations over $\mathcal{M}_{S4}$ which do not correspond to level valuations, hence they are deleted.
+If $\varphi$ only receives a designated value in the allowed rows, then it is declared to be valid in S4, otherwise it is declared to be non-valid.
+Formally:
 Definition 5.4 (Partial valuation in $\mathcal{M}_{S4}$).
@@
 and, for every $\alpha,\beta\in\Lambda$:
-[…]
-  • if $\alpha\to\beta\in\Lambda$ then $\tilde{v}_{p}(\alpha\to\beta)\in{\to}^{\mathcal{A}_{S4}}(\tilde{v}_{p}(\alpha),\tilde{v}_{p}(\beta))$.
+  • if $#\in\{\neg,\Box\}$ and $\alpha\in\Lambda$ then $\tilde{v}_{p}(\#\alpha)\in\#^{S4}(\tilde{v}_{p}(\alpha))$;
+  • if $\alpha\to\beta\in\Lambda$ then $\tilde{v}_{p}(\alpha\to\beta)\in{\to}^{S4}(\tilde{v}_{p}(\alpha),\tilde{v}_{p}(\beta))$.
 Definition 5.5 (Grätz’s partial level valuation).
-A partial valuation $v_{p}:\Lambda\rightarrow\{0,1,2\}$ in $\mathcal{M}_{S4}$ is a partial level valuation in $\mathcal{R}(\mathcal{M}_{S4})$ ($v_{p}\in PLV$) iff
+Let $\Lambda\subseteq For(\Sigma_{b})$ be a set of formulas closed under subformulas.
+A partial valuation $v_{p}:\Lambda\rightarrow\{0,1,2\}$ in $\mathcal{M}_{S4}$ is a partial level valuation over $\Lambda$ in $\mathcal{R}(\mathcal{M}_{S4})$ ($v_{p}\in PLV$) iff
   1. 1.
@@
 Remark 5.6.
-In [4, p.
-15] it is observed that, being $\Lambda$ finite, the set of $PLV$s with domain $\Lambda$ is well-defined (that is, it is not cyclic), given that the search space is finite: the set of all the partial valuations with domain $\Lambda$ is finite.
-Moreover, it is proved in [4] that this notion is analytic in $\mathcal{M}_{S4}$ (i.e, every partial level valuation in $\mathcal{M}_{S4}$ can be extended to a level valuation in $\mathcal{M}_{S4}$) and co-analytic (i.e, for any level valuation $v$ in $\mathcal{M}_{S4}$, the restriction of $v$ to a partial level valuation $v_{p}$ defined over a finite set $\Lambda$ closed under subformulas is a partial level valuation).
+The set $PLV$ plays a fundamental role in the proof of soundness and completeness of the truth-tables decision procedure introduced in [12]. Some comments on this notion are in order at this point.
+The existence (and uniqueness) of the set $PLV$ is the crucial fact used in his proof.
+In this respect, [12, p.
+15] presents an algorithm (Algorithm 3.4) for computing the set $PLV$ of partial level valuations as a (constructive) truth-table.
+He argues that this process is well-defined since the set of all partial valuations with domain $\Lambda$ is finite, so is its powerset.
+Uniqueness is tacitly assumed, given the nature of the algorithm.
+Moreover, he proves that this notion is analytic in $\mathcal{M}_{S4}$ (i.e., every partial level valuation in $\mathcal{M}_{S4}$ can be extended to a level valuation in $\mathcal{M}_{S4}$) and co-analytic (i.e., for any level valuation $v$ in $\mathcal{M}_{S4}$, the restriction of $v$ to a partial level valuation $v_{p}$ defined over a finite set $\Lambda$ closed under subformulas is a partial level valuation).
+This argument, justified by the uniqueness of the set $PLV$, allows us to conclude that the truth-table procedure corresponds exactly to the level valuation semantics.
+A similar and fully detailed argument will be given for $\mathbf{IPL}$ in Section 9.
+From the observations above, one can easily prove that the truth-table method for S4 is sound and complete, that is:
 Theorem 5.7.
-$\mathcal{R}(\mathcal{M}_{S4})$ is analytic and co-analytic with PLV.
-Using this, one can easily prove that the truth table method is sound and complete, that is,
-Theorem 3.9 ([4]).
-For every $\varphi$: $\models_{\mathcal{R}(\mathcal{M}_{S4})}\varphi$ if, and only if,for every partial valuation $v_{p}$ in $\mathcal{M}_{S4}$ defined over the set $\Lambda$ of subformulas of $\varphi$: if $v_{p}\in PLV$, then $v_{p}(\varphi)\in\mathcal{D}$.
-Corollary 3.9.1 ([4]).
-For every $\varphi$: $\vdash_{S4}^{b}\varphi$ if, and only if, for every partial valuation $v_{p}$ in $\mathcal{M}_{S4}$ defined over the set $\Lambda$ of subformulas of $\varphi$: if $v_{p}\in PLV$ then $v_{p}(\alpha)\in\mathcal{D}$.
+([12, Theorem 3.15]) Let $\varphi$ be a formula over $\Sigma_{b}$.
+Then, $\models_{\mathcal{R}(\mathcal{M}_{S4})}\varphi$ if, and only if, for every partial valuation $v_{p}$ in $\mathcal{M}_{S4}$ defined over the set $\Lambda$ of subformulas of $\varphi$: if $v_{p}\in PLV$, then $v_{p}(\varphi)\in\mathcal{D}$.
+Corollary 5.7.1.
+For every $\varphi$ over $\Sigma_{b}$: $\vdash_{S4}^{b}\varphi$ if, and only if, for every partial valuation $v_{p}$ in $\mathcal{M}_{S4}$ defined over the set $\Lambda$ of subformulas of $\varphi$, if $v_{p}\in PLV$ then $v_{p}(\alpha)\in\mathcal{D}$.
 Proof.
-It is an immediate consequence of Theorems 3.9 and 3.4. ∎
-Therefore, to construct a truth table, compute every partial valuation and remove any non-PLV.
-And this algorithmic process is a sound and complete decision procedure for S4.
-Obviously the soundness and completeness Theorem 3.9 (and so, the decision procedure) can be extended to inferences in S4 from a finite set of premises, taking into account that $\vdash_{S4}^{b}$ (and so $\models_{\mathcal{R}(\mathcal{M}_{S4})}$) satisfies the deduction metatheorem.
-3.2 Reducing Grätz’s RNmatrix and algorithm for S4
+It is an immediate consequence of Theorems 5.7 and 5.3. ∎
+Therefore, to construct a truth-table, compute every partial valuation and remove any non-$PLV$ in a finite number of steps.
+This algorithmic process is a sound and complete decision procedure for S4.
+Clearly, the soundness and completeness Theorem 5.7 (and hence the decision procedure) can be extended to inferences in S4 from a finite set of premises, since $\vdash_{S4}^{b}$ (and thus $\models_{\mathcal{R}(\mathcal{M}_{S4})}$) satisfies the deduction metatheorem.
+6 Reducing Grätz’s RNmatrix and algorithm for S4
 As suggested by Grätz ([12], p.
@@
 Along this section we will deal exclusively with signature $\Sigma$ for S4.
-[…]
-$\neg^{\mathcal{A}_{S4}}$
+It is easy to see that the reduced Nmatrix for S4 over signature $\Sigma$ (recall Example 2.6 and Table 1) obtained in this way is $\mathcal{M}^{\prime}_{S4}=\langle\mathcal{V},\mathcal{D},\mathcal{O}^{\prime}\rangle$, where $\mathcal{O}^{\prime}=\{\neg^{S4},\Box^{S4},\rightarrow^{\prime S4},\lor^{\prime S4},\land^{\prime S4}\}$ such that the tables for $^{\prime S4}$, for $#\in\{\to,\vee,\land\}$, are displayed in Table 2.
+$\neg^{S4}$
 0 $\{1,2\}$
@@
 2 $\{0\}$
-$\Box^{\mathcal{A}_{S4}}$
+$\Box^{S4}$
 0 $\{0\}$
@@
 2 $\{2\}$
-$\rightarrow^{\mathcal{A}^{\prime}_{S4}}$ 0 1 2
+$\rightarrow^{\prime S4}$ 0 1 2
 0 $\{1,2\}$ $\{1,2\}$ $\{2\}$
@@
 2 $\{0\}$ $\{1\}$ $\{2\}$
-$\lor^{\mathcal{A}^{\prime}_{S4}}$ 0 1 2
+$\lor^{\prime S4}$ 0 1 2
 0 $\{0\}$ $\{1,2\}$ $\{2\}$
@@
 2 $\{2\}$ $\{2\}$ $\{2\}$
-$\land^{\mathcal{A}^{\prime}_{S4}}$ 0 1 2
+$\land^{\prime S4}$ 0 1 2
 0 $\{0\}$ $\{0\}$ $\{0\}$
@@
 Table 2: Tables for $\mathcal{M}^{\prime}_{S4}$.
-Of course this move requires adjusting the soundness and completeness proofs of Grätz to the new RNmatrix defined over signature $\Sigma$.
+Note that the Nmatrices in Table 2 were obtained by computing truth-tables, and therefore filtered sets of partial valuations.
+Because of this, some rows were excluded.
+In particular, if $v(\beta)=2$ and $v(\alpha\to\beta)=1$, we cannot find a partial valuation $w$ such that $w(\alpha\to\beta)=0$ and $w(\beta)=2$.
+Therefore, such a partial valuation cannot be a partial level valuation and we can safely remove it from the definition of implication.
+This was also discussed in [12, p.
+17].
+Of course considering the reduced Nmatrix over signature $\Sigma$ requires adjusting the soundness and completeness proofs of Grätz to the new RNmatrix defined over $\Sigma$.
 Let $\vdash_{S4}$ be the consequence relation of the standard Hilbert calculus $\mathcal{H}_{S4}$ for S4 over $\Sigma$.
@@
 Consider the family of level $\mathcal{L}^{\prime S4}_{k}$ ($k\geq 0$) as in Definition 4.1, but now starting from ${\mathcal{L}^{\prime}}_{0}^{S4}=Val(\mathcal{M}^{\prime}_{S4})$ in item i), and considering formulas $\alpha\in For(\Sigma)$ in item ii).
-Let $\mathcal{L}^{\prime}_{S4}=\cap^{\infty}_{k\geq 0}\mathcal{L}^{\prime S4}_{k}$ and $\mathcal{R}(\mathcal{M}^{\prime}_{S4})=\langle\mathcal{M}^{\prime}_{S4},\mathcal{L}^{\prime}_{S4}\rangle$.
+Let $\mathcal{L}^{\prime}_{S4}=\bigcap^{\infty}_{k\geq 0}\mathcal{L}^{\prime S4}_{k}$ and $\mathcal{R}(\mathcal{M}^{\prime}_{S4})=\langle\mathcal{M}^{\prime}_{S4},\mathcal{L}^{\prime}_{S4}\rangle$.
 Remark 6.1.
@@
 Proof.
-It coincides with the proof for $\mathcal{R}(\mathcal{M}_{S4})$ found in [4, Theorems 3.4 and 3.5].
+It coincides with the proof for $\mathcal{R}(\mathcal{M}_{S4})$ found in [12, Theorems 3.4 and 3.5].
 The only detail to be checked is that any axiom of $\mathcal{H}_{S4}$ is still valid in the reduced Nmatrix $\mathcal{M}^{\prime}_{S4}$.
@@
 ∎
-For completeness, recall that, given a Tarskian and finitary logic L and a formula $\varphi$, a set $\Delta$ of formulas is said to be $\varphi$-saturated in L if $\Delta\nvdash_{\bf L}\varphi$, but $\Delta,\alpha\vdash_{\bf L}\varphi$, whenever $\alpha\notin\Delta$.
+The proof of completeness requires the use of a fundamental notion: $\varphi$-saturated sets.
+Remark 6.3.
+Recall that, given a Tarskian and finitary logic L and a formula $\varphi$, a set $\Delta$ of formulas is said to be $\varphi$-saturated in L if $\Delta\nvdash_{\bf L}\varphi$, but $\Delta,\alpha\vdash_{\bf L}\varphi$, whenever $\alpha\notin\Delta$.
 It is a well-known result that, if $\Gamma\nvdash_{\bf L}\varphi$, there exists a $\varphi$-saturated set $\Delta$ in L such that $\Gamma\subseteq\Delta$.
@@
 Proof.
-[…]
+The unary multioperators interpreting $\neg$ and $\Box$ in $\mathcal{M}^{\prime}_{S4}$ and in the Nmatrix $\mathcal{M}_{S4}$ for S4 considered in [12] coincide, while $\to^{\prime S4}$ is different to the corresponding multioperator in $\mathcal{M}_{S4}$ (which is precisely $\to^{S4}$).
+In turn, the multioperators $\lor^{\prime S4}$ and $\land^{\prime S4}$ were not considered in [12, Lemma 3.7].
+Thus, it is only required to prove that $v_{\Delta}(\alpha\#\beta)\in\#^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$ for every $\alpha,\beta$ and $#\in\{\lor,\land,\to\}$.
 Disjunction: Suppose that $v_{\Delta}(\alpha\vee\beta)=2$.
@@
 This implies that $\alpha\in\Delta$ or $\beta\in\Delta$, hence $v_{\Delta}(\alpha)\in\mathcal{D}$ or $v_{\Delta}(\beta)\in\mathcal{D}$.
-This implies that $v_{\Delta}(\alpha\vee\beta)=2\in\vee^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+This implies that $v_{\Delta}(\alpha\vee\beta)=2\in\vee^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 Suppose now that $v_{\Delta}(\alpha\vee\beta)=1$.
@@
 This means that $v_{\Delta}(\alpha)=1$ and $v_{\Delta}(\beta)\in\{0,1\}$, or vice versa.
-Then, $v_{\Delta}(\alpha\vee\beta)=1\in\{1,2\}=\vee^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+Then, $v_{\Delta}(\alpha\vee\beta)=1\in\{1,2\}=\vee^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 Finally, if $v_{\Delta}(\alpha\vee\beta)=0$ then $\alpha\vee\beta\notin\Delta$, hence $\alpha\notin\Delta$ and $\beta\notin\Delta$.
-This means that $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\vee\beta)=0\in\{0\}=\vee^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+This means that $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\vee\beta)=0\in\{0\}=\vee^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 Conjunction: Suppose that $v_{\Delta}(\alpha\land\beta)=2$.
@@
 This implies that $\Box\alpha\in\Delta$ and $\Box\beta\in\Delta$, hence $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=2$.
-Thus, $v_{\Delta}(\alpha\land\beta)=2\in\{2\}=\land^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+Thus, $v_{\Delta}(\alpha\land\beta)=2\in\{2\}=\land^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 Now, assume that $v_{\Delta}(\alpha\land\beta)=1$.
@@
 That is, $v_{\Delta}(\alpha),v_{\Delta}(\beta)\in\{1,2\}$ and either $v_{\Delta}(\alpha)=1$ or $v_{\Delta}(\beta)=1$.
-Then, $v_{\Delta}(\alpha\land\beta)=1\in\{1\}=\land^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+Then, $v_{\Delta}(\alpha\land\beta)=1\in\{1\}=\land^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 Finally, if $v_{\Delta}(\alpha\land\beta)=0$ then $\alpha\land\beta\notin\Delta$, hence $\alpha\notin\Delta$ or $\beta\notin\Delta$.
-This means that $v_{\Delta}(\alpha)=0$ or $v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\land\beta)=0\in\{0\}=\land^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+This means that $v_{\Delta}(\alpha)=0$ or $v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\land\beta)=0\in\{0\}=\land^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 Implication: Suppose that $v_{\Delta}(\alpha\to\beta)=2$.
@@
 If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)=0$.
-Thus, $v_{\Delta}(\alpha\to\beta)=2\in{\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+Thus, $v_{\Delta}(\alpha\to\beta)=2\in{\to}^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 If $\beta\in\Delta$, suppose first that $\Box\alpha\in\Delta$.
 Then, by axiom K it follows that $\Box\beta\in\Delta$.
-Hence, $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=2$ and so $v_{\Delta}(\alpha\to\beta)=2\in\{2\}={\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+Hence, $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=2$ and so $v_{\Delta}(\alpha\to\beta)=2\in\{2\}={\to}^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 Otherwise, if $\Box\alpha\notin\Delta$ then $v_{\Delta}(\alpha)\neq 2$ and $v_{\Delta}(\beta)\neq 0$.
-From this, $v_{\Delta}(\alpha\to\beta)=2\in{\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+From this, $v_{\Delta}(\alpha\to\beta)=2\in{\to}^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 Suppose now that $v_{\Delta}(\alpha\to\beta)=1$.
@@
 If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)=0$ and $v_{\Delta}(\beta)\in\{0,1\}$.
-Thus, $v_{\Delta}(\alpha\to\beta)=1\in\{1,2\}={\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
-Otherwise, if $\beta\in\Delta$ then $v_{\Delta}(\beta)=1$ and so $v_{\Delta}(\alpha\to\beta)=1\in{\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+Thus, $v_{\Delta}(\alpha\to\beta)=1\in\{1,2\}={\to}^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+Otherwise, if $\beta\in\Delta$ then $v_{\Delta}(\beta)=1$ and so $v_{\Delta}(\alpha\to\beta)=1\in{\to}^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 Finally, if $v_{\Delta}(\alpha\to\beta)=0$ then $\alpha\to\beta\notin\Delta$, hence $\alpha\in\Delta$ and $\beta\notin\Delta$.
-This means that $v_{\Delta}(\alpha)\in\{1,2\}$ and $v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\to\beta)=0\in\{0\}={\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+This means that $v_{\Delta}(\alpha)\in\{1,2\}$ and $v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\to\beta)=0\in\{0\}={\to}^{\prime S4}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 This completes the proof. ∎
@@
 Proof.
-It is analogous to the proof for $\mathcal{R}(\mathcal{M}_{S4})$ found in [4, Lemma 3.9 and Theorem 3.10], but now by using our Lemma 3.12 instead of [4, Lemma 3.7]. ∎
-Now, Definitions 3.5 and 3.6 can be easily adapted to $\mathcal{M}^{\prime}_{S4}$, obtaining so the set $PLV^{\prime}$ of partial level valuations in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over finite sets of formulas.
+It is analogous to the proof for $\mathcal{R}(\mathcal{M}_{S4})$ found in [12, Lemma 3.9 and Theorem 3.10], but now by using our Lemma 6.4 instead of [12, Lemma 3.7]. ∎
+Now, Definitions 5.4 and 5.5 can be easily adapted to $\mathcal{M}^{\prime}_{S4}$, obtaining so the set $PLV^{\prime}$ of partial level valuations in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over finite sets of formulas.
 However, it will be convenient to give a bit more detailed presentation of these sets, by specifying explicitly the domain of each valuation.
@@
 Consider the sets
-[…]
+$CS(\Sigma)=\{\Lambda\subseteq For(\Sigma)\mid\mbox{ $\Lambda$ is non-empty and closed under subformulas}\},$
 […]
@@
 Let $\Lambda\in CS(\Sigma)$. A partial valuation in $\mathcal{M}^{\prime}_{S4}$ is a function $\tilde{v}_{p}:\Lambda\rightarrow\mathcal{V}$ such that, for every $\alpha,\beta\in\Lambda$:
-[…]
-[…]
+  • if $#\in\{\neg,\Box\}$ and $\alpha\in\Lambda$ then $\tilde{v}_{p}(\#\alpha)\in\#^{S4}(\tilde{v}_{p}(\alpha))$;
+  • if $#\in\{\to,\vee,\land\}$ and $\alpha\#\beta\in\Lambda$ then $\tilde{v}_{p}(\alpha\#\beta)\in\#^{\prime S4}(\tilde{v}_{p}(\alpha),\tilde{v}_{p}(\beta))$.
 Let $PV(\Lambda)$ be the set of partial valuations in $\mathcal{M}^{\prime}_{S4}$ with domain $\Lambda$.
@@
 Remark 6.9.
-Given $v,w$ and $\alpha$ let $\mathsf{P}_{\Lambda}(v,w,\alpha)$ iff $w(\alpha)=0$ and, $\forall\beta\in\Lambda$, $w(\beta)=2$ whenever $v(\beta)=2$. If $\Lambda\in FCS(\Sigma)$ then
-$\begin{array}[]{lll}PLV(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in PLV(\Lambda)$}\big{)}\}.\end{array}$
+Given $v,w$ and $\alpha$ let $\mathsf{P}^{4}_{\Lambda}(v,w,\alpha)$ iff $w(\alpha)=0$ and, $\forall\beta\in\Lambda$, $w(\beta)=2$ whenever $v(\beta)=2$. If $\Lambda\in FCS(\Sigma)$ then
+$\begin{array}[]{lll}PLV(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big(\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}^{4}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in PLV(\Lambda)$}\big)\}.\end{array}$
 As observed in Remark 5.6 for $\mathcal{M}_{S4}$, since $\Lambda$ is finite then $PV(\Lambda)$ is finite, hence the definition above is not cyclic.
+To be more precise, the following result can be proven:
+Proposition 6.10 (Existence and uniqueness of partial level valuations in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$).
+Let $\Lambda\in FCS(\Sigma)$. Then, there exists a unique set $PLV(\Lambda)\subseteq PV(\Lambda)$ satisfying the condition of Remark 6.9. Moreover, $PLV(\Lambda)\neq\emptyset$.
+Proof.
+The proof can be easily adapted from the corresponding one for Proposition 9.7 for $\mathbf{IPL}$ to be given in Section 9. ∎
 The next step is to show analyticity and co-analyticity w.r.t.
 level valuations in $\mathcal{M}^{\prime}_{S4}$.
-In order to do this it is considered, adapting [4, Definition 3.12], the intermediary notion of partial’ level valuation in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$.
+In order to do this it is considered, adapting [12, Definition 3.12], the intermediary notion of partial’ level valuation in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$.
 Definition 6.11 (Partial’ level valuation over $\Lambda$).
@@
 Remark 6.12.
-Consider the predicate $\mathsf{P}_{\Lambda}(v,w,\alpha)$ introduced in Remark 3.17. Clearly,
-$\begin{array}[]{lll}PLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $w\in\mathcal{L}^{\prime}_{S4}$}\big{)}\}.\end{array}$
+Consider the predicate $\mathsf{P}^{4}_{\Lambda}(v,w,\alpha)$ introduced in Remark 6.9. Clearly,
+$\begin{array}[]{lll}PLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big(\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}^{4}_{\Lambda}(\tilde{v}_{p},w,\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $w\in\mathcal{L}^{\prime}_{S4}$}\big)\}.\end{array}$
 Lemma 6.13 (Co-analyticity lemma for $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$).
@@
 Proof.
-It is an adaptation of the proof of [4, Lemma 3.13].
+It is an adaptation of the proof of [12, Lemma 3.13].
 Suppose, by contradiction, that $\tilde{v}_{p}\notin PLV^{\prime}(\Lambda)$.
@@
 By Theorem 6.5, $\{\Box\beta\mid\beta\in\Gamma\}\vdash_{S4}\alpha$.
-Observe that $\nvdash_{S4}\alpha$: otherwise, $v(\alpha)=2$, by Corollary  3.11.1.
+Observe that $\nvdash_{S4}\alpha$: otherwise, $v(\alpha)=2$, by Corollary 6.2.1.
 Hence, there exist $\beta_{1},\ldots,\beta_{n}\in\Gamma$ such that $\vdash_{S4}\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots))$.
@@
 Definition 6.14.
-[…]
+The complexity $\textsf{c}_{1}(\alpha)$ of a formula $\alpha\in For(\Sigma)$ is defined as follows: $\textsf{c}_{1}(p)=0$ if $p$ is a propositional variable; $\textsf{c}_{1}(\neg\alpha)=\textsf{c}_{1}(\Box\alpha)=\textsf{c}_{1}(\alpha)+1$; and $\textsf{c}_{1}(\alpha\#\beta)=\textsf{c}_{1}(\alpha)+\textsf{c}_{1}(\beta)+1$, for $#\in\{\to,\vee,\land\}$.
+Let $For(\Sigma)_{n}=\{\alpha\in For(\Sigma)\mid\textsf{c}_{1}(\alpha)\leq n\}$ for every $n\geq 0$.
 Remark 6.15.
 Observe that it is possible to define an enumeration $\alpha_{1},\alpha_{2},\ldots\alpha_{m},\ldots$ (for $m\in\omega^{2}$) of $For(\Sigma)$ such that $\textsf{c}_{1}(\alpha_{i})\leq\textsf{c}_{1}(\alpha_{j})$ if $i\leq j$ and, for every $i$ such that $\textsf{c}_{1}(\alpha_{i})>0$, if $\beta$ is a strict subformula of $\alpha_{i}$ then $\beta=\alpha_{j}$ for some $j<i$.
-Such an enumeration of $For(\Sigma)$ can be defined as follows: every formula have and index in $\omega^{2}$, which is a denumerable ordinal, such that all the formulas with complexity 00 (which form a denumerable set) are placed first, with indexes in $I_{0}:=\omega$; after this, all the formulas with complexity $1$ (which form a denumerable set) are placed with an index in $I_{1}:=\omega\cdot 2\setminus\omega=\{\omega,\omega+1,\ldots\}$; in general, the formulas with complexity $n$ (which form a denumerable set) have an index in $I_{n}:=\omega\cdot(n+1)\setminus\omega\cdot n=\{\omega\cdot n,\omega\cdot n+1,\ldots\}$.
+Such an enumeration of $For(\Sigma)$ can be defined as follows: every formula have and index in $\omega^{2}$, which is a denumerable ordinal, such that all the formulas with complexity 0 (which form a denumerable set) are placed first, with indexes in $I_{0}:=\omega$; after this, all the formulas with complexity $1$ (which form a denumerable set) are placed with an index in $I_{1}:=\omega\cdot 2\setminus\omega=\{\omega,\omega+1,\ldots\}$; in general, the formulas with complexity $n$ (which form a denumerable set) have an index in $I_{n}:=\omega\cdot(n+1)\setminus\omega\cdot n=\{\omega\cdot n,\omega\cdot n+1,\ldots\}$.
+Based on the enumeration of $For(\Sigma)$ given in the previous remark, consider the following sets:
+Definition 6.16.
+Let $\alpha_{1},\alpha_{2},\ldots\alpha_{m},\ldots$ (for $m\in\omega^{2}$) be an enumeration of $For(\Sigma)$ as in Remark 6.15.
+For every $n,m\in\omega$ let ${\Lambda}_{n}:=\Lambda\cup For(\Sigma)_{n}$; ${\Lambda}_{n}^{0}:={\Lambda}_{n}$; and ${\Lambda}_{n}^{m+1}={\Lambda}_{n}^{m}\cup\{\alpha_{\omega\cdot(n+1)+m}\}$.
+Remark 6.17.
+By definition, ${\Lambda}_{n}^{m+1}={\Lambda}_{n}\cup\{\alpha_{\omega\cdot(n+1)},\alpha_{\omega\cdot(n+1)+1},\ldots,\alpha_{\omega\cdot(n+1)+m}\}$, and $\alpha_{\omega\cdot(n+1)+m}\in\Lambda$ if and only if ${\Lambda}_{n}^{m+1}={\Lambda}_{n}^{m}$.
+Hence, ${\Lambda}_{n}^{m+1}$ is obtained from ${\Lambda}_{n}$ by adding the first $m+1$ formulas (of the given enumeration) with complexity $n+1$, and so ${\Lambda}_{n}^{m}$ adds at most $m$ formulas to ${\Lambda}_{n}$, for $m\in\omega$.
+It is worth noting that ${\Lambda}_{n}^{m}\in CS(\Sigma)$ for every $n,m\in\omega$; in particular, ${\Lambda}_{n}\in CS(\Sigma)$ for every $n\in\omega$.
+Clearly ${\Lambda}_{n}^{m}\subseteq{\Lambda}_{n}^{m+1}$, ${\Lambda}_{n+1}=\bigcup_{m\in\omega}{\Lambda}_{n}^{m}$, and $For(\Sigma)=\bigcup_{n\in\omega}{\Lambda}_{n}$.
 Lemma 6.18 (Analyticity lemma for $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$).
@@
 Proof.
-It is an adaptation of the proof of [4, Lemma 3.14].
+It is an adaptation of the proof of [12, Lemma 3.14].
 In this case, more modifications are required because of the conjunction and disjunction connectives, and by the fact of considering a fixed set $\Lambda$.
@@
 Then $\tilde{v}_{p}(\alpha)=2$ and so $\tilde{v}_{p}\in\mathcal{L}^{\prime S4}_{n+1}$.
-From this, $\tilde{v}_{p}\in\mathcal{L}^{\prime}_{S4}$ and the result clearly holds.
+From this, $\tilde{v}_{p}\in\mathcal{L}^{\prime}_{S4}$ and the result clearly holds by taking $v:=\tilde{v}_{p}$.
 Suppose now that $\Lambda\neq\emptyset$ is a proper subset of $For(\Sigma)$.
@@
 The valuation $v$ will be defined by induction on the complexity $n$ of the formulas.
-Observe that $\Lambda\cup\mathcal{P}\in CS(\Sigma)$ and, for every $n\in\omega$, $\Lambda_{n}\in CS(\Sigma)$ for $\Lambda_{n}:=\Lambda\cup For(\Sigma)_{n}$.
-In addition, for every $m\in\omega^{2}$ and $n\geq 0$, if $\alpha_{m}$ is such that $\textsf{c}_{1}(\alpha_{m})=n+1$, then $\Lambda_{n}^{m}\in CS(\Sigma)$, where $\Lambda_{n}^{m}=\Lambda_{n}\cup\{\alpha_{m}\}$.
-Clearly $\Lambda_{n+1}=\cup_{k\in\omega}\Lambda_{n}^{\omega\cdot(n+1)+k}$.
-Base $n=0$: Define $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v(p)=0$ if $p$ is a propositional variable which does not belong to $\Lambda$.
-Clearly, $v\in PLV^{\prime}(\Lambda\cup\mathcal{P})$ up to this stage.
-Inductive step: Assume that $v(\alpha)$ was defined for every $\alpha\in\Lambda_{n}$ such that $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v\in PLV^{\prime}(\Lambda_{n})$ up to this stage, for a given $n\geq 0$ (Induction Hypothesis, IH).
-Let $\alpha=\alpha_{m}$ (for $m\in I_{n+1}$) such that $\alpha\not\in\Lambda$ and $\textsf{c}_{1}(\alpha)=n+1$.
-Observe that $v(\beta)$ was already defined, for every strict subformula $\beta$ of $\alpha$.
-Since $v\in PLV^{\prime}(\Lambda_{n})$ then, for every $\delta\in\Lambda_{n}$ such that $v(\delta)=1$ there exists $w\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}_{\Lambda_{n}}(v,w,\delta)$.
-Such a $w$ will be denoted by $w^{\delta}$ (observe that it is possible to have more that one $w^{\delta}$ for each $\delta$, and it is possible to have $w^{\delta}=w^{\gamma}$ for $\delta\neq\gamma$).
-Recall that $\Lambda_{n}^{m}=\Lambda_{n}\cup\{\alpha_{m}\}$ is closed under subformulas.
-There are four cases to analyze:
-[…]
-[…]
-(2.1) There is $\delta\in\Lambda_{n}$ with $v(\delta)=1$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=0$.
-Define $v(\alpha)=1$.
-Hence $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\delta}$.
-(2.2) There is $\delta\in\Lambda_{n}$ with $v(\delta)=1$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=1$.
-By Lemma 3.20, $v^{\prime}=w^{\delta}_{|\Lambda_{n}^{m}}\in PLV^{\prime}(\Lambda_{n}^{m})$, hence there exists $w^{\prime\prime}\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}_{\Lambda_{n}^{m}}(v^{\prime},w^{\prime\prime},\alpha)$.
-Define $v(\alpha)=1$.
-Hence, $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\prime\prime}$.
-(2.3) For every $\delta\in\Lambda_{n}$ such that $v(\delta)=1$, $w^{\delta}(\alpha)=2$ for every $w^{\delta}$. Define $v(\alpha)=2$. Clearly, $v\in PLV^{\prime}(\Lambda_{n}^{m})$.
-[…]
+More precisely, with terminology as in Definition 6.16 and from the observations in Remark 6.17, an extension $v_{n}^{m}$ of $\tilde{v}_{p}$ (i.e., $\tilde{v}_{p}\subseteq v_{n}^{m}$) will be defined for every $(n,w)\in\omega\times\omega$ such that $v_{n}^{m}\in PLV^{\prime}({\Lambda}_{n}^{m})$ and $v_{i}^{j}\subseteq v_{n}^{m}$ if $(i,j)\leq(n,m)$, where: $(i,j)\leq(n,m)$ iff $i\leq n$ or $i=n$ and $j\leq m$ (and where $\omega\times\omega$ denotes, as usual, the Cartesian product of $\omega$ with itself).
+Base $n=m=0$: Observe that ${\Lambda}_{0}={\Lambda}_{0}^{0}=\Lambda\cup\mathcal{P}$.
+Define $v_{0}^{0}(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v_{0}^{0}(p)=0$ if $p$ is a propositional variable which does not belong to $\Lambda$.
+Clearly, $\tilde{v}_{p}\subseteq v_{0}^{0}$ and $v_{0}^{0}\in PLV^{\prime}({\Lambda}_{0})=PLV^{\prime}({\Lambda}_{0}^{0})$.
+Let $v_{0}:=v_{0}^{0}$.
+Inductive step: Assume that $v_{i}^{j}(\alpha)$ was defined for every $\alpha\in{\Lambda}_{i}^{j}$ such that $v_{i}^{j}(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, $v_{i}^{j}\in PLV^{\prime}({\Lambda}_{i}^{j})$, $v_{i}^{j}\subseteq v_{k}^{r}$ if $(i,j)\leq(k,r)\leq(n,m)$, for given $n,m\geq 0$ (Induction Hypothesis, IH).
+Now it will be shown how to extend $v_{n}^{m}$ to a function $v_{n}^{m+1}$ with domain ${\Lambda}_{n}^{m+1}={\Lambda}_{n}^{m}\cup\{\alpha_{\omega\cdot(n+1)+m}\}$.
+That is, it will be shown how to define a value for $\alpha_{\omega\cdot(n+1)+m}$ while preserving the values assigned by $v_{n}^{m}$, in such a manner that $v_{n}^{m+1}\in PLV^{\prime}({\Lambda}_{n}^{m+1})$.
+To start with, define $v_{n}^{m+1}(\beta)=v_{n}^{m}(\beta)$ for every $\beta\in{\Lambda}_{n}^{m}$.
+Let $\alpha=\alpha_{\omega\cdot(n+1)+m}$.
+If $\alpha\in\Lambda$ then ${\Lambda}_{n}^{m+1}={\Lambda}_{n}^{m}$.
+Then $v_{n}^{m+1}:=v_{n}^{m}$, and the required hypothesis are obviously satisfied by $v_{n}^{m+1}$.
+Suppose now that $\alpha\not\in\Lambda$.
+Observe that $v_{n}^{m+1}(\beta)=v_{n}^{m}(\beta)$ was already defined, for every strict subformula $\beta$ of $\alpha$.
+Since $v_{n}^{m}\in PLV^{\prime}({\Lambda}_{n}^{m})$ then, for every $\delta\in{\Lambda}_{n}^{m}$ such that $v_{n}^{m}(\delta)=1$ there exists $w\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}^{4}_{{\Lambda}_{n}^{m}}(v_{n}^{m},w,\delta)$.
+Such a $w$ will be denoted by $w^{\delta}$ (observe that it is possible to have more than one $w^{\delta}$ for each $\delta$, and it is possible to have $w^{\delta}=w^{\gamma}$ for $\delta\neq\gamma$).
+Recall that ${\Lambda}_{n}^{m}$ is closed under subformulas.
+There are three cases to analyze:
+(1) $\alpha=\#\beta$ such that $^{S4}v_{n}^{m}(\beta)=\{0\}$ (for some $#\in\{\neg,\Box\}$), or $\alpha=\beta\#\gamma$ such that $^{\prime S4}(v_{n}^{m}(\beta),v_{n}^{m}(\gamma))=\{0\}$ (for some $#\in\{\land,\vee,\to\}$).
+Then, define $v_{n}^{m+1}(\alpha)=0$.
+Observe that $v_{n}^{m+1}\in PLV^{\prime}({\Lambda}_{n}^{m+1})$.
+(2) $\alpha=\#\beta$ such that $^{S4}v_{n}^{m}(\beta)\in\mathcal{D}$ (for some $#\in\{\neg,\Box\}$), or $\alpha=\beta\#\gamma$ such that $^{\prime S4}(v_{n}^{m}(\beta),v_{n}^{m}(\gamma))\in\mathcal{D}$ (for some $\in\{\vee$$\to\}$).
+There are three subcases to analyze:
+(2.1) There is $\delta\in{\Lambda}_{n}^{m}$ with $v_{n}^{m}(\delta)=1$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=0$.
+Define $v_{n}^{m+1}(\alpha)=1$.
+Hence $v_{n}^{m+1}\in PLV^{\prime}(\Lambda_{n}^{m+1})$ such that $w^{\alpha}=w^{\delta}$.
+(2.2) There is $\delta\in\Lambda_{n}^{m}$ with $v_{n}^{m}(\delta)=1$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=1$.
+By Lemma 6.13, $v^{\prime}=w^{\delta}_{|\Lambda_{n}^{m}}\in PLV^{\prime}(\Lambda_{n}^{m})$, hence there exists $w^{\prime\prime}\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}^{4}_{\Lambda_{n}^{m}}(v^{\prime},w^{\prime\prime},\alpha)$.
+Define $v_{n}^{m+1}(\alpha)=1$.
+Hence, $v_{n}^{m+1}\in PLV^{\prime}(\Lambda_{n}^{m+1})$ such that $w^{\alpha}=w^{\prime\prime}$.
+(2.3) For every $\delta\in\Lambda_{n}^{m}$ such that $v_{n}^{m}(\delta)=1$, $w^{\delta}(\alpha)=2$ for every $w^{\delta}$.
+Define $v_{n}^{m+1}(\alpha)=2$.
+Clearly, $v_{n}^{m+1}\in PLV^{\prime}(\Lambda_{n}^{m+1})$.
+(3) $\alpha=\beta\#\gamma$ such that $^{\prime S4}(v_{n}^{m}(\beta),v_{n}^{m}(\gamma))=\{1\}$ (for some $\in\{\land$$\to\}$). There are two subcases to analyze:
 (3.1) $\alpha=\beta\to\gamma$.
-Then $v(\beta)=2$ and $v(\gamma)=1$.
-Define $v(\alpha)=1$.
-Then $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\gamma}$.
+Then $v_{n}^{m}(\beta)=2$ and $v_{n}^{m}(\gamma)=1$.
+Define $v_{n}^{m+1}(\alpha)=1$.
+Then $v_{n}^{m+1}\in PLV^{\prime}(\Lambda_{n}^{m+1})$ such that $w^{\alpha}=w^{\gamma}$.
 Indeed, since $w^{\gamma}(\beta)=2$ and $w^{\gamma}(\gamma)=0$ then $w^{\gamma}(\alpha)=0$.
 (3.2) $\alpha=\beta\land\gamma$.
-Then $v(\beta)=1$ and $v(\gamma)\in\mathcal{D}$ or vice versa.
+Then $v_{n}^{m}(\beta)=1$ and $v_{n}^{m}(\gamma)\in\mathcal{D}$ or vice versa.
 Then, there exists $w\in\mathcal{L}^{\prime}_{S4}$ (where $w=w^{\beta}$ or $w=w^{\gamma}$) such that $w(\beta)=0$ or $w(\gamma)=0$ and so $w(\alpha)=0$.
-Define $v(\alpha)=1$.
-Then $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\beta}$ or $w^{\alpha}=w^{\gamma}$.
-[…]
-We have shown in cases (1)-(4) how to extend the domain of $v$ to the additional formula $\alpha_{m}$, for every $\alpha_{m}$ such that $\textsf{c}_{1}(\alpha_{m})=n+1$, showing that $v\in PLV^{\prime}(\Lambda_{n}^{m})$.
-Consider now the simultaneous extension of $v$ with the values assigned to every $\alpha_{m}$ as prescribed above.
-We obtain so an extension of $v$ to $\Lambda_{n+1}=\cup_{m}\Lambda_{n}^{m}$.
-By the very definition of partial’ level valuations it follows that $v\in PLV^{\prime}(\Lambda_{n+1})$, and $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$.
-This completes the inductive step.
+Define $v_{n}^{m+1}(\alpha)=1$.
+Then $v_{n}^{m+1}\in PLV^{\prime}(\Lambda_{n}^{m+1})$ such that $w^{\alpha}=w^{\beta}$ or $w^{\alpha}=w^{\gamma}$.
+(4) $\alpha=\Box\beta$ such that $\Box^{S4}v(\beta)=\{2\}$ or $\alpha=\beta\#\gamma$ such that $^{\prime S4}(v_{n}^{m}(\beta),v_{n}^{m}(\gamma))=\{2\}$ (for some $#\in\{\vee,\land,\to\}$).
+Define $v_{n}^{m+1}(\alpha)=2$.
+It is clear that $v_{n}^{m+1}\in PLV^{\prime}(\Lambda_{n}^{m+1})$.
+We have shown in cases (1)-(4) how to extend the domain of $v_{n}^{m}$ to the additional formula $\alpha_{\omega\cdot(n+1)+m}$, showing that the resulting function $v_{n}^{m+1}$ is in $PLV^{\prime}(\Lambda_{n}^{m+1})$, and $\tilde{v}_{p}\subseteq v_{i}^{j}\subseteq v_{k}^{r}$, for every $(i,j)\leq(k,r)\leq(n,m+1)$.
+That is, we show how to increase the superscript $m$.
+In order to increase the subscript $n$, let $v_{n+1}=v_{n+1}^{0}:=\bigcup_{m\in\omega}v_{n}^{m}$.
+By the procedure described above, it is immediate to see that $v_{n+1}^{0}$ is in $PLV^{\prime}(\bigcup_{m\in\omega}{\Lambda}_{n}^{m})=PLV^{\prime}({\Lambda}_{n+1})=PLV^{\prime}({\Lambda}_{n+1}^{0})$, and $\tilde{v}_{p}\subseteq v_{i}^{j}\subseteq v_{k}^{r}$, for every $(i,j)\leq(k,r)\leq(n+1,0)$.
+By repeating the process, it is possible to define $v_{n}^{m}$ for every $(n,m)\in\omega\times\omega$ with the required properties.
+Finally, let $v:=\bigcup_{n\in\omega}v_{n}$.
+From the manner in which the construction was made, it follows that $v$ is in $PLV^{\prime}(\bigcup_{n\in\omega}{\Lambda}_{n})=PLV^{\prime}(For(\Sigma))$, and $\tilde{v}_{p}\subseteq v$.
+By the first part of the proof, we conclude that $v\in\mathcal{L}^{\prime}_{S4}$.
 ∎
@@
 Suppose that $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=1$.
-Then, there exists $w\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)$.
+Then, there exists $w\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}^{4}_{\Lambda}(\tilde{v}_{p},w,\alpha)$.
 Let $\tilde{w}_{p}=w_{|\Lambda}$.
-By Lemma 3.20, $\tilde{w}_{p}\in PLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)$.
-Now, suppose that there exists $\tilde{w}^{\prime}_{p}\in PLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}^{\prime}_{p},\alpha)$.
+By Lemma 6.13, $\tilde{w}_{p}\in PLV^{\prime}(\Lambda)$ such that $\mathsf{P}^{4}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)$.
+Now, suppose that there exists $\tilde{w}^{\prime}_{p}\in PLV^{\prime}(\Lambda)$ such that $\mathsf{P}^{4}_{\Lambda}(\tilde{v}_{p},\tilde{w}^{\prime}_{p},\alpha)$.
 By Lemma 6.18, there exists $v^{\prime}\in\mathcal{L}^{\prime}_{S4}$ such that the restriction of $v^{\prime}$ to $\Lambda$ coincides with $\tilde{w}^{\prime}_{p}$.
-Hence, $\mathsf{P}_{\Lambda}(\tilde{v}_{p},v^{\prime},\alpha)$.
+Hence, $\mathsf{P}^{4}_{\Lambda}(\tilde{v}_{p},v^{\prime},\alpha)$.
 This shows that
-$\begin{array}[]{lll}PLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in PLV^{\prime}(\Lambda)$}\big{)}\}.\end{array}$ From this, being $PV(\Lambda)$ finite, it follows that $PLV(\Lambda)=PLV^{\prime}(\Lambda)$.
+$\begin{array}[]{lll}PLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big(\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}^{4}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in PLV^{\prime}(\Lambda)$}\big)\}.\end{array}$ From this, it follows that $PLV(\Lambda)=PLV^{\prime}(\Lambda)$, by Proposition 6.10.
 ∎
@@
 This means that Grätz’s algorithm can be adapted to the reduced (and expanded with disjunction and conjunction) RNmatrix $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$.
-4 An RNmatrix for IPL
+7 First steps towards an RNmatrix for $\mathbf{IPL}$
+As mentioned in Section 1, Kearns’s level valuations fail to provide an effective decision procedure for S4, while Grätz method produces an algorithm to decide that logic.
+Clearly, the composition of Gödel’s translation from $\mathbf{IPL}$ to S4 with Grätz’s decision procedure for S4 naturally produces a decision procedure for $\mathbf{IPL}$.
+Let $\Sigma=(\{\neg,\Box\},\{\land,\vee,\to\})$ be the full signature of S4, and let $\Omega=(\{\neg\},\{\land,\lor,\rightarrow\})$ be the signature for intuitionistic propositional logic $\mathbf{IPL}$.
 Definition 7.1.
-Consider the signature Ω={{¬},{,,}}normal-$\Omega=\{\{\neg\},\{\land,\lor,\rightarrow\}\}$ for intuitionistic propositional logic (IPL).
-The set of intuitionistic formulas will be denoted by as $For(\Omega)$.
-4.1 Gödel-McKinsey-Tarski Box Translation
+The complexity $\textsf{c}(\alpha)$ of a formula $\alpha\in For(\Omega)$ is defined as follows: $\textsf{c}(p)=0$ if $p$ is a propositional variable; $\textsf{c}(\neg\alpha)=\textsf{c}(\alpha)+1$; and $\textsf{c}(\alpha\#\beta)=\textsf{c}(\alpha)+\textsf{c}(\beta)+1$, for $#\in\{\to,\vee,\land\}$.
+Recall the Gödel-Tarski-McKinsey Box translation $t:For(\Omega)\rightarrow For(\Sigma)$ defined as follows:
+  1. i)
+    $t(p)=\Box p$, if $p$ is a propositional variable;
+  2. ii)
+    $t(\neg\alpha)=\Box\neg t(\alpha)$;
+  3. iii)
+    $t(\alpha\#\beta)=\Box(t(\alpha)\#t(\beta))$, where $#\in\{\rightarrow,\lor,\land\}$.
+It is widely known that this translation is faithful (see [25]; for a more recent proof see [26]), that is:
+Theorem 7.2 ([25]).
+Let $\varphi$ be a formula in $For(\Omega)$. Then: $\varphi$ is valid in $\mathbf{IPL}$ if and only if $t(\varphi)$ is valid in S4.
+As seen in Section 6, Grätz’s results and algorithm can be extended to S4 defined over $\Sigma$, now based on $\mathcal{M}^{\prime}_{S4}$.
+Then, by Theorem 7.2 and by Theorem 6.20, a formula $\alpha$ in $For(\Omega)$ is valid in $\mathbf{IPL}$ iff the reduced truth-table obtained by taking the partial valuations in $\mathcal{M}^{\prime}_{S4}$ for the set of subformulas of the formula $t(\alpha)$ in $For(\Sigma)$ shows that $t(\alpha)$ is a valid formula in S4.
+This produces obviously a decision procedure for $\mathbf{IPL}$, based on an RNmatrix for S4 and a translation between $\mathbf{IPL}$ and S4.
+It is a challenging question how to define a direct decision procedure for $\mathbf{IPL}$ based on a finite-valued RNmatrix for $\mathbf{IPL}$, with level valuations (and with an associate truth-tables decision procedure) specifically designed for this logic.
+Moreover, the soundness and completeness of the RNmatrix (and the decision procedure obtained from it) should not rely on the corresponding soundness and completeness results of Gödel’s [11] and Grätz’s constructions.
+A solution to this question will be given in the next two sections.
+8 An RNmatrix for $\mathbf{IPL}$
+This section describes a new semantics for $\mathbf{IPL}$ based on a 3-valued RNmatrix with a very intuitive interpretation.
+We obtain this semantics by abstracting the composed procedure for $\mathbf{IPL}$ mentioned at the end of the previous section but defined in a direct way.
+That is, we describe a ‘pure’ RNmatrix for $\mathbf{IPL}$ (i.e., one that does not depend on any other construction), which turns out to be a suitable semantics for $\mathbf{IPL}$.
+Moreover, Section 9 presents an algorithm for deciding validity in $\mathbf{IPL}$ by removing spurious rows from the truth-tables generated by the procedure.
+Hence, the soundness and completeness of the semantics and the associated decision procedure are proven independently of Gödel’s 1933 results (proved in [25]) and those in [12].
+8.1 Inducing the Nmatrix $\mathcal{M}_{IPL}$ for $\mathbf{IPL}$
+The truth-tables for intuitionistic logic can be defined based on the Box translation $t$ described above, combined with a semi-translation $t_{0}$ to be defined now.
 Definition 8.1.
-(Box translation function) Let $f_{\Box}:For(\Omega)\rightarrow For(\Sigma)$ be a function such that:
+The semi-translation function $t_{0}:For(\Omega)\rightarrow For(\Sigma)$ is defined as follows:
   1. i)
-    $f_{\Box}(p)=\Box p$, if $p$ is a propositional variable;
+    $t_{0}(p)=p$, if $p$ is a propositional variable;
   2. ii)
-    $f_{\Box}(\neg\alpha)=\Box\neg f_{\Box}(\alpha)$;
+    $t_{0}(\neg\alpha)=\neg t(\alpha)$;
   3. iii)
-[…]
-It is widely known that this translation is faithful [9].
-4.2 Truth tables for intuitionistic logic
-We introduce the truth tables for intuitionistic logic based on the presented translation.
-4.2.1 Intuitionistic non-deterministic valuations
+    $t_{0}(\alpha\#\beta)=t(\alpha)\#t(\beta)$, where $#\in\{\rightarrow,\lor,\land\}$.
+Lemma 8.2.
+For every $\alpha\in For(\Omega)$, $t(\alpha)=\Box(t_{0}(\alpha))$.
+Proof.
+Straightforward, by induction on the complexity of $\alpha$. ∎
 Remark 8.3.
-If $v$ is a valuation in $\mathcal{M}_{S4}$, then, for every $\alpha\in For(\Sigma)$, $v(\Box\alpha)=\Box^{\mathcal{A}_{S4}}(v(\alpha))$.
-This follows from the fact that the multioperation $\Box^{\mathcal{A}_{S4}}$ is deterministic.
+Observe that the interpretation of $\Box$ in $\mathcal{M}^{\prime}_{S4}$ is deterministic (see Table 2).
+Let us denote by $\Box^{4}(a)$ the only element of ${\Box^{\prime}}^{S4}(a)$, for $a\in\mathcal{V}$.
+Hence, if $v$ is a valuation in $\mathcal{M}^{\prime}_{S4}$ then $v(\Box\alpha)=\Box^{4}(v(\alpha))$, for every $\alpha\in For(\Sigma)$.
+By Lemma 8.2, $v(t(\alpha))=\Box^{4}(v(t_{0}(\alpha)))$, for every $\alpha\in For(\Omega)$.
+To check the validity in $\mathbf{IPL}$ of $\alpha\in For(\Omega)$, one must construct a truth-table in $\mathcal{M}^{\prime}_{S4}$ for $t(\alpha)\in For(\Sigma)$.
+Assume that the set of subformulas of $\alpha$ is arranged as a sequence $\alpha_{1},\ldots,\alpha_{n}=\alpha$ in non-decreasing order of complexity (the complexity of formulas in $For(\Omega)$ is defined as usual, see Definition 7.1 above).
+But then, the columns of the truth-table for $t(\alpha)$ in $\mathcal{M}^{\prime}_{S4}$ can be organized in pairs of formulas of the form $\langle t_{0}(\alpha_{1});t(\alpha_{1})\rangle,\ldots,\langle t_{0}(\alpha_{n});t(\alpha_{n})\rangle=\langle t_{0}(\alpha);t(\alpha)\rangle$.
+Indeed, the sequence of formulas $t_{0}(\alpha_{1}),t(\alpha_{1}),t_{0}(\alpha_{2}),\ldots,t_{0}(\alpha_{n}),t(\alpha_{n})=t(\alpha)$ in $For(\Sigma)$ induced by the sequence of pairs of formulas is the sequence of formulas that can be considered to construct a truth-table for $t(\alpha)$.
+[^5] Rigorously speaking, if $\alpha$ depends on the propositional variables $p_{1},\ldots,p_{k}$ then usually it is considered $\alpha_{i}=p_{i}$ (for $1\leq i\leq k$), and then $t(\alpha_{i})=\Box p_{i}$ is placed at a latter stage, say $\alpha_{i+k}$.
+Hence, the first $2k$ columns of a ‘standard’ table for $t(\alpha)$ starts with formulas $p_{1},\ldots,p_{k},\Box p_{1},\ldots\Box p_{k}$ instead of $p_{1},\Box p_{1},p_{2},\Box p_{2},\ldots p_{k},\Box p_{k}$.
+After this, the pairs produce the final columns of a table for $t(\alpha)$.
+For instance, in order to check $\alpha=p\vee\neg p$ it is considered the sequence of pairs $\langle p;\Box p\rangle$, $\langle\neg\Box p;\Box\neg\Box p\rangle$, $\langle\Box p\lor\Box\neg\Box p;\Box(\Box p\lor\Box\neg\Box p)\rangle$ based on the sequence $p,\neg p,p\vee\neg p$ for $\alpha$.
+Observe that $t(\alpha)=\Box(\Box p\lor\Box\neg\Box p)$.
+This suggests considering for $\mathbf{IPL}$ an Nmatrix (derived from $\mathcal{M}^{\prime}_{S4}$) with domain $\left\{\langle c;\Box^{4}(c)\rangle\mid c\in\mathcal{V}\right\}=\{\langle 2;2\rangle,\langle 1;0\rangle,\langle 0;0\rangle\}$ and multioperations defined as follows:
+  1. 1.
+    $\neg^{IPL}(\langle a;b\rangle)=\left\{\langle c;\Box^{4}(c)\rangle\mid c\in{\neg^{\prime}}^{S4}(b)\right\}$
+  2. 2.
+    $^{IPL}(\langle a;b\rangle,\langle c;d\rangle)=\left\{\langle e;\Box^{4}(e)\rangle\mid e\in{\#^{\prime}}^{S4}(b,d)\right\}$
+where $#\in\{\rightarrow,\lor,\land\}$.
+Let us now consider a label for the truth-values above as follows: $\mathbf{T}:=\langle 2;2\rangle$, $\mathbf{U}:=\langle 1;0\rangle$ and $\mathbf{F}:=\langle 0;0\rangle$.
+The observations above lead us to the following:
+Definition 8.4 (Intuitionistic Nmatrix $\mathcal{M}_{IPL}$).
+The intuitionistic Nmatrix over signature $\Omega$ is given by $\mathcal{M}_{IPL}=\langle V_{IPL},\mathcal{D}_{IPL},\mathcal{O}_{IPL}\rangle$, where $V_{IPL}=\{\mathbf{T},\mathbf{U},\mathbf{F}\}$, $\mathcal{D}_{IPL}=\{\mathbf{T}\}$, and $\mathcal{O}_{IPL}(\#)=\#^{IPL}$ for each connective #\# in $\Omega$ is defined as in Figure 3.
+Figure 3: Tables for $\mathcal{M}_{IPL}$.
+Observe that, in a given valuation over $\mathcal{M}_{IPL}$, a formula is false when receives a non-designated value, $\mathbf{U}$ or $\mathbf{F}$.
+By looking at the tables of the Nmatrix, apparently both values have the same behavior.
+But it is not the case.
+The first thing to note in Figure 3 is that the truth-value $\mathbf{U}$, which intuitively means that the epistemic status of $\alpha$ is unknown, only appears in the definitions of $\neg^{IPL}$ and $\to^{IPL}$.
+As it is going to be more detailed in the completeness proof, the intended meaning is the following (see also Remark 8.15 below):
+$v(\neg\alpha)=\mathbf{U}$
+only happens when both $\neg\alpha$ and $\alpha$ are false; that is, when there is no proof of (or knowledge or information about) neither $\alpha$ nor $\neg\alpha$, hence excluded middle fails for $\alpha$ (intuitionistic, non-classical scenario);
+$v(\neg\alpha)=\mathbf{F}$
+means that $\neg\alpha$ is false, but $\alpha$ is true; that is, although $\neg\alpha$ is unknown, $\alpha$ is known, therefore excluded middle holds for $\alpha$ (classical scenario);
+$v(\alpha\to\beta)=\mathbf{U}$
+only happens when both $\alpha$ and $\alpha\to\beta$ are false; that is, there is no proof of (or knowledge or information about) neither $\alpha$ nor $\alpha\to\beta$, hence Peirce’s law $\alpha\vee(\alpha\to\beta)$ fails for $\alpha$ and $\alpha\to\beta$ (intuitionistic, non-classical scenario);
+$v(\alpha\to\beta)=\mathbf{F}$
+means that $\alpha\to\beta$ is false, but $\alpha$ is true; that is, although $\alpha\to\beta$ is unknown, $\alpha$ is known, therefore Peirce’s law holds for $\alpha$ and $\alpha\to\beta$ (classical scenario).
+In turn, the behavior of conjunction and disjunction is ‘classical’ (in terms of designated/non-designated values), and that is to be expected: the inference rules for conjunction and disjunction are identical in classical and intuitionistic logic.
+Because of this, $\mathbf{U}$ never appears as a value assigned to a conjunction or a disjunction.
+The third point to note is that, if we disregard the truth-value $\mathbf{U}$, the remaining Nmatrix corresponds exactly to classical logic.
+In fact, it is straightforward to see how classical logic can be fully recovered from the intuitionistic Nmatrix.
+This recovery can be achieved by refining $\mathcal{M}_{IPL}$ (in the sense of [3] and [4]) to the truth-values $\mathbf{T},\mathbf{F}$}.
+The resulting is an Nmatrix with zero degree of non-determinism, equivalent to the standard (deterministic) matrix for classical logic.
+8.2 Level valuations for $\mathcal{M}_{IPL}$
+The following step is the definition of a suitable set of valuations (level valuations) over the Nmatrix for $\mathbf{IPL}$ just defined.
+The construction of the sets of level valuations given by Kearns and Grätz is based on the idea of super designated truth-values, that is: designated values $a$ such that any element of $\tilde{\Box}(a)$ is still designated, where $\tilde{\Box}$ is the multioperator associated to $\Box$.
+Thus, if a formula gets a designated value for every $n$th-level valuation, it must receive a super-designated value in every $(n+1)$th-level valuation.
+This guarantees the validity of the Necessitation inference rule, level by level.
+In this way, if a formula is provable in S4 with $n$ applications of the necessitation rule, then it is valid in $\mathcal{L}^{S4}_{n}$.
+From this, soundness of syntactical proofs in S4 w.r.t.
+the RNmatrix semantics follows easily.
+Necessitation is usually called a global inference rule: in order to be sound, the hypothesis of the rule must be provable, not merely assumed.
+However, $\mathbf{IPL}$ has no global inference rules in its usual Hilbert-style presentation (being Modus Ponens the only inference rule, which is local in opposition to global).
+Taking into account that we are abstracting the composition of the Box translation with Grätz’s truth-tables algorithm for S4, the set of $n$th-level valuations will be constructed based on partial level valuations (Definition 5.5) instead of the Kearns-like Definition 5.1 of level valuations.
+Thus, the requirement that every valuation (row) $v_{p}$ such $v_{p}(\alpha)=1$ must be supported by another valuation (row) $w_{p}$ with $w_{p}(\alpha)=0$ while preserving the values $2$ can be easily adapted to the present framework, by changing $1$, 0 and $2$ by $\mathbf{U}$, $\mathbf{F}$ and $\mathbf{T}$, respectively (see Definition 8.6 below).
+However, as observed above, there is no counterpart to Necessitation rule in the usual Hilbert-style calculus for $\mathbf{IPL}$, therefore the level valuations seem, in principle, unrelated to syntactical proofs in $\mathbf{IPL}$.
+The key is to consider a standard sequent calculus presentation for $\mathbf{IPL}$, and so it can be established, in a similar way to the case of S4, a relationship between syntactical proofs in $\mathbf{IPL}$ and nthnth-level valuations: if a sequent is proved in $\mathbf{IPL}$ with a derivation of height $n$ then it is valid w.r.t.
+nthnth-level valuations for $\mathbf{IPL}$ (see Proposition 8.11).
+Using this, the soundness of the sequent calculus for $\mathbf{IPL}$ w.r.t.
+its RNmatrix semantics can be easily obtained.
+In addition, intuitionistic level valuations have a very interesting interpretation.
+Indeed, a suitable and very natural $3$-valued characteristic function associated to $\varphi$-saturated sets in $\mathbf{IPL}$ will be introduced in Definition 8.14.
+Hence, in Proposition 8.17 it will be proven that these characteristic functions correspond to level valuations.
+The relationship between level valuations and $3$-valued characteristic functions of $\varphi$-saturated sets also has an interesting consequence in terms of complexity: at level 0 (and hence at any level) the valuations can only assign the classical truth-values $\mathbf{T}$ and $\mathbf{F}$ to propositional variables.
+Indeed, atomic formulas contain neither negation nor implication connectives, so it makes no sense to assign them the value $\mathbf{U}$, which, as discussed above, is exclusively related to the failure of the law of excluded middle and Peirce’s law.
+This feature will reduce the complexity of the algorithm, as we will be discuss in Section 10.
 Definition 8.5.
-We define the function $h:For(\Omega)\rightarrow For(\Sigma)$ as follows:
+Let $Val^{nd}(\mathcal{M}_{IPL})$ be the set of valuations over $\mathcal{M}_{IPL}$, and let $Val^{nd}_{+}(\mathcal{M}_{IPL})$ be the set of valuations $v$ in $Val^{nd}(\mathcal{M}_{IPL})$ such that $v(p)\in\{\mathbf{T},\mathbf{F}\}$ for every propositional variable $p$.
+Definition 8.6.
+(Level valuations for $\mathbf{IPL}$) The set of $n$th-level valuations $\mathcal{L}^{IPL}_{n}$ for $\mathbf{IPL}$, where $n\in\omega$, is defined as follows:
   1. i)
-    $h(p)=p$, if $p$ is a propositional variable;
+    $\mathcal{L}_{0}^{IPL}=Val^{nd}_{+}(\mathcal{M}_{IPL})$;
   2. ii)
-    $h(\neg\alpha)=\neg f_{\Box}(\alpha)$;
-  3. iii)
-[…]
-We call $h$ as the semi-translation function.
-Lemma 4.5.
-For every $\alpha\in For(\Omega)$, $f_{\Box}(\alpha)=\Box(h(\alpha))$.
-Proof.
-Let $\alpha$ be an intuitionistic formula. We proceed by case analysis.
-Case $\alpha=p_{n}$
-In this case, $f_{\Box}(\alpha)=f_{\Box}(p_{n})=\Box p_{m}$. Also, $\Box(h(\alpha))=\Box(h(p_{n}))=\Box p_{m}$. Note that $n=m$ in both cases. Hence, $f_{\Box}(p_{n})=\Box(h(p_{n}))$.
-Case $\alpha=\neg\beta$
-By Definition 4.4, $\Box(h(\neg\beta))=\Box(\neg f_{\Box}(\beta))$. Moreover, by Definition 4.2, $f_{\Box}(\neg\beta)=\Box(\neg f_{\Box}(\beta))$. Hence, $f_{\Box}(\neg\beta)=\Box(h(\neg\beta))$.
-[…]
-By Definition 4.4, we have that for each ##\##, $\Box(h(\beta\#\gamma))=\Box(f_{\Box}(\beta)\#f_{\Box}(\gamma))$.
-Moreover, by Definition 4.2, $f_{\Box}(\beta\#\gamma)=\Box(f_{\Box}(\beta)\#f_{\Box}(\gamma))$.
-Therefore, $f_{\Box}(\beta\#\gamma)=\Box(h(\beta\#\gamma))$.
-Remark 4.6.
-By Lemma 4.5 and Remark 4.3, $v(f_{\Box}(\alpha))=\Box^{\mathcal{A}_{S4}}(v(h(\alpha)))$, for every $\alpha\in For(\Omega)$.
-Definition 4.7.
-(Truth function) Let $v$ be a valuation function in $\mathcal{M}_{S4}$ and $\alpha\in For(\Omega)$. We define the function $\tilde{v}:For(\Omega)\rightarrow V\times V$ as follows:
-$\tilde{v}(\alpha)\triangleq\langle v(h(\alpha));v(f_{\Box}(\alpha))\rangle$
-We will say that $\tilde{v}$ is induced by $v$ and will call $\tilde{v}$ as the truth function.
-By Remark 4.6, we get $\tilde{v}(\alpha)=\langle v(h(\alpha));\Box^{\mathcal{A}_{S4}}(v(h(\alpha)))\rangle$. This motivates the following:
-Definition 4.8 (Intuitionistic truth values).
-The set of intuitionistic truth values is given by
-$V_{IPL}=\left\{\langle c;\Box^{\mathcal{A}_{S4}}(c)\rangle\mid c\in\mathcal{V}\right\}=\{\mathbf{F},\mathbf{U},\mathbf{T}\}$
-where $\mathbf{F}=\langle 0;0\rangle$, $\mathbf{U}=\langle 1;0\rangle$ and $\mathbf{T}=\langle 2;2\rangle$.
-Therefore, the truth function 4.7 is a function $\tilde{v}:For(\Omega)\rightarrow\{\mathbf{F},\mathbf{U},\mathbf{T}\}$.
-[…]
-Definition 4.9 (Intuitionistic multioperations).
-Let $\langle a,b\rangle,\langle c,d\rangle\in V_{IPL}$. Define the intuitionistic multioperations as follows:
-  1. 1.
-    $\neg^{\prime\mathcal{A}_{IPL}}(\langle a;b\rangle)=\left\{\langle c;\Box^{\mathcal{A}_{S4}}(c)\rangle\mid c\in\neg^{\mathcal{A}_{S4}}(b)\right\}$
-  2. 2.
-    #$\#^{\prime\mathcal{A}_{IPL}}(\langle a;b\rangle,\langle c;d\rangle)=\left\{\langle e;\Box^{\mathcal{A}_{S4}}(e)\rangle\mid e\in\#^{\mathcal{A}^{\prime}_{S4}}(b,d)\right\}$
-[…]
-Definition 4.10.
-The intuitionistic Nmatrix over signature $\Omega$ is given by $\mathcal{M}_{IPL}=\langle V_{IPL},\mathcal{D}_{IPL},\mathcal{O}_{IPL}\rangle$, where $\mathcal{D}_{IPL}=\{\mathbf{T}\}$, $\mathcal{O}_{IPL}=\{\neg^{\mathcal{A}_{IPL}},\rightarrow^{\mathcal{A}_{IPL}},\lor^{\mathcal{A}_{IPL}},\land^{\mathcal{A}_{IPL}}\}$ and each $o\in\mathcal{O}_{IPL}$ is defined in Table 3.
-$\neg^{\mathcal{A}_{IPL}}$
+    $\mathcal{L}^{IPL}_{n+1}=\{v\in\mathcal{L}^{IPL}_{n}\mid\forall\alpha\in For(\Omega)$, if $v(\alpha)=\mathbf{U}$ then there exists $w\in\mathcal{L}^{IPL}_{n}$ such that $w(\alpha)=\mathbf{F}$ and, for every $\beta\in For(\Omega)$: $v(\beta)=\mathbf{T}$ implies $w(\beta)=\mathbf{T}\}$.
+The set of level valuations in $\mathcal{M}_{IPL}$ is defined as the intersection of the sets of $n$th-level valuations, for $n\geq 0$:
+$\mathcal{L}_{IPL}=\bigcap^{\infty}_{n\geq 0}\mathcal{L}^{IPL}_{n}$
+Definition 8.7 (RNmatrix for IPL).
+Let $\mathcal{M}_{IPL}$ be the Nmatrix defined in Definition 8.4. The RNmatrix for IPL is the pair $\mathcal{R}(\mathcal{M}_{IPL})=\langle\mathcal{M}_{IPL},\mathcal{L}_{IPL}\rangle$.
+8.3 Soundness and Completeness of $\mathbf{IPL}$ w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$
+From now on, LI will stand for a standard sequent calculus for IPL over $\Omega$.
+In fact, the sequent calculus LI presented here is obtained from Gentzen’s system LJ (see, for instance, [30, Chapter 1]) with minor modifications.
+Concretely, we consider sequents formed by pairs of finite multisets of formulas instead of pairs of sequences (which allows to eliminate the exchange rule), and by adding contexts to the axiom (which allows to eliminate the left weakening rule).
+Certain rules can also be simplified in this setting.
+Clearly, both calculi are equivalent.
+By a sequent we mean an expression $\Gamma\Rightarrow\Delta$ such that $\Gamma$ and $\Delta$ are (non simultaneously empty) finite multisets, and $\Delta$ contains at most one formula.
+The calculus LI is given by the following axiom and inference rules:
+Axiom
+$(Ax)\,\,\alpha,\Gamma\Rightarrow\alpha$
+Structural Rules
+$(contr\Rightarrow)\,\displaystyle\frac{\alpha,\alpha,\Gamma\Rightarrow\Delta}{\alpha,\Gamma\Rightarrow\Delta}\hskip 56.9055pt(\Rightarrow w)\,\displaystyle\frac{\Gamma\Rightarrow}{\Gamma\Rightarrow\alpha}$
+$(cut)\,\displaystyle\frac{\Gamma\Rightarrow\alpha\hskip 14.22636pt\alpha,\Gamma\Rightarrow\Delta}{\Gamma\Rightarrow\Delta}$
+Logical Rules
 […]
@@
 […]
-$\rightarrow^{\mathcal{A}_{IPL}}$ F U T
-[…]
-[…]
-[…]
-$\lor^{\mathcal{A}_{IPL}}$ F U T
-[…]
-[…]
-[…]
-$\land^{\mathcal{A}_{IPL}}$ F U T
-[…]
-[…]
-[…]
-Table 3: Tables for #$\#^{\mathcal{A}_{IPL}}$.
-Lemma 4.11.
-Let #$\#\in\Omega$. Then, #$\#^{\mathcal{A}_{IPL}}=\#^{\prime\mathcal{A}_{IPL}}$. That is, the multioperators in Definition 4.9 coincide with the ones in Definition 4.10.
+$\mbox{($\neg\Rightarrow$)}\,\,\displaystyle\frac{\Gamma\Rightarrow\alpha}{\neg\alpha,\Gamma\Rightarrow}\hskip 42.67912pt\mbox{($\Rightarrow\neg$)}\,\,\displaystyle\frac{\alpha,\Gamma\Rightarrow}{\Gamma\Rightarrow\neg\alpha}$
+The consequence relation of LI will be denoted by $\vdash_{IPL}$.
+Hence, for any set $\Gamma\cup\{\alpha\}\subseteq For(\Omega)$, $\Gamma\vdash_{IPL}\alpha$ iff there exists a finite set $\Gamma_{0}\subseteq\Gamma$ such that the sequent $\Gamma_{0}\Rightarrow\alpha$ is derivable in LI.
+The logic generated by LI, with consequence relation $\vdash_{IPL}$, is precisely $\mathbf{IPL}$.
+Remark 8.8.
+Being a presentation of $\mathbf{IPL}$, the deduction metatheorem (DM) holds for the consequence relation $\vdash_{IPL}$ associated to LI:
+$(DM)\ \ \Gamma,\alpha\vdash_{IPL}\beta\ \ \mbox{ iff }\ \ \Gamma\vdash_{IPL}\alpha\to\beta$
+for every set of formulas $\Gamma\cup\{\alpha,\beta\}$.
+The validity of (DM) can be checked directly by using ($Ax$) and the rules for implication.
+Indeed, the ‘only if’ part follows by ($\Rightarrow\to$).
+For the ‘if’ part, it can be observed that $\Gamma_{0},\alpha,\alpha\to\beta\Rightarrow\beta$ is provable in LI for every finite multiset $\Gamma_{0}$ and formulas $\alpha$ and $\beta$: this is a consequence of applying ($\to\Rightarrow$) to the axioms $\Gamma_{0},\alpha\Rightarrow\alpha$ and $\Gamma_{0},\alpha,\beta\Rightarrow\beta$.
+From this, the result follows from the hypothesis $\Gamma_{0}\Rightarrow\alpha\to\beta$ (for a finite set $\Gamma_{0}\subseteq\Gamma$) and (cutcut).
+Definition 8.9.
+Let $\Pi$ be a derivation of a sequent $\Gamma\Rightarrow\Delta$ in LI (that is, $\Pi$ is a tree of sequents with root $\Gamma\Rightarrow\Delta$ and axioms as leaves).
+The depth of $\Pi$ is its height (as a tree).
+If a sequent $\Gamma\Rightarrow\Delta$ admits a derivation in LI of depth $k$ (for $k\geq 0$), we write $\Gamma\vdash_{IPL}^{k}\Delta$.
+The proof of soundness of LI w.r.t. the RNmatrix $\mathcal{R}(\mathcal{M}_{IPL})$ requires to establish, in a precise way, soundness of derivations of depth $k$ w.r.t. valuations of level $k$.
+Definition 8.10.
+Let $k\geq 0$ and $v\in\mathcal{L}_{k}^{IPL}$.
+We say that $v$ satisfies a formula $\alpha$, written as $v\models\alpha$, if $v(\alpha)=\mathbf{T}$.
+If $\Gamma$ is a non-empty (multi)set of formulas, we say that $v$ satisfies $\Gamma$, denoted by $v\models\Gamma$, if $v\models\gamma$ for every $\gamma\in\Gamma$.
+We say that $v$ satisfies a sequent $\Gamma\Rightarrow\Delta$, denoted by $v\models\Gamma\Rightarrow\Delta$, if the following holds:
+(1) $\Delta=\emptyset$. In this case, $v\not\models\Gamma$, that is, $v(\beta)\in\{\mathbf{U},\mathbf{F}\}$ for some $\beta\in\Gamma$; or
+(2) $\Delta=\{\gamma\}$, for some formula $\gamma$. In this case, $v\models\gamma$ whenever $v\models\Gamma$.
+A sequent $\Gamma\Rightarrow\Delta$ is said to be valid in $\mathcal{L}_{k}^{IPL}$, denoted by $\Gamma\models_{k}\Delta$, if it is satisfied by any $v\in\mathcal{L}_{k}^{IPL}$.
+Proposition 8.11.
+For every $k\geq 0$ and every sequent $\Gamma\Rightarrow\Delta$: if $\Gamma\vdash_{IPL}^{k}\Delta$, then $\Gamma\models_{k}\Delta$.
 Proof.
-It is straightforward to verify that, for every $z,w\in V_{IPL}$:
-  1. 1.
-    $\neg^{\mathcal{A}_{IPL}}(z)=\neg^{\prime\mathcal{A}_{IPL}}(z)$; and
-  2. 2.
-[…]
-Definition 4.12.
-[…]
-Definition 4.13.
-Given $n\geq 0$, the set of formulas over $\Omega$ with complexity less or equal than $n$ will be denoted by $For(\Omega)_{n}$.
-That is: $For(\Omega)_{n}=\{\alpha\in For(\Omega)\mid\textsf{c}(\alpha)\leq n\}$.
-Definition 4.14.
-Let $Val(\mathcal{M}_{IPL})$ be the set of valuations over $\mathcal{M}_{IPL}$, and let $Val_{+}(\mathcal{M}_{IPL})$ be the set of valuations $v$ in $Val(\mathcal{M}_{IPL})$ such that $v(p)\in\{\mathbf{T},\mathbf{F}\}$ for every propositional variable $p$.
-Definition 4.15.
-(Level valuations for IPL) Let $Val(\mathcal{M}_{IPL})$ be the set of valuations over $\mathcal{M}_{IPL}$. We define the level $\mathcal{L}^{IPL}_{k}$, where $k\in\mathbb{N}$, as follows:
-  1. i)
-    $\mathcal{L}_{0}^{IPL}=Val_{+}(\mathcal{M}_{IPL})$;
-  2. ii)
-    $\mathcal{L}^{IPL}_{k+1}=\{v\in\mathcal{L}^{IPL}_{k}\mid\forall\alpha\in For(\Omega)_{k+1}$, if $v(\alpha)=\mathbf{U}$ then there exists $w\in\mathcal{L}^{IPL}_{k}$ such that $w(\alpha)=\mathbf{F}$ and: $v(\beta)=\mathbf{T}$ implies $w(\beta)=\mathbf{T}$ for every $\beta\in For(\Omega)\}$.
-Definition 4.16.
-The set of level valuations in $\mathcal{M}_{IPL}$ is defined as the intersection of every level valuation:
-$\mathcal{L}_{IPL}=\cap^{\infty}_{k\geq 0}\mathcal{L}^{IPL}_{k}$
-Definition 4.17 (RNmatrix for IPL).
-Let $\mathcal{M}_{IPL}$ be the Nmatrix defined in Definition 4.10. The RNmatrix for IPL is the pair $\mathcal{R}(\mathcal{M}_{IPL})=\langle\mathcal{M}_{IPL},\mathcal{L}_{IPL}\rangle$.
-From now on, $H_{IPL}$ will denote the standard Hilbert calculus for IPL over $\Omega$, given by the following axioms, and with Modus Ponens (MP) as the only inference rule:
-$\begin{array}[]{ll}\textsf{(Ax1)}&\alpha\to(\beta\to\alpha)\\[5.69054pt] \textsf{(Ax2)}&(\alpha\to(\beta\to\gamma))\to((\alpha\to\beta)\to(\alpha\to\gamma))\\[5.69054pt] \textsf{(Ax3)}&\alpha\to(\beta\to(\alpha\land\beta))\\[5.69054pt] \textsf{(Ax4)}&(\alpha\land\beta)\to\alpha\\[5.69054pt] \textsf{(Ax5)}&(\alpha\land\beta)\to\beta\\[5.69054pt] \textsf{(Ax6)}&\alpha\to(\alpha\lor\beta)\\[5.69054pt] \textsf{(Ax7)}&\beta\to(\alpha\lor\beta)\\[5.69054pt] \textsf{(Ax8)}&(\alpha\to\gamma)\to((\beta\to\gamma)\to((\alpha\lor\beta)\to\gamma))\\[5.69054pt] \textsf{(Ax9)}&(\beta\to\alpha)\to((\beta\to\neg\alpha)\to\neg\beta)\\[5.69054pt] \textsf{(Ax10)}&\alpha\to(\neg\alpha\to\beta)\\[5.69054pt] \end{array}$
-The consequence relation of $H_{IPL}$ will be denoted by $\vdash_{IPL}$.
-Observe that derivations in $H_{IPL}$ from premises are defined as usual in Hilbert calculi.
-That is, premises can be used in the derivations in the same way as axioms.
-The proof of soundness of $H_{IPL}$ w.r.t. the RNmatrix $\mathcal{R}(\mathcal{M}_{IPL})$ will be postponed to Subsection 4.2.2. We will concentrate now on the completeness proof.
+By induction on $k$.
+Base $k=0$: Suppose that $\Gamma\vdash_{IPL}^{0}\Delta$.
+Then $\Gamma\Rightarrow\Delta$ is an instance of the axiom ($Ax$), that is: $\Gamma\Rightarrow\Delta=\alpha,\Gamma^{\prime}\Rightarrow\alpha$ for some multiset $\Gamma^{\prime}$.
+Let $v\in\mathcal{L}_{0}^{IPL}$ such that $v\models\Gamma^{\prime}\cup\{\alpha\}$.
+In particular, $v\models\alpha$, hence $v\models\Gamma\Rightarrow\Delta$.
+From this, $\Gamma\models_{0}\Delta$.
+Inductive step: Suppose that, for any sequent $\Gamma\Rightarrow\Delta$, if $\Gamma\vdash_{IPL}^{j}\Delta$ then $\Gamma\models_{j}\Delta$, for every $j\leq k$, for a given $k\geq 0$ (Induction Hypothesis, IH).
+Let $\Gamma\Rightarrow\Delta$ be a sequent such that $\Gamma\vdash_{IPL}^{k+1}\Delta$.
+If $\Gamma\Rightarrow\Delta$ is an instance of the axiom, it can be proven as above that $\Gamma\models_{k+1}\Delta$.
+Suppose now that $\Gamma\Rightarrow\Delta$ was obtained from other sequents by applying a rule of LI in a derivation of depth $k+1$.
+We have the following cases:
+($contr\Rightarrow$): Suppose that $\Gamma\Rightarrow\Delta=\alpha,\Gamma^{\prime}\Rightarrow\Delta$ is obtained from $\alpha,\alpha,\Gamma^{\prime}\Rightarrow\Delta$ by applying rule ($contr\Rightarrow$).
+Then $\alpha,\alpha,\Gamma^{\prime}\vdash_{IPL}^{k}\Delta$ and so $\alpha,\alpha,\Gamma^{\prime}\models_{k}\Delta$, by (IH).
+But the latter clearly implies that $\alpha,\Gamma^{\prime}\models_{k}\Delta$ and so $\alpha,\Gamma^{\prime}\models_{k+1}\Delta$, given that $\mathcal{L}_{k+1}^{IPL}\subseteq\mathcal{L}_{k}^{IPL}$.
+($\Rightarrow w$): Suppose that $\Gamma\Rightarrow\Delta=\Gamma\Rightarrow\alpha$ is obtained from $\Gamma\Rightarrow$ by ($\Rightarrow w$).
+Then $\Gamma\vdash_{IPL}^{k}$ and so, by (IH), $\Gamma\models_{k}$.
+This means that, for every $v\in\mathcal{L}_{k}^{IPL}$, $v\not\models\Gamma$ and so $v\models\Gamma\Rightarrow\alpha$.
+From this it follows that $\Gamma\models_{k+1}\Delta$, given that $\mathcal{L}_{k+1}^{IPL}\subseteq\mathcal{L}_{k}^{IPL}$.
+(cutcut): Assume that $\Gamma\Rightarrow\Delta$ follows from $\Gamma\Rightarrow\alpha$ and $\alpha,\Gamma\Rightarrow\Delta$ by (cutcut) rule.
+Then $\Gamma\vdash_{IPL}^{k}\alpha$ and $\alpha,\Gamma\vdash_{IPL}^{k}\Delta$ and so, by (IH), $\Gamma\models_{k}\alpha$ and $\alpha,\Gamma\models_{k}\Delta$.
+Suppose first that $\Delta=\emptyset$.
+Let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma$; in particular, $v\in\mathcal{L}_{k}^{IPL}$ and so $v\models\Gamma\Rightarrow\alpha$.
+From this it follows that $v\models\alpha$.
+But this is impossible, given that $v\models\Gamma\cup\{\alpha\}\Rightarrow$, by (IH).
+From this, $v\not\models\Gamma$ for every $v\in\mathcal{L}_{k+1}^{IPL}$.
+That is, $v\models\Gamma\Rightarrow$, for every $v\in\mathcal{L}_{k+1}^{IPL}$.
+Now, suppose that $\Delta=\{\gamma\}$ for some formula $\gamma$, and let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma$.
+Once again, since $v\models\Gamma\Rightarrow\alpha$ we infer that $v\models\alpha$.
+But then $v\models\Gamma\cup\{\alpha\}$ and so $v\models\gamma$, given that $v\models\Gamma\cup\{\alpha\}\Rightarrow\gamma$.
+This means that $v\models\Gamma\Rightarrow\Delta$, for every $v\in\mathcal{L}_{k+1}^{IPL}$.
+That is, $\Gamma\models_{k+1}\Delta$.
+($\vee\Rightarrow$): Assume that $\Gamma\Rightarrow\Delta=\alpha\vee\beta,\Gamma^{\prime}\Rightarrow\Delta$ follows from $\alpha,\Gamma^{\prime}\Rightarrow\Delta$ and $\beta,\Gamma^{\prime}\Rightarrow\Delta$ by ($\vee\Rightarrow$).
+By (IH), $\alpha,\Gamma^{\prime}\models_{k}\Delta$ and $\beta,\Gamma^{\prime}\models_{k}\Delta$.
+Suppose first that $\Delta=\emptyset$.
+By hypothesis, if $v\in\mathcal{L}_{k+1}^{IPL}$ is such that $v\models\Gamma^{\prime}$, then $v\not\models\alpha$ and $v\not\models\beta$, and so $v\not\models\alpha\vee\beta$.
+This means that $v\models\Gamma^{\prime}\cup\{\alpha\vee\beta\}\Rightarrow$, for every $v\in\mathcal{L}_{k+1}^{IPL}$.
+Now, if $\Delta=\{\gamma\}$ for some formula $\gamma$, let $v\in\mathcal{L}_{k+1}^{IPL}$.
+By hypothesis, if $v\models\Gamma^{\prime}$ and $v\models\alpha$, then $v\models\gamma$; and if $v\models\Gamma^{\prime}$ and $v\models\beta$, then $v\models\gamma$.
+From this, if $v\models\Gamma^{\prime}$ and $v\models\alpha\vee\beta$ then, by definition of $\mathcal{M}_{IPL}$, either $v\models\alpha$ or $v\models\beta$ and so, by hypothesis, $v\models\gamma$.
+That is, $v\models\Gamma\Rightarrow\Delta$ for every $v\in\mathcal{L}_{k+1}^{IPL}$ and so $\Gamma\models_{k+1}\Delta$.
+($\Rightarrow\vee_{i}$): It is enough to observe that, if either $v\models\alpha$ or $v\models\beta$ then $v\models\alpha\vee\beta$, by definition of $\mathcal{M}_{IPL}$.
+($\wedge\Rightarrow$): Suppose that $\Gamma\Rightarrow\Delta=\alpha\land\beta,\Gamma^{\prime}\Rightarrow\Delta$ follows from $\alpha,\beta,\Gamma^{\prime}\Rightarrow\Delta$ by ($\wedge\Rightarrow$).
+Assume first that $\Delta=\emptyset$.
+Then, if $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma^{\prime}$ it follows that $v\not\models\{\alpha,\beta\}$, since $\mathcal{L}_{k+1}^{IPL}\subseteq\mathcal{L}_{k}^{IPL}$ and $\alpha,\beta,\Gamma^{\prime}\models_{k}\Delta$ by (IH).
+From this, either $v\not\models\alpha$ or $v\not\models\beta$.
+By definition of $\mathcal{M}_{IPL}$, $v\not\models\alpha\land\beta$.
+That is, $v\not\models\Gamma^{\prime}\cup\{\alpha\land\beta\}$ for every $v\in\mathcal{L}_{k+1}^{IPL}$, i.e., $\alpha\wedge\beta,\Gamma^{\prime}\models_{k+1}$.
+In turn, if $\Delta=\{\gamma\}$ for some formula $\gamma$, let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma^{\prime}\cup\{\alpha\land\beta\}$.
+By definition of $\mathcal{M}_{IPL}$, $v\models\Gamma^{\prime}\cup\{\alpha,\beta\}$ and so $v\models\gamma$, by (IH) and the fact that $v\in\mathcal{L}_{k}^{IPL}$.
+That is, $v\models\Gamma^{\prime}\cup\{\alpha\land\beta\}\Rightarrow\Delta$ for every $v\in\mathcal{L}_{k+1}^{IPL}$.
+($\Rightarrow\wedge$): It is enough to observe that, if both $v\models\alpha$ and $v\models\beta$ then $v\models\alpha\wedge\beta$, by definition of $\mathcal{M}_{IPL}$.
+($\to\Rightarrow$) Suppose that $\Gamma\Rightarrow\Delta=\alpha\to\beta,\Gamma^{\prime}\Rightarrow\Delta$ is obtained from $\Gamma^{\prime}\Rightarrow\alpha$ and $\beta,\Gamma^{\prime}\Rightarrow\Delta$ by ($\to\Rightarrow$).
+By (IH), $\Gamma^{\prime}\models_{k}\alpha$ and $\beta,\Gamma^{\prime}\models_{k}\Delta$.
+Assume first that $\Delta=\emptyset$.
+Let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma^{\prime}\cup\{\alpha\to\beta\}$.
+Then, $v\in\mathcal{L}_{k}^{IPL}$ such that $v\models\Gamma^{\prime}$ and so $v\models\alpha$, by hypothesis.
+Since $v\models\alpha\to\beta$ then $v\models\beta$, by definition of $\mathcal{M}_{IPL}$.
+That is, $v\in\mathcal{L}_{k}^{IPL}$ is such that $v\models\Gamma^{\prime}\cup\{\beta\}$, which contradicts the fact that $\beta,\Gamma^{\prime}\models_{k}$.
+From this, $v\not\models\Gamma^{\prime}\cup\{\alpha\to\beta\}$ for every $v\in\mathcal{L}_{k+1}^{IPL}$, which means that $\alpha\to\beta,\Gamma^{\prime}\models_{k+1}$.
+Now, suppose that $\Delta=\{\gamma\}$ for some formula $\gamma$.
+Let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma^{\prime}\cup\{\alpha\to\beta\}$.
+By reasoning as above, $v\in\mathcal{L}_{k}^{IPL}$ is such that $v\models\Gamma^{\prime}\cup\{\beta\}$.
+But then $v\models\gamma$, by hypothesis.
+This implies that $v\models\Gamma^{\prime}\cup\{\alpha\to\beta\}\Rightarrow\gamma$ for every $v\in\mathcal{L}_{k+1}^{IPL}$.
+($\Rightarrow\to$): Assume that $\Gamma\Rightarrow\Delta=\Gamma\Rightarrow\alpha\to\beta$ is obtained from $\alpha,\Gamma\Rightarrow\beta$ by ($\Rightarrow\to$).
+By (IH), $\alpha,\Gamma\models_{k}\beta$.
+Let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma$.
+Suppose that $v(\alpha\to\beta)=\mathbf{F}$.
+By definition of $\mathcal{M}_{IPL}$, $v(\alpha)=\mathbf{T}$ and $v(\beta)\in\{\mathbf{U},\mathbf{F}\}$.
+That is, $v\models\Gamma\cup\{\alpha\}$ but $v\not\models\beta$, which contradicts the fact that $\alpha,\Gamma\models_{k}\beta$, given that $v\in\mathcal{L}_{k}^{IPL}$.
+Hence, $v(\alpha\to\beta)\neq\mathbf{F}$.
+Suppose now that $v(\alpha\to\beta)=\mathbf{U}$.
+By Definition 8.6 of level valuations, there exists $w\in\mathcal{L}_{k}^{IPL}$ such that $w(\alpha\to\beta)=\mathbf{F}$ and $w\models\Gamma$.
+By reasoning as above, $w(\alpha)=\mathbf{T}$ and $w(\beta)\in\{\mathbf{U},\mathbf{F}\}$.
+But then $w\models\Gamma\cup\{\alpha\}$ and $w\not\models\beta$, which contradicts the fact that $\alpha,\Gamma\models_{k}\beta$.
+From this, it follows that $v(\alpha\to\beta)=\mathbf{T}$.
+That is, $v\models\Gamma\Rightarrow\alpha\to\beta$ for every $v\in\mathcal{L}_{k+1}^{IPL}$.
+($\neg\Rightarrow$): Assume that $\Gamma\Rightarrow\Delta=\neg\alpha,\Gamma^{\prime}\Rightarrow$ is obtained from $\Gamma^{\prime}\Rightarrow\alpha$ by ($\neg\Rightarrow$).
+By (IH), $\Gamma^{\prime}\models_{k}\alpha$.
+Let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma^{\prime}\cup\{\neg\alpha\}$.
+Then, $v\in\mathcal{L}_{k}^{IPL}$ such that $v\models\Gamma^{\prime}$ and so $v\models\alpha$, by hypothesis, which contradicts the fact that $v\models\neg\alpha$.
+From this, $v\not\models\Gamma^{\prime}\cup\{\neg\alpha\}$ for every $v\in\mathcal{L}_{k+1}^{IPL}$, which means that $\neg\alpha,\Gamma^{\prime}\models_{k+1}$.
+($\Rightarrow\neg$): Assume that $\Gamma\Rightarrow\Delta=\Gamma\Rightarrow\neg\alpha$ is obtained from $\alpha,\Gamma\Rightarrow$ by ($\Rightarrow\neg$).
+By (IH), $\alpha,\Gamma\models_{k}$.
+Let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma$.
+Suppose that $v(\neg\alpha)=\mathbf{F}$.
+By definition of $\mathcal{M}_{IPL}$, $v(\alpha)=\mathbf{T}$, hence $v\models\Gamma\cup\{\alpha\}$, which contradicts the fact that $\alpha,\Gamma\models_{k}$, since $v\in\mathcal{L}_{k}$.
+Now, suppose that $v(\neg\alpha)=\mathbf{U}$.
+By Definition 8.6 of level valuations, there exists $w\in\mathcal{L}_{k}^{IPL}$ such that $w(\neg\alpha)=\mathbf{F}$ and $w\models\Gamma$.
+As proven above, $w(\alpha)=\mathbf{T}$ and so $w\models\Gamma\cup\{\alpha\}$, contradicting the fact that $\alpha,\Gamma\models_{k}$.
+From this, it follows that $v(\neg\alpha)=\mathbf{T}$.
+That is, $v\models\Gamma\Rightarrow\neg\alpha$ for every $v\in\mathcal{L}_{k+1}^{IPL}$.
+This concludes the proof. ∎
+Theorem 8.12 (Soundness of LI w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$).
+For every set $\Gamma\cup\{\varphi\}$ of formulas over $\Omega$: $\Gamma\vdash_{IPL}\varphi$ implies that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$.
+Proof.
+Suppose that $\Gamma\vdash_{IPL}\varphi$.
+By definition, there exists a finite set $\Gamma_{0}\subseteq\Gamma$ such that the sequent $\Gamma_{0}\Rightarrow\varphi$ is derivable in LI.
+But then, there exists $k\geq 0$ such that $\Gamma_{0}\vdash_{IPL}^{k}\varphi$.
+By Proposition 8.11, $\Gamma_{0}\models_{k}\varphi$.
+This implies that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$, given that $\mathcal{L}_{IPL}\subseteq\mathcal{L}_{k}^{IPL}$.
+∎
+We will concentrate now on the completeness proof of LI w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$.
 Remark 8.13.
-Recall from Subsection 3.2 the notion of $\varphi$-saturated set in L.
-If $\Delta$ is a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$ (that is, IPL) then: $\alpha\in\Delta$ implies that $\neg\alpha\notin\Delta$; $\alpha\vee\beta\in\Delta$ iff $\alpha\in\Delta$ or $\beta\in\Delta$; $\alpha\land\beta\in\Delta$ iff $\alpha\in\Delta$ and $\beta\in\Delta$; $\alpha\to\beta\in\Delta$ and $\alpha\in\Delta$ implies that $\beta\in\Delta$; and $\beta\in\Delta$ implies that $\alpha\to\beta\in\Delta$.
+Recall from Remark 6.3 that, in any Tarskian and finitary logic L, if $\Gamma\nvdash_{\bf L}\varphi$ then there exists a $\varphi$-saturated set $\Delta$ in L such that $\Gamma\subseteq\Delta$.
+This property holds, in particular, in the logic generated by LI (that is, IPL).
+Moreover, it is easy to prove that, if $\Delta$ is a $\varphi$-saturated set in IPL then: $\alpha\in\Delta$ implies that $\neg\alpha\notin\Delta$; $\alpha\vee\beta\in\Delta$ iff $\alpha\in\Delta$ or $\beta\in\Delta$; $\alpha\land\beta\in\Delta$ iff $\alpha\in\Delta$ and $\beta\in\Delta$; $\alpha\to\beta\in\Delta$ and $\alpha\in\Delta$ implies that $\beta\in\Delta$; and $\beta\in\Delta$ implies that $\alpha\to\beta\in\Delta$.
 Definition 8.14 (Valuation associated to a $\varphi$-saturated set in IPL).
-Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$ (that is, IPL).
-The valuation associated to $\Delta$ in IPL is the function $v_{\Delta}:For(\Omega)\to V_{IPL}$ defined inductively as follows (here, $p$ is a propositional variable):
+Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by LI (that is, IPL).
+The valuation associated to $\Delta$ in IPL is the function $v_{\Delta}:For(\Omega)\to V_{IPL}$ defined inductively as follows (here, $p$ is a propositional variable and $\alpha,\beta\in For(\Omega)$):
 […]
@@
 Remark 8.15.
-Observe that $v_{\Delta}(\neg\alpha)=\mathbf{U}$ whenever $\alpha$ is undetermined in $\Delta$, that is: $\neg\alpha\notin\Delta$ and $\alpha\notin\Delta$.
-Hence, the failure of excluded middle (which is a principle valid in classical logic but not in IPL) occurs only for formulas $\alpha$ with such undetermined status in $\Delta$.
-Analogously, $v_{\Delta}(\alpha\to\beta)=\mathbf{U}$ whenever $\alpha$ and $\alpha\to\beta$ are jointly undetermined in $\Delta$, that is: neither $\alpha$ nor $\alpha\to\beta$ belong to $\Delta$.
-Hence, the failure of $\alpha\vee(\alpha\to\beta)$ (which is a principle valid in classical logic but not in IPL) occurs only for formulas $\alpha$ and $\alpha\to\beta$ with such joint undetermined status in $\Delta$.
+The observations about the meaning of $\mathbf{U}$ made after Definition 8.4 can be recast now.
+Observe that $v_{\Delta}(\neg\alpha)=\mathbf{U}$ whenever $\alpha$ is unknown in $\Delta$, that is: $\neg\alpha\notin\Delta$ and $\alpha\notin\Delta$.
+Hence, the failure of excluded middle $\alpha\vee\neg\alpha$ (a principle valid in classical logic but not in IPL) occurs only for formulas $\alpha$ with such undetermined status in $\Delta$.
+Analogously, $v_{\Delta}(\alpha\to\beta)=\mathbf{U}$ whenever $\alpha$ and $\alpha\to\beta$ are jointly unknown in $\Delta$, that is: neither $\alpha$ nor $\alpha\to\beta$ belong to $\Delta$.
+Hence, the failure of Peirce’s law $\alpha\vee(\alpha\to\beta)$ (a principle valid in classical logic but not in IPL) occurs only for formulas $\alpha$ and $\alpha\to\beta$ with such joint undetermined status in $\Delta$.
+The function $v_{\Delta}$ can be seen as a 3-valued characteristic map associated to $\Delta$, in which the third value $\mathbf{U}$ reflects the undetermined status of a sentence that does not belong to $\Delta$.
+Only sentences of the form $\neg\alpha$ or $\alpha\to\beta$ can be unknown or undetermined (in this sense) w.r.t.
+a $\varphi$-saturated set in $\mathbf{IPL}$.
 Proposition 8.16.
-Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$.
-Then, the valuation $v_{\Delta}$ associated to $\Delta$ in IPL is a valuation in the Nmatrix $\mathcal{M}_{IPL}$ which belongs to $Val_{+}(\mathcal{M}_{IPL})$ (recall Definition 4.14) such that, for every $\alpha$: $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$.
+Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by LI.
+Then, the valuation $v_{\Delta}$ associated to $\Delta$ in IPL is a valuation in the Nmatrix $\mathcal{M}_{IPL}$ which belongs to $Val^{nd}_{+}(\mathcal{M}_{IPL})$ (recall Definition 8.5) such that, for every $\alpha$: $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$.
 Proof.
@@
 Moreover, $v_{\Delta}(p)\in\{\mathbf{T},\mathbf{F}\}$ for every propositional variable $p$.
-Negation: Let us prove first that $v_{\Delta}(\neg\alpha)\in\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$ for every $\alpha$.
+Negation: Let us prove first that $v_{\Delta}(\neg\alpha)\in\neg^{IPL}(v_{\Delta}(\alpha))$ for every $\alpha$.
 If $v_{\Delta}(\neg\alpha)=\mathbf{T}$ then $\neg\alpha\in\Delta$, by definition.
 Hence, $\alpha\notin\Delta$ and so $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$.
-From this, $v_{\Delta}(\neg\alpha)=\mathbf{T}\in\{\mathbf{U},\mathbf{T}\}=\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$.
+From this, $v_{\Delta}(\neg\alpha)=\mathbf{T}\in\{\mathbf{U},\mathbf{T}\}=\neg^{IPL}(v_{\Delta}(\alpha))$.
 Now, if $v_{\Delta}(\neg\alpha)=\mathbf{U}$ then $\alpha\notin\Delta$, by definition.
-Hence $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and so $v_{\Delta}(\neg\alpha)=\mathbf{U}\in\{\mathbf{U},\mathbf{T}\}=\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$.
+Hence $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and so $v_{\Delta}(\neg\alpha)=\mathbf{U}\in\{\mathbf{U},\mathbf{T}\}=\neg^{IPL}(v_{\Delta}(\alpha))$.
 Finally, if $v_{\Delta}(\neg\alpha)=\mathbf{F}$ then $\alpha\in\Delta$, by definition.
-Hence $v_{\Delta}(\alpha)=T$ and so $v_{\Delta}(\neg\alpha)=\mathbf{F}\in\{\mathbf{F}\}=\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$.
+Hence $v_{\Delta}(\alpha)=\mathbf{T}$ and so $v_{\Delta}(\neg\alpha)=\mathbf{F}\in\{\mathbf{F}\}=\neg^{IPL}(v_{\Delta}(\alpha))$.
 Implication: There are 3 cases:
 (I1) Suppose that $v_{\Delta}(\alpha\to\beta)=\mathbf{T}$. Then $\alpha\to\beta\in\Delta$, by definition.
-– If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{T}\in{\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
-– If $\alpha\in\Delta$ then $\beta\in\Delta$, by MP.
-Hence $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=\mathbf{T}$ and so $v_{\Delta}(\alpha\to\beta)=\mathbf{T}\in\{\mathbf{T}\}={\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+– If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{T}\in{\to}^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+– If $\alpha\in\Delta$ then $\beta\in\Delta$, by (DM).
+Hence $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=\mathbf{T}$ and so $v_{\Delta}(\alpha\to\beta)=\mathbf{T}\in\{\mathbf{T}\}={\to}^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 (I2) Suppose that $v_{\Delta}(\alpha\to\beta)=\mathbf{U}$.
@@
 Then $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$.
-Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{U}\in\{\mathbf{U},\mathbf{T}\}={\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{U}\in\{\mathbf{U},\mathbf{T}\}={\to}^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 (I3) Suppose that $v_{\Delta}(\alpha\to\beta)=\mathbf{F}$.
@@
 Then $v_{\Delta}(\alpha)=\mathbf{T}$ and $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$.
-Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{F}\in\{\mathbf{F}\}={\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{F}\in\{\mathbf{F}\}={\to}^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 Conjunction: There are 2 cases:
@@
 From this, $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=\mathbf{T}$.
-Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{T}\in\{\mathbf{T}\}=\land^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{T}\in\{\mathbf{T}\}=\land^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 (C2) Suppose that $v_{\Delta}(\alpha\land\beta)=\mathbf{F}$.
@@
 From this, $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ or $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$.
-Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{F}\in\{\mathbf{F}\}=\land^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{F}\in\{\mathbf{F}\}=\land^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 Disjunction: There are 2 cases:
@@
 From this, $v_{\Delta}(\alpha)=\mathbf{T}$ or $v_{\Delta}(\beta)=\mathbf{T}$.
-Hence, $v_{\Delta}(\alpha\vee\beta)=\mathbf{T}\in\{\mathbf{T}\}=\vee^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+Hence, $v_{\Delta}(\alpha\vee\beta)=\mathbf{T}\in\{\mathbf{T}\}=\vee^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 (D2) Suppose that $v_{\Delta}(\alpha\vee\beta)=\mathbf{F}$.
@@
 From this, $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$.
-Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{F}\in\{\mathbf{F}\}=\vee^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
+Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{F}\in\{\mathbf{F}\}=\vee^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
 This concludes the proof. ∎
 Proposition 8.17.
-Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$.
+Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by LI.
 Then, the valuation $v_{\Delta}$ associated to $\Delta$ in IPL is a level valuation in $\mathcal{M}_{IPL}$ such that, for every $\alpha$: $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$.
@@
 By Proposition 8.16, $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$ for every $\alpha$. It will be shown, by induction on $n$, that $v_{\Delta}\in\mathcal{L}^{IPL}_{n}$ for every $n\geq 0$.
-Base $n=0$: $v_{\Delta}\in\mathcal{L}^{IPL}_{0}=Val_{+}(\mathcal{M}_{IPL})$, by Proposition 4.21.
-[…]
+Base $n=0$: $v_{\Delta}\in\mathcal{L}^{IPL}_{0}=Val^{nd}_{+}(\mathcal{M}_{IPL})$, by Proposition 8.16.
+Inductive step: Assume that, for every $\varphi^{\prime}$ and for every $\varphi^{\prime}$-saturated set $\Delta^{\prime}$, $v_{\Delta^{\prime}}\in\mathcal{L}^{IPL}_{k}$ for every $k\leq n$, for a given $n\geq 0$ (Induction Hypothesis, IH).
+Let us prove that $v_{\Delta}\in\mathcal{L}^{IPL}_{n+1}$.
+By IH, $v_{\Delta}\in\mathcal{L}^{IPL}_{n}$.
+Let $\alpha\in For(\Omega)$ such that $v_{\Delta}(\alpha)=\mathbf{U}$.
+By Definition 8.14, $\alpha\neq p$ for every propositional variable $p$; and $\alpha\neq\beta\#\gamma$ for $\in\{\land$$\vee\}$ and every formulas $\beta,\gamma$.
+Then, there are two cases to analyze:
 (1) $\alpha=\neg\beta$ for some $\beta$.
@@
 Suppose that $\Delta,\beta\vdash_{IPL}\neg\beta$.
-By the deduction metatheorem, $\Delta\vdash_{IPL}\beta\to\neg\beta$.
-By (Ax9) and by MP, $\Delta,\beta\to\beta,\beta\to\neg\beta\vdash_{IPL}\neg\beta$.
-But $\beta\to\beta\in\Delta$ and $\beta\to\neg\beta\in\Delta$, hence $\Delta\vdash_{IPL}\neg\beta$, that is, $\neg\beta\in\Delta$, a contradiction.
+By the deduction metatheorem (DM), $\Delta\vdash_{IPL}\beta\to\neg\beta$, and so the sequent $\Delta_{0}\Rightarrow\beta\to\neg\beta$ is provable in LI, for some finite set $\Delta_{0}\subseteq\Delta$.
+But, for every finite multiset $\Gamma$, the sequent $\Gamma,\beta\to\neg\beta\Rightarrow\neg\beta$ is provable in LI.
+Indeed, from axiom $\Gamma,\beta\Rightarrow\beta$ it follows $\Gamma,\beta,\neg\beta\Rightarrow$, by ($\neg\Rightarrow$).
+From this, and by using axiom $\Gamma,\beta\Rightarrow\beta$ once again, by ($\to\Rightarrow$) we infer $\Gamma,\beta,\beta\to\neg\beta\Rightarrow$.
+Finally, $\Gamma,\beta\to\neg\beta\Rightarrow\neg\beta$ follows by ($\Rightarrow\neg$).
+In particular, $\Delta_{0},\beta\to\neg\beta\Rightarrow\neg\beta$ is provable in LI and so, by (cutcut) with $\Delta_{0}\Rightarrow\beta\to\neg\beta$, it follows that $\Delta_{0}\Rightarrow\neg\beta$ is provable in LI.
+That is, $\Delta\vdash_{IPL}\neg\beta$, hence $\neg\beta\in\Delta$, a contradiction.
 From this, $\Delta,\beta\nvdash_{IPL}\neg\beta$.
@@
 Suppose that $\Delta,\beta\vdash_{IPL}\beta\to\gamma$.
-By MP, $\Delta,\beta\vdash_{IPL}\gamma$ and so, by the deduction metatheorem, $\Delta\vdash_{IPL}\beta\to\gamma$, a contradiction.
+By (DM), $\Delta,\beta\vdash_{IPL}\gamma$ and so, by (DM) again, $\Delta\vdash_{IPL}\beta\to\gamma$, a contradiction.
 From this, $\Delta,\beta\nvdash_{IPL}\beta\to\gamma$.
@@
 From the previous analysis we conclude that $v_{\Delta}\in\mathcal{L}^{IPL}_{n+1}$. This shows that $v_{\Delta}\in\mathcal{L}^{IPL}_{n}$ for every $n\geq 0$ and so $v_{\Delta}\in\mathcal{L}_{IPL}$. ∎
-Theorem 4.23 (Completeness of $\mathcal{H}_{IPL}$ w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$).
+Theorem 8.18 (Completeness of LI w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$).
 For every $\Gamma$ and $\varphi$: $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$ implies that $\Gamma\vdash_{IPL}\varphi$.
@@
 Suppose that $\Gamma\nvdash_{IPL}\varphi$.
-Then, there exists a $\varphi$-saturated set $\Delta$ in IPL such that $\Gamma\subseteq\Delta$.
+Then, there exists a $\varphi$-saturated set $\Delta$ in LI such that $\Gamma\subseteq\Delta$.
 By Proposition 8.17, $v_{\Delta}$ is a level valuation in $\mathcal{M}_{IPL}$ such that $v_{\Delta}(\alpha)=\mathbf{T}$ for every $\alpha\in\Gamma$, but $v_{\Delta}(\varphi)\neq\mathbf{T}$.
@@
 ∎
-4.2.2 Intuitionistic partial (level) valuations and intuitionistic truth tables
+Remark 8.19.
+Observe that the semantical characterization of $\mathbf{IPL}$ w.r.t.
+level valuation semantics is quite interesting from a theoretical point of view.
+However, it has little interest from a practical point of view, given that level valuations are infinitary objects, and it is far from obvious how to check validity of a formula in $\mathbf{IPL}$ by using $\mathcal{R}(\mathcal{M}_{IPL})$.
+The next step is the definition of an (algorithmic) way to construct intuitionistic truth-tables, based precisely on level valuations.
+From soundness and completeness of $\mathbf{IPL}$ w.r.t.
+$\mathcal{R}(\mathcal{M}_{IPL})$, several important consequences can be obtained.
+Some of these consequence will be extremely useful to prove the soundness and completeness of the (non-deterministic) truth-tables decision method for $\mathbf{IPL}$ to be defined in the next section.
+Corollary 8.19.1 (Finitariness for $\mathcal{R}(\mathcal{M}_{IPL})$).
+Let $\Gamma\cup\{\varphi\}\subseteq For(\Omega)$.
+If $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$ then $\Gamma_{0}\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$ for some finite subset $\Gamma_{0}$ of $\Gamma$.
+Proof.
+It follows from Theorems 8.18, 8.12 and the fact that LI is finitary. ∎
+Corollary 8.19.2 (Deduction metatheorem for $\mathcal{R}(\mathcal{M}_{IPL})$).
+Let $\Gamma\cup\{\alpha,\beta\}\subseteq For(\Omega)$.
+Then, $\Gamma,\alpha\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$ if, and only if, $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$.
+Proof.
+It follows from Theorems 8.18, 8.12 and the fact that $\bf LI$} satisfies the deduction metatheorem (recall Remark 8.8). ∎
+Proposition 8.20.
+Let $v\in\mathcal{L}_{IPL}$ and let $\Gamma=\{\alpha\in For(\Omega)\mid v(\alpha)=\mathbf{T}\}$.
+Then, $\Gamma$ is a closed theory in $\mathbf{IPL}$, i.e.: if $\Gamma\vdash_{IPL}\alpha$ then $\alpha\in\Gamma$.
+Proof.
+Let $\Gamma=\{\alpha\in For(\Omega)\mid v(\alpha)=\mathbf{T}\}$ and suppose that $\Gamma\vdash_{IPL}\alpha$.
+By soundness of $\mathbf{IPL}$ w.r.t.
+level valuation semantics (Theorem 8.12), $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha$.
+Given that $v\in\mathcal{L}_{IPL}$ such that $v\models\Gamma$, it follows that $v\models\alpha$.
+That is, $v(\alpha)=\mathbf{T}$ and so $\alpha\in\Gamma$.
+∎
+The following result will be crucial for proving soundness and completeness of intuitionistic truth-tables:
+Proposition 8.21.
+Let $v\in\mathcal{L}_{IPL}$, and $\alpha\in For(\Omega)$ such that $v(\alpha)=\mathbf{U}$.
+Then, there exists $w\in\mathcal{L}_{IPL}$ such that $w(\alpha)=\mathbf{F}$ and $w(\gamma)=\mathbf{T}$ for every formula $\gamma$ such that $v(\gamma)=\mathbf{T}$.
+Proof.
+$v\in\mathcal{L}_{IPL}$, and $\alpha\in For(\Omega)$ such that $v(\alpha)=\mathbf{U}$.
+Since, in particular, $v\in Val^{nd}_{+}(\mathcal{M}_{IPL})$ (recall Definitions 8.5 and 8.6), $v(\alpha)=\mathbf{U}$ implies that $\alpha$ cannot be a propositional variable.
+By definition of the conjunction and disjunction tables in $\mathcal{M}_{IPL}$ (Figure 3), and since $v(\alpha)=\mathbf{U}$, it follows that $\alpha$ cannot be either a conjunction or a disjunction.
+Let $\Gamma=\{\gamma\in For(\Omega)\mid v(\gamma)=\mathbf{T}\}$.
+Then, $\alpha\not\in\Gamma$.
+By the observations above, there are only two cases to consider:
+(1) $\alpha=\neg\beta$ for some $\beta$.
+As noted above, $\alpha\not\in\Gamma$, i.e., $\neg\beta\not\in\Gamma$.
+Hence $\Gamma\nvdash_{IPL}\neg\beta$, given that $\Gamma$ is closed by Proposition 8.20.
+If $\beta\in\Gamma$ then $v(\beta)=\mathbf{T}$ and so, by definition of $\mathcal{M}_{IPL}$, $v(\alpha)=v(\neg\beta)=\mathbf{F}$, a contradiction.
+Then, $\beta\not\in\Gamma$.
+Suppose that $\Gamma,\beta\vdash_{IPL}\neg\beta$.
+As in the proof of Proposition 8.17 we infer that $\Gamma\vdash_{IPL}\neg\beta$.
+But this lead us to a contradiction since, as pointed out before, $\Gamma\nvdash_{IPL}\neg\beta$.
+From this, $\Gamma,\beta\nvdash_{IPL}\neg\beta$.
+Thus, there exists a $\neg\beta$-saturated set $\Delta$ in $\mathbf{IPL}$ such that $\Gamma\cup\{\beta\}\subseteq\Delta$.
+By Proposition 8.17, the function $v_{\Delta}$ is a level valuation such that $v_{\Delta}(\alpha)=\mathbf{F}$ (since $v_{\Delta}(\beta)=\mathbf{T}$) and $v_{\Delta}(\gamma)=\mathbf{T}$ for every formula $\gamma$ such that $v(\gamma)=\mathbf{T}$, as required.
+(2) $\alpha=\beta\to\delta$ for some $\beta$ and $\delta$.
+Since $\Gamma,\delta\vdash_{IPL}\beta\to\delta$ and $\Gamma\nvdash_{IPL}\beta\to\delta$ then $\Gamma\nvdash_{IPL}\delta$.
+Suppose that $\Gamma,\beta\vdash_{IPL}\beta\to\delta$.
+By (DM), $\Gamma,\beta\vdash_{IPL}\delta$ and so, by (DM) again, $\Gamma\vdash_{IPL}\beta\to\delta$, a contradiction.
+From this, $\Gamma,\beta\nvdash_{IPL}\beta\to\delta$.
+Thus, there exists a $(\beta\to\delta)$-saturated set $\Delta$ in $\mathbf{IPL}$ such that $\Gamma\cup\{\beta\}\subseteq\Delta$.
+By reasoning as above, we conclude that $\delta\not\in\Delta$.
+By Proposition 8.17, $v_{\Delta}\in\mathcal{L}_{IPL}$ is such that $v_{\Delta}(\alpha)=\mathbf{F}$ (since $v_{\Delta}(\beta)=\mathbf{T}$ and $v(\delta)\neq\mathbf{T}$) and $v_{\Delta}(\gamma)=\mathbf{T}$ for every formula $\gamma$ such that $v(\gamma)=\mathbf{T}$, as required.
+9 A truth-tables procedure for $\mathbf{IPL}$
+After obtaining a suitable RNmatrix for $\mathbf{IPL}$, the next step is the definition of a decision procedure from this.
+As in the case of S4, this procedure is given by means of branching truth-tables with an algorithm for deleting inadequate rows.
+9.1 Intuitionistic partial (level) valuations and intuitionistic truth-tables
 Definition 9.1.
 Consider the sets
-[…]
+$CS(\Omega)=\{\Lambda\subseteq For(\Omega)\mid\mbox{ $\Lambda$ is non-empty and closed under subformulas}\},$
 […]
@@
   • if $\alpha\in\mathcal{P}\cap\Lambda$ then $\tilde{v}_{p}(\alpha)\in\{\mathbf{T},\mathbf{F}\}$;
-  • if $\neg\alpha\in\Lambda$ then $\tilde{v}_{p}(\neg\alpha)\in\neg^{\mathcal{A}_{IPL}}(\tilde{v}_{p}(\alpha))$;
-[…]
+  • if $\neg\alpha\in\Lambda$ then $\tilde{v}_{p}(\neg\alpha)\in\neg^{IPL}(\tilde{v}_{p}(\alpha))$;
+  • if $#\in\{\to,\vee,\land\}$ and $\alpha\#\beta\in\Lambda$ then $\tilde{v}_{p}(\alpha\#\beta)\in\#^{IPL}(\tilde{v}_{p}(\alpha),\tilde{v}_{p}(\beta))$.
 Let $iPV(\Lambda)$ be the set of partial valuations in $\mathcal{M}_{IPL}$ with domain $\Lambda$.
 Definition 9.3 (Intuitionistic partial level valuation over $\Lambda$).
-Let $\Lambda\in FCS(\Omega)$. A partial valuation $\tilde{v}_{p}\in iPV(\Lambda)$ is an intuitionistic partial level valuation in $\mathcal{M}_{IPL}$ over $\Lambda$ iff
+Let $\Lambda\in FCS(\Omega)$. An intuitionistic partial level valuation in $\mathcal{M}_{IPL}$ over $\Lambda$ is a partial valuation $\tilde{v}_{p}\in iPV(\Lambda)$ satisfying the following:
   1.
-$\forall\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$, there exists an intuitionistic partial level valuation $\tilde{w}_{p}$ in $\mathcal{M}_{IPL}$ over $\Lambda$ such that $\tilde{w}_{p}(\alpha)=\mathbf{F}$ and, $\forall\beta\in\Lambda$, $\tilde{w}_{p}(\beta)=\mathbf{T}$ whenever $\tilde{v}_{p}(\beta)=bT$.
+$\forall\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$, there exists an intuitionistic partial level valuation $\tilde{w}_{p}$ in $\mathcal{M}_{IPL}$ over $\Lambda$ such that $\tilde{w}_{p}(\alpha)=\mathbf{F}$ and, $\forall\beta\in\Lambda$, $\tilde{w}_{p}(\beta)=\mathbf{T}$ whenever $\tilde{v}_{p}(\beta)=\mathbf{T}$.
 The set of intuitionistic partial level valuations in $\mathcal{M}_{IPL}$ over $\Lambda$ will be denoted by $iPLV(\Lambda)$.
 Remark 9.4.
-Given $v,w$ and $\alpha$ let $\mathsf{P}_{\Lambda}^{i}(v,w,\alpha)$ iff $w(\alpha)=\mathbf{F}$ and, $\forall\beta\in\Lambda$, $w(\beta)=\mathbf{T}$ whenever $v(\beta)=\mathbf{T}$.
+Given $v,w$ and $\alpha$ let $\mathsf{P}_{\Lambda}(v,w,\alpha)$ iff $w(\alpha)=\mathbf{F}$ and, $\forall\beta\in\Lambda$, $w(\beta)=\mathbf{T}$ whenever $v(\beta)=\mathbf{T}$.
 If $\Lambda\in FCS(\Omega)$ then
-$\begin{array}[]{lll}iPLV(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in iPLV(\Lambda)$}\big{)}\}.\end{array}$
-As observed in Remark 3.7 for $\mathcal{M}_{S4}$, since $\Lambda$ is finite then $iPV(\Lambda)$ is finite, hence the definition of $iPLV(\Lambda)$ is not cyclic.
+$\begin{array}[]{lll}iPLV(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big(\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in iPLV(\Lambda)$}\big)\}.\end{array}$
+Remark 9.5.
+As in the case of S4 discussed in Remark 5.6, the impredicative definition of $iPLV(\Lambda)$ does not lead to circularity.
+Indeed, by using an algorithm (see Algorithm 1 below) similar to the Algorithm 3.4 given in [12, p.
+15] for $\mathcal{M}_{S4}$, Definition 9.3 can be transformed into a predicative one (see Definition 9.11 and Proposition 9.19).
+This is the consequence of the fact that $iPLV(\Lambda)$ is obtained as the output of Algorithm 1, being the unique subset of $iPV(\Lambda)$ satisfying the condition of Remark 9.4, as it will be shown in Proposition 9.7.
+The set $iPLV(\Lambda)$ is precisely a (intuitionistic) truth-table, in the wider sense proposed here, based on $\mathcal{R}(\mathcal{M}_{IPL})$.
+It is obtained by means of a mechanical procedure, where its elements correspond to the rows of the table, and its columns correspond to the values assigned in each row to the elements of $\Lambda$.
+Notation 9.6.
+Given $\Lambda\in FCS(\Omega)$ with $n\geq 1$ elements, consider an arrangement $\alpha_{1},\ldots,\alpha_{n}$ of this set such that the complexity of $\alpha_{i}$ is less than or equal to the complexity of $\alpha_{j}$ if $i<j$ (so the atomic formulas occurring in $\Lambda$ appear as the first elements of the sequence, in arbitrary order).
+In this way, any subformula of $\alpha_{j}$ appears as $\alpha_{i}$ for some unique $i\leq j$.
+Suppose that the set $iPV(\Lambda)$ has $m\geq 1$ elements.
+Consider now an arrangement $v_{p}^{1},\ldots,v_{p}^{m}$ of its elements, ordered by lexicographical order, by stipulating that $\mathbf{F}<\mathbf{U}<\mathbf{T}$.
+In this way, $iPV(\Lambda)$ can be represented as a 2-dimensional array (i.e., an array of arrays, or a matrix) $V$ such that each row $V[i]$, for $1\leq i\leq m$, corresponds to the ithith element $v_{p}^{i}$ of $iPV(\Lambda)$.
+In turn, $V[i][j]$ is the truth-value assigned to the formula $\alpha_{j}\in\Lambda$ by the partial valuation $V[i]=v_{p}^{i}$.
+That is, $V[i][j]=v_{p}^{i}(\alpha_{j})$.
+This notation will be used in Algorithm 1, to be described in the sequel, also for subsets of $iPV(\Lambda)$.
+Algorithm 1 (in Figure 4) for obtaining $iPLV(\Lambda)$.
+Let $\Lambda\in FCS(\Omega)$, and consider the notation stipulated in Notation 9.6.
+Hence, $\Lambda=\{\alpha_{1},\ldots,\alpha_{n}\}$ and $iPV(\Lambda)=\{v_{p}^{1},\ldots,v_{p}^{m}\}$ (which can also be seen as sequences).
+Let $V:=iPLV(\Lambda)$.
+For each element $v_{p}^{i}$ of $V$, if $v_{p}^{i}(\alpha_{j})=\mathbf{U}$, determine if there is some $v_{p}^{i^{\prime}}$ such that $\mathsf{P}_{\Lambda}(v_{p}^{i},v_{p}^{i^{\prime}},\alpha_{j})$.
+This is represented by the function isSupported(V,i,j)(V,i,j) in Figure 4.
+If, for some $j$, isSupported(V,i,j)(V,i,j) is false (meaning no valuation in $V$ supports $v_{p}^{i}$ w.r.t.
+$\alpha_{j}$) then $v_{p}^{i}$ is removed from the candidate set for $iPLV(\Lambda)$ and will not be considered in subsequent algorithm iterations.
+After checking all the elements of $V$ using the elements of $V$ itself, let $IPV$ be the set of remaining valuations.
+If no valuation was removed, i.e., if $IPV=V$, the algorithm stops at this point and returns $IPV$ as being the set $iPLV(\Lambda)$ of partial level valuations.
+The set $IPV$ is the result of a refinement function applied to $V$, represented by refine(V)(V) in Figure 4.
+In case $IPV\neq V$, it means that $IPV$ is a proper subset of $V$ (i.e., some valuations were effectively removed in this round of the process).
+The set $IPV$ needs to be checked again with the function refine, to guarantee that every row is still supported.
+Then, the set $V$ is updated to be equal to $IPV$ (and so $V$ is, again, the universe of partial valuations to be checked), it is renumbered as a sequence $V[1],\ldots,V[m^{\prime}]$ (where $m^{\prime}<m$ is the cardinal or $IPV$) and then we repeat the process with the updated $V$ until no changes occur.
+The set $iPLV(\Lambda)$ is then defined to be the fixpoint $V=IPV$ (see Figure 4).
+Figure 4: Algorithm for obtaining the set $iPLV(\Lambda)$.
+Next result shows that, thanks to Algorithm 1, the set $iPLV(\Lambda)$ is well-defined, it is unique, and it is non-empty:
+Proposition 9.7 (Existence and uniqueness of intuitionistic partial level valuations).
+Let $\Lambda\in FCS(\Omega)$. Then, there exists a unique set $iPLV(\Lambda)\subseteq iPV(\Lambda)$ satisfying the condition of Remark 9.4:
+$\begin{array}[]{lll}iPLV(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big(\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in iPLV(\Lambda)$}\big)\}.\end{array}$
+Moreover, the set $iPLV(\Lambda)$ is non-empty, having at least $2^{k}$ elements, where $k$ is the number of different propositional variables occurring in $\Lambda$.
+Proof.
+Given $\Lambda\in FCS(\Omega)$, define the sequences $(\alpha_{1},\ldots,\alpha_{n})$ and $(v_{p}^{1},\ldots,v_{p}^{m})$ representing $\Lambda$ and $iPV(\Lambda)$ as in Notation 9.6, and then execute Algorithm 1 as described.
+Notice that $iPV(\Lambda)$ is a finite set, and the algorithm operates on a subset $V\subseteq iPV(\Lambda)$ in each round.
+If the algorithm does not stop in a given round, it deletes some elements from $V$ and restarts the process with a new subset that is strictly smaller than the previous one.
+This shows that the process will eventually terminate.
+There are two possible termination conditions: either every element of $V$ is eliminated, or a set $IPV=V$ is produced.
+In the second case, every element of $IPV$ is supported (as no removal is possible at this point); hence, $iPLV(\Lambda):=IPV$ satisfies the condition of Remark 9.4.
+In turn, the first possibility (arriving to a set $IPV=\emptyset$) is impossible by the following reason: every partial valuation $v:\Lambda\rightarrow V_{IPL}$ such that $v(\alpha)\in\{\mathbf{T},\mathbf{F}\}$ for every $\alpha\in\Lambda$ (i.e., every classical partial valuation) is never removed from $iPV(\Lambda)$ by means of Algorithm 1.
+Indeed, any classical valuation $v$ never satisfies that $v(\alpha_{j})=\mathbf{U}$ for some $j$, hence it will never be removed.
+Since there are exactly $2^{k}$ classical valuations in $iPV(\Lambda)$ (where $k$ is the number of different propositional variables occurring in $\Lambda$), the set $IPV$ produced by Algorithm 1 is non-empty, having at least $2^{k}$ elements.
+We have proven, therefore, the existence of a (non-empty) subset $IPV$ of $iPV(\Lambda)$ satisfying the condition of Remark 9.4.
+To prove the uniqueness of such a set of partial valuations, observe that the order of the valuations $V[i]$ in the array $V$ is irrelevant for Algorithm 1.
+The algorithm’s execution does not depend on this initial ordering.
+In each round, every valuation $V[i]$ is tested against every formula $\alpha_{j}$ in $\Lambda$.
+If $V[i][j]=\mathbf{U}$, the function isSupported(V,i,j)(V,i,j) checks all other valuations $V[i^{\prime}]$ in $V$ to determine whether $\mathsf{P}_{\Lambda}(V[i],V[i^{\prime}],\alpha_{j})$ holds.
+If this condition fails for every $i^{\prime}$, $V[i]$ is removed in the next step.
+This procedure is clearly independent of the lexicographical order used in the formal presentation of Algorithm 1: that order was chosen solely to provide a precise description of the algorithm’s steps.
+In other words, regardless of the initial arrangement of $iPV(\Lambda)$, the first round will eliminate the same valuations and produce a unique set $V_{1}$ for the second stage.
+The same argument applies: any arrangement of $V_{1}$ yields the same subset $V_{2}$ after the second round.
+By induction, the algorithm’s output — the fixpoint $V_{s}=IPV$ reached in the final step $s$ — must be unique.
+This completes the proof. ∎
+Remark 9.8.
+Observe that it is possible to have a set $iPLV(\Lambda)$ depending on $k$ different propositional variables $p_{1},\ldots,p_{k}$ having exactly $2^{k}$ elements.
+It is enough considering a formula $\alpha(p_{1},\ldots,p_{k})$ constructed by using exclusively conjunction and disjunction.
+Let $\Lambda$ be the set of subformulas of $\alpha$.
+Clearly, the truth-table $iPLV(\Lambda)$ for $\alpha$ only contains classical values $\mathbf{T}$ and $\mathbf{F}$, coinciding so with $iPV(\Lambda)$.
+It has, therefore, exactly $2^{k}$ elements.
+Moreover, by expanding this table with a column for $\alpha\to\alpha$, i.e., by considering $iPV(\Lambda\cup\{\alpha\to\alpha\})$, the last column only contains values $\mathbf{U}$ or $\mathbf{T}$.
+But then the rows assigning $\mathbf{U}$ to $\alpha\to\alpha$ are not supported by any other row, hence they are removed in the first round of Algorithm 1, arriving so to a set $iPLV(\Lambda\cup\{\alpha\to\alpha\})$ with $2^{k}$ elements.
 Definition 9.9.
-An intuitionistic truth table in $\mathcal{M}_{IPL}$ is an exhaustive list of all intuitionistic partial level valuations for a given finite domain closed under subformulas.
+An intuitionistic truth-table in $\mathcal{M}_{IPL}$ is an exhaustive list of every and only intuitionistic partial level valuations for a given finite domain closed under subformulas.
 Remark 9.10.
 Each of these partial valuations corresponds to a row of the table.
-The following is an example of the algorithm for searching the (intuitionistic) partial level valuations.
-[^5] This example, as the ones presented in Appendix A, were automatically generated by a Coq module for RNmatrices written by R.
-Leme, see  https://github.com/renatoleme/Forest We start with the full table for $\neg\neg(p\lor\neg p)$.
-Table 4: Initial table.
-Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
-(1) F U F U U
-(2) F U F U T
-(3) F U F T F
-(4) F T T F U
-(5) F T T F T
-(6) T F T F U
-(7) T F T F T
-Table 5: First cycle.
-Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
-(1) $\times$ 6, 7 $\times$ 4, 5, 6, 7 3
-(2) $\times$ 7 $\times$ 5, 7 $\times$
-(3) $\times$ $\emptyset$ $\times$ $\times$ $\times$
-(4) $\times$ $\times$ $\times$ $\times$ $\emptyset$
-(5) $\times$ $\times$ $\times$ $\times$ $\times$
-(6) $\times$ $\times$ $\times$ $\times$ $\emptyset$
-(7) $\times$ $\times$ $\times$ $\times$ $\times$
-The cells marked with the multiplication symbol ($\times$) in a collumn $\alpha$ of a partial valuation $v$ indicates that $v(\alpha)\neq\mathbf{U}$, so we can ignore that cell.
-The cells marked with the empty set symbol ($\emptyset$) in a collumn $\alpha$ of a partial valuation $v$ indicates that $v(\alpha)=\mathbf{U}$ and there is no other line $w$ to validate $v$, so that line must be removed.
-Table 6: Intermediate table.
-Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
-(1) F U F U U
-(2) F U F U T
-(5) F T T F T
-(7) T F T F T
-Table 7: Second cycle.
-Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
-(1) $\times$ 7 $\times$ 5, 7 $\emptyset$
-(2) $\times$ 7 $\times$ 5, 7 $\times$
-(5) $\times$ $\times$ $\times$ $\times$ $\times$
-(7) $\times$ $\times$ $\times$ $\times$ $\times$
-Table 8: Final table.
-Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
-(2) F U F U T
-(5) F T T F T
-(7) T F T F T
-Table 9: Validators.
-Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
-(2) $\times$ 7 $\times$ 5, 7 $\times$
-(5) $\times$ $\times$ $\times$ $\times$ $\times$
-(7) $\times$ $\times$ $\times$ $\times$ $\times$
-Note that, in the validators table, there is no occurrence of $\emptyset$. So, there is no further row to remove and we can stop the searching algorithm.
-4.2.3 Soundness and completeness of the intuitionistic truth tables
-Now, it will be shown that the generalized notion of truth tables for IPL proposed in Subsection 4.2.2 is sound and complete.
-As in the case of S4, this will be shown by proving first analyticity and co-analyticity w.r.t.
-level valuations in $\mathcal{M}_{IPL}$.
-In order to do this it is considered, adapting [4, Definition 3.12], the intermediary notion of intuitionistic partial’ level valuation in $\mathcal{R}(\mathcal{M}_{IPL})$ over $\Lambda$.
+Example
+In what follows, we present an example of a truth-table produced by Algorithm 1 which keep track of the dependencies of each partial valuation [^6] This example was automatically generated by a Rocq module for RNmatrices written by R.
+Leme, see https://github.com/renatoleme/Forest .
+The first table (Table 5(a)) is the full table for $\neg\neg(p\lor\neg p)$, a valid intuitionistic formula, and was obtained by applying the Nmatrix for $\mathbf{IPL}$ in the usual way.
+(a) Initial table.
+(b) First cycle of dependencies.
+(c) Intermediate table.
+(d) Second cycle of dependencies.
+(e) Final table.
+(f) Final dependency table.
+Figure 5: Example of executions of refinement algorithm.
+Note that each partial valuation is identified by its ID, which is an index $n\in\mathbb{N}$.
+The dependency tables are obtained in two steps: first, find every $v$ such that there is some $\alpha$ such that $v(\alpha)=\mathbf{U}$.
+Then, for each of these partial valuations, find every valuation $w$ such that $w(\alpha)=\mathbf{F}$ and, for every $\beta$, $w(\beta)=\mathbf{T}$ whenever $v(\beta)=\mathbf{T}$.
+In Figure 5, the tables on the left correspond to an execution of Algorithm 1 (see Figure 4), while the tables on the right represent the dependency tables.
+Whenever $v(\alpha)=\mathbf{U}$ occurs in a table on the left, the corresponding table on the right contains a set of supports (i.e., partial valuations $w$ such that $w(\alpha)=\mathbf{F}$ and $w(\beta)=\mathbf{T}$ whenever $v(\beta)=\mathbf{T}$) in the correspondent cell.
+If no such support exists for that case, then the corresponding partial valuation is removed in the subsequent execution of Algorithm 1, which can be observed in the transition between successive tables on the left.
+By removing rows $3$, 44, and $6$ from Table 5(a), we obtain Table 5(b).
+In this new table, the row $1$, which was previously supported by the row $3$, no longer has that support available.
+Running the algorithm again with the updated table, we find that the row $1$ must now also be removed.
+This results in Table 5(e).
+At this point, Table 5(f) contains no occurrences of $\emptyset$, which means that every instance of $\mathbf{U}$ is now adequately supported.
+Hence, there are no more rows to remove and the algorithm is terminated.
+Finally, by inspecting the resulting Table 5(e), we can confirm that the formula is indeed valid.
+Algorithm 1 is fundamentally the same as the one presented by Grätz ([12], Algorithm $3.4$), but is applied to the new RNmatrix for $\mathbf{IPL}$.
+As a result, it remains within the same complexity class using exponential space.
+However, it has a lower upper bound because the atomic valuation considers only two truth-values: $\mathbf{F}$ and $\mathbf{T}$.
+Section 10 provides a further discussion on this topic.
+9.2 Soundness and completeness of the intuitionistic truth-tables
+Now, it will be shown that the generalized notion of truth-tables for IPL proposed in Subsection 9.1 is sound and complete.
+This will be established by first proving analyticity and co-analyticity with respect to level valuations in $\mathcal{M}_{IPL}$.
+To this end, and adapting [12, Definition 3.12], we introduce the intermediary notion of intuitionistic partial level valuation in $\mathcal{R}(\mathcal{M}_{IPL})$ over $\Lambda$.
 Definition 9.11 (Intuitionistic partial’ level valuation over $\Lambda$).
-Let $\Lambda\in CS(\Omega)$. An intuitionistic partial valuation $\tilde{v}_{p}\in iPV(\Lambda)$ is an intuitionistic partial’ level valuation in $\mathcal{R}(\mathcal{M}_{IPL})$ over $\Lambda$ iff
+Let $\Lambda\in CS(\Omega)$ and $v_{p}\in iPV(\Lambda)$. Then, $\tilde{v}_{p}$ is an intuitionistic partial’ level valuation in $\mathcal{R}(\mathcal{M}_{IPL})$ over $\Lambda$ iff
   1.
@@
 The set of intuitionistic partial’ level valuations in $\mathcal{R}(\mathcal{M}_{IPL})$ over $\Lambda$ will be denoted by $iPLV^{\prime}(\Lambda)$.
+Definition 9.11 above is similar to Definition 9.3, except for two important aspects: (1) the domain $\Lambda$ of the partial valuation $\tilde{v}_{p}$ no longer needs to be finite, and (2) whenever there is a formula $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$, the support $w$ must be a level valuation.
+This difference is central for proving both co-analyticity and analyticity.
+Aspect (1) is crucial for analyticity (Lemma 9.18), since we aim to prove that any domain closed under subformulas (possibly infinite) can be extended to a domain over the entire set of formulas.
+Aspect (2), in turn, is essential for co-analyticity (Lemma 9.13), as it relies on the previous result obtained in Proposition 8.21, which establishes the existence of a certain level valuation under appropriate circumstances.
+The notion of $iPLV^{\prime}$ is well-suited for proof-theoretic purposes, but it is not sufficient for obtaining a decision procedure, since it depends on the set of level valuations.
+For this reason, we later prove that, for finite domains (which correspond to truth-tables), Definition 9.11 and Definition 9.3 are equivalent (see Proposition 9.19).
+This result, a direct consequence of Proposition 9.7, allows us to replace the notion of $iPLV^{\prime}$ with its finitary counterpart $iPLV$, showing that, in order to decide the validity of a formula (or argument), it is sufficient to restrict the analysis to the finite universe of its subformulas (see Theorem 9.22).
 Remark 9.12.
-Recall the predicate $\mathsf{P}_{\Lambda}^{i}(v,w,\alpha)$ introduced in Remark 4.27. Clearly,
-$\begin{array}[]{lll}iPLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},w,\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $w\in\mathcal{L}_{IPL}$}\big{)}\}.\end{array}$
+Recall the predicate $\mathsf{P}_{\Lambda}(v,w,\alpha)$ introduced in Remark 9.4. Clearly,
+$\begin{array}[]{lll}iPLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big(\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $w\in\mathcal{L}_{IPL}$}\big)\}.\end{array}$
 Lemma 9.13 (Co-analyticity lemma for $\mathcal{R}(\mathcal{M}_{IPL})$).
@@
 Proof.
-It is an adaptation of the proof of [4, Lemma 3.13].
-Suppose, by contradiction, that $\tilde{v}_{p}\notin iPLV^{\prime}(\Lambda)$.
-Let $\Gamma:=\{\beta\in\Lambda\mid\tilde{v}_{p}(\beta)=\mathbf{T}\}$.
-Then, there exists $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$ and, for every $w\in\mathcal{L}_{IPL}$, $w(\alpha)=\mathbf{F}$ implies that $w(\beta)\neq\mathbf{T}$ for some $\beta\in\Gamma$.
-By contraposition, this means that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha$.
-But $v(\beta)=\mathbf{T}$ for every $\beta\in\Gamma$ and $v(\alpha)=\mathbf{U}$, a contradiction.
-From this we infer that $\tilde{v}_{p}\in iPLV^{\prime}(\Lambda)$.
+Let $v\in\mathcal{L}_{IPL}$ and $\tilde{v}_{p}:=v_{|\Lambda}$.
+Clearly, $\tilde{v}_{p}\in iPV(\Lambda)$.
+Let $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$.
+Hence, $v(\alpha)=\mathbf{U}$.
+By Proposition 8.21, there exists $w\in\mathcal{L}_{IPL}$ such that $w(\alpha)=\mathbf{F}$ and $w(\beta)=\mathbf{T}$ for every $\beta$ such that $v(\beta)=\mathbf{T}$.
+In particular, $\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)$ holds.
+From this, $\tilde{v}_{p}\in iPLV^{\prime}(\Lambda)$ by Remark 9.12.
 ∎
+A key concern regarding truth-tables is whether a given partial valuation can always be extended to a full valuation, i.e., whether a domain closed under subformulas can be extended to a complete domain.
+This property, called analyticity, is central for proving the completeness of the method (see Theorem 9.22).
+To establish analyticity, we start with a domain over a set closed under subformulas and conservatively add new formulas to this domain.
+At the limit, we obtain the function given by the union of all such extensions, which corresponds to a function defined over the set of all formulas.
+However, a formula can be evaluated only if all of its subformulas are already evaluated.
+To ensure that this condition always holds, we consider a suitable ordering on the set of formulas.
+The technique will be entirely analogous to the case of S4, recall Remark 6.15.
+Definition 9.14.
+Given $n\geq 0$, the set of formulas over $\Omega$ with complexity less or equal than $n$ will be denoted by $For(\Omega)_{n}$.
+That is: $For(\Omega)_{n}=\{\alpha\in For(\Omega)\mid\textsf{c}(\alpha)\leq n\}$.
 Remark 9.15.
-Observe that, as it was done in Remark 3.22 for S4, it is possible to define an enumeration $\alpha_{1},\alpha_{2},\ldots\alpha_{m}\ldots$ (for $m\in\omega^{2}$) of $For(\Omega)$ such that $\textsf{c}(\alpha_{i})\leq\textsf{c}(\alpha_{j})$ if $i\leq j$ and, for every $i$ such that $\textsf{c}(\alpha_{i})>0$, if $\beta$ is a strict subformula of $\alpha_{i}$ then $\beta=\alpha_{j}$ for some $j<i$.
-Such an enumeration of $For(\Omega)$ can be defined as follows: every formula have and index in $\omega^{2}$, which is a denumerable ordinal, such that all the formulas with complexity 00 (which form a denumerable set) are placed first, with indexes in $I_{0}:=\omega$; after this, all the formulas with complexity $1$ (which form a denumerable set) are placed with an index in $I_{1}:=\omega\cdot 2\setminus\omega=\{\omega,\omega+1,\ldots\}$; in general, the formulas with complexity $n$ (which form a denumerable set) have an index in $I_{n}:=\omega\cdot(n+1)\setminus\omega\cdot n=\{\omega\cdot n,\omega\cdot n+1,\ldots\}$.
+Given the ordinal $\omega$, let $\omega^{2}:=\omega\cdot\omega$ be its ordinal multiplication.
+Observe that it is possible to define an enumeration $\alpha_{1},\alpha_{2},\ldots\alpha_{m},\ldots$ (for $m\in\omega^{2}$) of $For(\Omega)$ such that $\textsf{c}(\alpha_{i})\leq\textsf{c}(\alpha_{j})$ if $i\leq j$ and, for every $i$ such that $\textsf{c}(\alpha_{i})>0$, if $\beta$ is a strict subformula of $\alpha_{i}$ then $\beta=\alpha_{j}$ for some $j<i$.
+Such an enumeration of $For(\Omega)$ can be defined as follows: every formula have and index in $\omega^{2}$, which is a denumerable ordinal, such that all the formulas with complexity 0 (which form a denumerable set) are placed first, with indexes in $I_{0}:=\omega$; after this, all the formulas with complexity $1$ (which form a denumerable set) are placed with an index in $I_{1}:=\omega\cdot 2\setminus\omega=\{\omega,\omega+1,\ldots\}$; in general, the formulas with complexity $n$ (which form a denumerable set) have an index in $I_{n}:=\omega\cdot(n+1)\setminus\omega\cdot n=\{\omega\cdot n,\omega\cdot n+1,\ldots\}$.
+Based on the enumeration of $For(\Omega)$ given in the previous remark, consider the following sets:
+Definition 9.16.
+Let $\alpha_{1},\alpha_{2},\ldots\alpha_{m},\ldots$ (for $m\in\omega^{2}$) be an enumeration of $For(\Omega)$ as in Remark 9.15.
+For every $n,m\in\omega$ let $\bar{\Lambda}_{n}:=\Lambda\cup For(\Omega)_{n}$; $\bar{\Lambda}_{n}^{0}:=\bar{\Lambda}_{n}$; and $\bar{\Lambda}_{n}^{m+1}=\bar{\Lambda}_{n}^{m}\cup\{\alpha_{\omega\cdot(n+1)+m}\}$.
+Remark 9.17.
+By definition, $\bar{\Lambda}_{n}^{m+1}=\bar{\Lambda}_{n}\cup\{\alpha_{\omega\cdot(n+1)},\alpha_{\omega\cdot(n+1)+1},\ldots,\alpha_{\omega\cdot(n+1)+m}\}$, and $\alpha_{\omega\cdot(n+1)+m}\in\Lambda$ if and only if $\bar{\Lambda}_{n}^{m+1}=\bar{\Lambda}_{n}^{m}$.
+Hence, $\bar{\Lambda}_{n}^{m+1}$ is obtained from $\bar{\Lambda}_{n}$ by adding the first $m+1$ formulas (of the given enumeration) with complexity $n+1$, and so $\bar{\Lambda}_{n}^{m}$ adds at most $m$ formulas to $\bar{\Lambda}_{n}$, for $m\in\omega$.
+It is worth noting that $\bar{\Lambda}_{n}^{m}\in CS(\Omega)$ for every $n,m\in\omega$; in particular, $\bar{\Lambda}_{n}\in CS(\Omega)$ for every $n\in\omega$.
+Clearly $\bar{\Lambda}_{n}^{m}\subseteq\bar{\Lambda}_{n}^{m+1}$, $\bar{\Lambda}_{n+1}=\bigcup_{m\in\omega}\bar{\Lambda}_{n}^{m}$, and $For(\Omega)=\bigcup_{n\in\omega}\bar{\Lambda}_{n}$.
 Lemma 9.18 (Analyticity lemma for $\mathcal{R}(\mathcal{M}_{IPL})$).
@@
 Proof.
-It is an adaptation of the proof of [4, Lemma 3.14] to $\mathcal{R}(\mathcal{M}_{IPL})$.
+It is an adaptation of the proof of the Analyticity lemma for $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ (recall Lemma 6.18) to $\mathcal{R}(\mathcal{M}_{IPL})$.
 Suppose first that $\Lambda=For(\Omega)$.
 By induction on $n\in\omega$, it will be shown that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n}$ for every $n$.
-By Definition 4.25 of partial valuation in $\mathcal{M}_{IPL}$, $\tilde{v}_{p}\in Val_{+}(\mathcal{M}_{IPL})={\mathcal{L}}_{0}^{IPL}$.
-Suppose that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n}$ for a given $n\geq 0$, and let $\alpha\in For(\Omega)_{n+1}$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$.
-By Definition 4.30 of intuitionistic partial’ level valuation, there exists a level valuation $w$ in $\mathcal{L}_{IPL}$ (and so, $w\in\mathcal{L}^{IPL}_{n}$) such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},w,\alpha)$.
+By Definition 9.2 of partial valuation in $\mathcal{M}_{IPL}$, $\tilde{v}_{p}\in Val^{nd}_{+}(\mathcal{M}_{IPL})={\mathcal{L}}_{0}^{IPL}$.
+Suppose that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n}$ for a given $n\geq 0$, and let $\alpha\in For(\Omega)$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$.
+By Definition 9.11 of intuitionistic partial’ level valuation, there exists a level valuation $w$ in $\mathcal{L}_{IPL}$ (and so, $w\in\mathcal{L}^{IPL}_{n}$) such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)$.
 Since $\Lambda=For(\Omega)$, this means that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n+1}$.
-From this, $\tilde{v}_{p}\in\mathcal{L}_{IPL}$ and the result clearly holds.
+From this, $\tilde{v}_{p}\in\mathcal{L}_{IPL}$ and the result clearly holds by taking $v:=\tilde{v}_{p}$.
 Suppose now that $\Lambda\neq\emptyset$ is a proper subset of $For(\Omega)$.
@@
 The valuation $v$ will be defined by induction on the complexity $n$ of the formulas.
-Observe that $\Lambda\cup\mathcal{P}\in CS(\Omega)$ and, for every $n\in\omega$, $\bar{\Lambda}_{n}\in CS(\Omega)$ for $\bar{\Lambda}_{n}:=\Lambda\cup For(\Omega)_{n}$.
-In addition, for every $m\in\omega^{2}$ and $n\geq 0$, if $\alpha_{m}$ is such that $\textsf{c}(\alpha_{m})=n+1$, then $\bar{\Lambda}_{n}^{m}\in CS(\Sigma)$, where $\bar{\Lambda}_{n}^{m}=\bar{\Lambda}_{n}\cup\{\alpha_{m}\}$.
-Clearly $\bar{\Lambda}_{n+1}=\cup_{k\in\omega}\bar{\Lambda}_{n}^{\omega\cdot(n+1)+k}$.
-Base $n=0$: Define $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v(p)=\mathbf{F}$ if $p$ is a propositional variable which does not belong to $\Lambda$.
-Clearly, $v\in PLV^{\prime}(\Lambda\cup\mathcal{P})$ up to this stage.
-Inductive step: Assume that $v(\alpha)$ was defined for every $\alpha\in\bar{\Lambda}_{n}$ such that $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v\in iPLV^{\prime}(\bar{\Lambda}_{n})$ up to this stage, for a given $n\geq 0$ (Induction Hypothesis, IH).
-Let $\alpha=\alpha_{m}$ (for $m\in I_{n+1}$) such that $\alpha\not\in\Lambda$ and $\textsf{c}(\alpha)=n+1$.
-Observe that $v(\beta)$ was already defined, for every strict subformula $\beta$ of $\alpha$.
-Since $v\in iPLV^{\prime}(\bar{\Lambda}_{n})$ then, for every $\delta\in\bar{\Lambda}_{n}$ such that $v(\delta)=\mathbf{U}$ there exists $w\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\bar{\Lambda}_{n}}^{i}(v,w,\delta)$.
-Such a $w$ will be denoted by $w^{\delta}$ (observe that it is possible to have more that one $w^{\delta}$ for each $\delta$, and it is possible to have $w^{\delta}=w^{\gamma}$ for $\delta\neq\gamma$).
-Recall that $\bar{\Lambda}_{n}^{m}=\bar{\Lambda}_{n}\cup\{\alpha_{m}\}$ is closed under subformulas.
+More precisely, with terminology as in Definition 9.16 and from the observations in Remark 9.17, an extension $v_{n}^{m}$ of $\tilde{v}_{p}$ (i.e., $\tilde{v}_{p}\subseteq v_{n}^{m}$) will be defined for every $(n,w)\in\omega\times\omega$ such that $v_{n}^{m}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$ and $v_{i}^{j}\subseteq v_{n}^{m}$ if $(i,j)\leq(n,m)$, where: $(i,j)\leq(n,m)$ iff $i\leq n$ or $i=n$ and $j\leq m$ (and where $\omega\times\omega$ denotes, as usual, the Cartesian product of $\omega$ with itself).
+Base $n=m=0$: Observe that $\bar{\Lambda}_{0}=\bar{\Lambda}_{0}^{0}=\Lambda\cup\mathcal{P}$.
+Define $v_{0}^{0}(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v_{0}^{0}(p)=\mathbf{F}$ if $p$ is a propositional variable which does not belong to $\Lambda$.
+Clearly, $\tilde{v}_{p}\subseteq v_{0}^{0}$ and $v_{0}^{0}\in iPLV^{\prime}(\bar{\Lambda}_{0})=iPLV^{\prime}(\bar{\Lambda}_{0}^{0})$.
+Let $v_{0}:=v_{0}^{0}$.
+Inductive step: Assume that $v_{i}^{j}(\alpha)$ was defined for every $\alpha\in\bar{\Lambda}_{i}^{j}$ such that $v_{i}^{j}(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, $v_{i}^{j}\in iPLV^{\prime}(\bar{\Lambda}_{i}^{j})$, $v_{i}^{j}\subseteq v_{k}^{r}$ if $(i,j)\leq(k,r)\leq(n,m)$, for given $n,m\geq 0$ (Induction Hypothesis, IH).
+Now it will be shown how to extend $v_{n}^{m}$ to a function $v_{n}^{m+1}$ with domain $\bar{\Lambda}_{n}^{m+1}=\bar{\Lambda}_{n}^{m}\cup\{\alpha_{\omega\cdot(n+1)+m}\}$.
+That is, it will be shown how to define a value for $\alpha_{\omega\cdot(n+1)+m}$ while preserving the values assigned by $v_{n}^{m}$, in such a manner that $v_{n}^{m+1}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m+1})$.
+To start with, define $v_{n}^{m+1}(\beta)=v_{n}^{m}(\beta)$ for every $\beta\in\bar{\Lambda}_{n}^{m}$.
+Let $\alpha=\alpha_{\omega\cdot(n+1)+m}$.
+If $\alpha\in\Lambda$ then $\bar{\Lambda}_{n}^{m+1}=\bar{\Lambda}_{n}^{m}$.
+Then $v_{n}^{m+1}:=v_{n}^{m}$, and the required hypothesis are obviously satisfied by $v_{n}^{m+1}$.
+Suppose now that $\alpha\not\in\Lambda$.
+Observe that $v_{n}^{m+1}(\beta)=v_{n}^{m}(\beta)$ was already defined, for every strict subformula $\beta$ of $\alpha$.
+Since $v_{n}^{m}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$ then, for every $\delta\in\bar{\Lambda}_{n}^{m}$ such that $v_{n}^{m}(\delta)=\mathbf{U}$ there exists $w\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\bar{\Lambda}_{n}^{m}}(v_{n}^{m},w,\delta)$.
+Such a $w$ will be denoted by $w^{\delta}$ (observe that it is possible to have more than one $w^{\delta}$ for each $\delta$, and it is possible to have $w^{\delta}=w^{\gamma}$ for $\delta\neq\gamma$).
+Recall that $\bar{\Lambda}_{n}^{m}$ is closed under subformulas.
 There are three cases to analyze:
-(1) $\alpha=\neg\beta$ such that $\neg^{\mathcal{A}_{IPL}}v(\beta)=\{\mathbf{U},\mathbf{T}\}$, or $\alpha=\beta\to\gamma$ such that $\to^{\mathcal{A}_{IPL}}(v(\beta),v(\gamma))=\{\mathbf{U},\mathbf{T}\}$.
+(1) $\alpha=\neg\beta$ such that $\neg^{IPL}(v_{n}^{m}(\beta))=\{\mathbf{U},\mathbf{T}\}$, or $\alpha=\beta\to\gamma$ such that $\to^{IPL}(v_{n}^{m}(\beta),v_{n}^{m}(\gamma))=\{\mathbf{U},\mathbf{T}\}$.
 There are three subcases to analyze:
-(1.1) There is $\delta\in\bar{\Lambda}_{n}$ with $v(\delta)=\mathbf{U}$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=\mathbf{F}$.
-Define $v(\alpha)=\mathbf{U}$.
-Hence $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$ such that $w^{\alpha}=w^{\delta}$.
-(1.2) There is $\delta\in\bar{\Lambda}_{n}$ with $v(\delta)=\mathbf{U}$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=\mathbf{U}$.
-By Lemma 4.32, $v^{\prime}=w^{\delta}_{|\bar{\Lambda}_{n}^{m}}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$, hence there exists $w^{\prime\prime}\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\bar{\Lambda}_{n}^{m}}^{i}(v^{\prime},w^{\prime\prime},\alpha)$.
-Define $v(\alpha)=\mathbf{U}$.
-Hence, $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$ such that $w^{\alpha}=w^{\prime\prime}$.
-(1.3) For every $\delta\in\bar{\Lambda}_{n}$ such that $v(\delta)=\mathbf{U}$, $w^{\delta}(\alpha)=\mathbf{T}$ for every $w^{\delta}$.
-Define $v(\alpha)=\mathbf{T}$.
-Clearly, $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$.
-[…]
-[…]
-We have shown in cases (1)-(3) how to extend the domain of $v$ to the additional formula $\alpha_{m}$, for every $\alpha_{m}$ such that $\textsf{c}(\alpha_{m})=n+1$, showing that $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$.
-Consider now the simultaneous extension of $v$ with the values assigned to every $\alpha_{m}$ as prescribed above.
-We obtain so an extension of $v$ to $\bar{\Lambda}_{n+1}=\cup_{m}\bar{\Lambda}_{n}^{m}$.
-By the very definition of intuitionistic partial’ level valuations it follows that $v\in iPLV^{\prime}(\bar{\Lambda}_{n+1})$, and $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$.
-This completes the inductive step.
+(1.1) There is $\delta\in\bar{\Lambda}_{n}^{m}$ with $v_{n}^{m}(\delta)=\mathbf{U}$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=\mathbf{F}$.
+Define $v_{n}^{m+1}(\alpha)=\mathbf{U}$.
+Hence $v_{n}^{m+1}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m+1})$ such that $w^{\alpha}=w^{\delta}$.
+(1.2) There is $\delta\in\bar{\Lambda}_{n}^{m}$ with $v_{n}^{m}(\delta)=\mathbf{U}$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=\mathbf{U}$.
+By Proposition 8.21, there exists $w^{\prime\prime}\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{For(\Omega)}(w^{\delta},w^{\prime\prime},\alpha)$.
+Define $v_{n}^{m+1}(\alpha)=\mathbf{U}$.
+Hence, $v_{n}^{m+1}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m+1})$ such that $w^{\alpha}=w^{\prime\prime}$.
+(1.3) For every $\delta\in\bar{\Lambda}_{n}^{m}$ such that $v_{n}^{m}(\delta)=\mathbf{U}$, $w^{\delta}(\alpha)=\mathbf{T}$ for every $w^{\delta}$.
+Define $v_{n}^{m+1}(\alpha)=\mathbf{T}$.
+Clearly, $v_{n}^{m+1}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m+1})$.
+(2) $\alpha=\neg\beta$ such that $\neg^{IPL}(v_{n}^{m}(\beta))=\{\mathbf{F}\}$, or $\alpha=\beta\#\gamma$ such that $^{IPL}(v_{n}^{m}(\beta),v_{n}^{m}(\gamma))=\{\mathbf{F}\}$ (for some $#\in\{\vee,\land,\to\}$).
+Define $v_{n}^{m+1}(\alpha)=\mathbf{F}$.
+It is clear that $v_{n}^{m+1}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m+1})$.
+(3) $\alpha=\beta\#\gamma$ such that $^{IPL}(v_{n}^{m}(\beta),v_{n}^{m}(\gamma))=\{\mathbf{T}\}$ (for some $#\in\{\vee,\land,\to\}$).
+Define $v_{n}^{m+1}(\alpha)=\mathbf{T}$.
+It is clear that $v_{n}^{m+1}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m+1})$.
+We have shown in cases (1)-(3) how to extend the domain of $v_{n}^{m}$ to the additional formula $\alpha_{\omega\cdot(n+1)+m}\in\bar{\Lambda}_{n}^{m+1}$, showing that the resulting function $v_{n}^{m+1}$ is in $iPLV^{\prime}(\bar{\Lambda}_{n}^{m+1})$, and $\tilde{v}_{p}\subseteq v_{i}^{j}\subseteq v_{k}^{r}$, for every $(i,j)\leq(k,r)\leq(n,m+1)$.
+That is, we show how to increase the superscript $m$.
+In order to increase the subscript $n$, let $v_{n+1}=v_{n+1}^{0}:=\bigcup_{m\in\omega}v_{n}^{m}$.
+By the procedure described above, it is immediate to see that $v_{n+1}^{0}$ is in $iPLV^{\prime}(\bigcup_{m\in\omega}\bar{\Lambda}_{n}^{m})=iPLV^{\prime}(\bar{\Lambda}_{n+1})=iPLV^{\prime}(\bar{\Lambda}_{n+1}^{0})$, and $\tilde{v}_{p}\subseteq v_{i}^{j}\subseteq v_{k}^{r}$, for every $(i,j)\leq(k,r)\leq(n+1,0)$.
+By repeating the process, it is possible to define $v_{n}^{m}$ for every $(n,m)\in\omega\times\omega$ with the required properties.
+Finally, let $v:=\bigcup_{n\in\omega}v_{n}$.
+From the manner in which the construction was made, it follows that $v$ is in $iPLV^{\prime}(\bigcup_{n\in\omega}\bar{\Lambda}_{n})=iPLV^{\prime}(For(\Omega))$, and $\tilde{v}_{p}\subseteq v$.
+By the first part of the proof, we conclude that $v\in\mathcal{L}_{IPL}$.
 ∎
@@
 Suppose that $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$.
-Then, there exists $w\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},w,\alpha)$.
+Then, there exists $w\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)$.
 Let $\tilde{w}_{p}=w_{|\Lambda}$.
-By Lemma 4.32, $\tilde{w}_{p}\in iPLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}_{p},\alpha)$.
-Now, suppose that there exists $\tilde{w}^{\prime}_{p}\in iPLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}^{\prime}_{p},\alpha)$.
+By Lemma 9.13, $\tilde{w}_{p}\in iPLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)$.
+Now, suppose that there exists $\tilde{w}^{\prime}_{p}\in iPLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}^{\prime}_{p},\alpha)$.
 By Lemma 9.18, there exists $v^{\prime}\in\mathcal{L}_{IPL}$ such that the restriction of $v^{\prime}$ to $\Lambda$ coincides with $\tilde{w}^{\prime}_{p}$.
-Hence, $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},v^{\prime},\alpha)$.
+Hence, $\mathsf{P}_{\Lambda}(\tilde{v}_{p},v^{\prime},\alpha)$.
 This shows that
-$\begin{array}[]{lll}iPLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in iPLV^{\prime}(\Lambda)$}\big{)}\}.\end{array}$ From this, being $iPV(\Lambda)$ finite, it follows that $iPLV(\Lambda)=iPLV^{\prime}(\Lambda)$.
+$\begin{array}[]{lll}iPLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big(\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in iPLV^{\prime}(\Lambda)$}\big)\}.\end{array}$ From this, it follows that $iPLV(\Lambda)=iPLV^{\prime}(\Lambda)$, by Proposition 9.7.
 ∎
+Remark 9.20.
+In Proposition 9.7 the set $iPLV(\Lambda)$ was characterized constructively, as the output of an algorithm (Algorithm 1).
+In this subsection, we define an alternative subset $iPLV^{\prime}(\Lambda)\subseteq iPV(\Lambda)$ in a predicative way.
+This subset connects the RNmatrix semantics for $\mathbf{IPL}$ (based on level valuations) to the truth-table procedure given by $iPLV(\Lambda)$.
+Analyticity and co-analyticity of $iPLV^{\prime}(\Lambda)$ for level valuations allow us to show that, for a finite $\Lambda$, $iPLV^{\prime}(\Lambda)$ satisfies the condition of Remark 9.4.
+By uniqueness (Proposition 9.7), $iPLV^{\prime}(\Lambda)$ must coincide with $iPLV(\Lambda)$, i.e., with the truth-table generated by $\Lambda$.
+But then we obtain analyticity and co-analyticity of $iPLV(\Lambda)$ for level valuations, which implies the soundness and completeness of truth-tables for $\mathbf{IPL}$, see Theorem 9.22.
 Definition 9.21.
@@
 In particular, $\emptyset\models_{iPLV}\varphi$, denoted simply by $\models_{iPLV}\varphi$, if $\tilde{v}_{p}(\varphi)=\mathbf{T}$ for every $\tilde{v}_{p}\in iPLV(\Lambda)$, where $\Lambda$ is the set of subformulas of $\varphi$.
+Finally, we establish the correspondence between the restricted Nmatrix for $\mathbf{IPL}$ and the intuitionistic truth-tables:
 Theorem 9.22 (Soundness and completeness of intuitionistic truth-tables w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$).
@@
 ∎
-Proposition 4.38.
-Let $\Gamma\cup\{\alpha,\beta\}\subseteq For(\Omega)$. If $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$ then $\Gamma,\alpha\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$.
-Proof.
-Assume that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$, and let $v\in\mathcal{L}_{IPL}$ such that $v(\gamma)=v(\alpha)=\mathbf{T}$ for every $\gamma\in\Gamma$.
-By hypothesis, $v(\alpha\to\beta)=\mathbf{T}$.
-By definition of $\to^{\mathcal{A}_{IPL}}$, and given that $v\in Val(\mathcal{M}_{IPL})$, it follows from this that $v(\beta)=\mathbf{T}$.
-This shows that $\Gamma,\alpha\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$.
-∎
-Theorem 4.39 (Soundness of $\mathcal{H}_{IPL}$ w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$).
-For every $\Gamma$ and $\varphi$: $\Gamma\vdash_{IPL}\varphi$ implies that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$.
-Proof.
-Since the consequence relation $\mathcal{H}_{IPL}$ of the Hilbert calculus for IPL is finitary and satisfies the deduction metatheorem, and taking into account Proposition 4.38, it is enough to prove the result for $\Gamma=\emptyset$.
-We prove first the following two facts:
-Fact 1: If $\varphi$ is an instance of an axiom in $\mathcal{H}_{IPL}$ then $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$.
-Indeed, let $\varphi$ be an instance of an axiom in $\mathcal{H}_{IPL}$, and let $\Lambda$ be the set of subformulas of $\varphi$.
-Then, $\Lambda\in FCS(\Omega)$.
-It is routine to check that the truth-table for $\varphi$ over $\Lambda$ validates $\varphi$ (see Appendix A).
-That is: for every intuitionistic partial level valuation $\tilde{v}_{p}\in iPLV(\Lambda)$, $\tilde{v}_{p}(\varphi)=\mathbf{T}$.
-By Theorem 4.37 it follows that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$.
-Fact 2: If $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha$ and $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$ then $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$.
-Indeed, assume that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha$ and $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$.
-If $v\in\mathcal{L}_{IPL}$ then $v(\alpha)=v(\alpha\to\beta)=\mathbf{T}$.
-By definition of $\to^{\mathcal{A}_{IPL}}$, and by the fact that $v\in Val(\mathcal{M}_{IPL})$, it follows that $v(\beta)=\mathbf{T}$.
-This means that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$.
-Now, suppose that $\vdash_{IPL}\varphi$.
-By induction on the length of a derivation of $\varphi$ in $\mathcal{H}_{IPL}$, and by Fact 1 and Fact 2 it is immediate to prove that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$.
-∎
-Corollary 4.39.1 (Finitariness for $\mathcal{R}(\mathcal{M}_{IPL})$).
-Let $\Gamma\cup\{\varphi\}\subseteq For(\Omega)$.
-If $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$ then $\Gamma_{0}\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$ for some finite subset $\Gamma_{0}$ of $\Gamma$.
-Proof.
-It follows from Theorems 4.23, 4.39 and the fact that $\mathcal{H}_{IPL}$ is finitary. ∎
-Corollary 4.39.2 (Deduction metatheorem for $\mathcal{R}(\mathcal{M}_{IPL})$).
-Let $\Gamma\cup\{\alpha,\beta\}\subseteq For(\Omega)$. If $\Gamma,\alpha\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$ then $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$.
-Proof.
-It follows from Theorems 4.23, 4.39 and the fact that $\mathcal{H}_{IPL}$ satisfies the deduction metatheorem. ∎
-Corollary 4.39.3 (Deduction-detachment metatheorem for intuitionistic truth-tables).
+Corollary 9.22.1 (Deduction metatheorem for intuitionistic truth-tables).
 Let $\Gamma\cup\{\alpha,\beta\}$ be a finite subset of $For(\Omega)$. Then, $\Gamma,\alpha\models_{iPLV}\beta$ if, and only if, $\Gamma\models_{iPLV}\alpha\to\beta$.
 Proof.
-It follows from Theorem 4.37, Proposition 4.38 and Corollary 4.39.2. ∎
-It is worth noting that the latter result is useful to reduce the complexity of the truth-tables: instead of generating the (reduced) truth table for a formula $\delta=\beta_{1}\to(\beta_{2}\to(\ldots\to(\beta_{n}\to\varphi)\ldots))$ to be tested as a tautology, it suffices generating the (reduced) truth-table for $\{\beta_{1},\ldots,\beta_{n},\varphi\}$ and then determine if, in every row such that $\beta_{1},\ldots,\beta_{n}$ receive simultaneously the value $\mathbf{T}$, $\varphi$ also receives the value $\mathbf{T}$.
+It follows from Theorem 9.22 and Corollary 8.19.2. ∎
+It is worth noting that the latter result is useful to reduce the complexity of the truth-tables: instead of generating the (reduced) truth-table for a formula $\delta=\beta_{1}\to(\beta_{2}\to(\ldots\to(\beta_{n}\to\varphi)\ldots))$ to be tested as a tautology, it suffices generating the (reduced) truth-table for $\{\beta_{1},\ldots,\beta_{n},\varphi\}$ and then determine if, in every row such that $\beta_{1},\ldots,\beta_{n}$ receive simultaneously the value $\mathbf{T}$, $\varphi$ also receives the value $\mathbf{T}$.
 This reduces the number of rows required to test $\delta$.
-Corollary 4.39.4 (Soundness and completeness of $\mathcal{H}_{IPL}$ w.r.t. intuitionistic truth-tables).
+Finally, we arrive to the main result of the paper, recalling that LI is a standard sequent calculus for $\mathbf{IPL}$:
+Theorem 9.23 (Soundness and completeness of LI w.r.t. intuitionistic truth-tables).
 Let $\Gamma\cup\{\varphi\}$ be a finite subset of $For(\Omega)$. Then: $\Gamma\vdash_{IPL}\varphi$ if, and only if, $\Gamma\models_{iPLV}\varphi$.
 Proof.
-It follows from Theorems 4.39, 4.23 and 4.37. ∎
-The table is never empty.
-Lemma 4.40.
-Let $\Delta\in CS(\Sigma)$ and $\beta\in For(\Sigma)$.
-If $v_{p}:\Delta\to\mathcal{V}$ is $iPLV$ and $sub(\beta)\subseteq\Delta$, then $v_{p}$ has at least one extension $v^{\prime}_{p}:\Delta\cup\{\beta\}\to\mathcal{V}$ such that $v^{\prime}_{p}$ is $iPLV$.
-Proof.
-By hypothesis, $v_{p}$ is iPLV.
-Then, by analiticity of $\mathcal{R}(\mathcal{M}_{IPL})$, there is a level valuation $v:For(\Sigma)\to\mathcal{V}$ such that $v_{p}=v_{\mid\Delta}$, i.e, for every $\alpha\in\Delta$, $v(\alpha)=v_{p}(\alpha)$.
-More then that, $v$ is defined for $\beta$.
-Now, given that $sub(\beta)\subseteq\Delta$ and $\Delta\in CS(\Sigma)$, then $\Delta\cup\{\beta\}\in CS(\Sigma)$ and, by co-analiticity of $\mathcal{R}(\mathcal{M}_{IPL})$, $v^{\prime}_{p}=v_{\mid\Delta\cup\{\beta\}}$ is iPLV.
-∎
-5 Complexity
-Given a non-deterministic matrix $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ defined over a language $\mathcal{L}$ and a formula $\alpha\in\mathcal{L}$, we can estimate the total number of rows of the final truth table for $\alpha$ via the following upper bound.
-Let $sub(\alpha)\subseteq\mathcal{L}$ be the set of all sub-formulas of $\alpha$, $atoms(\alpha)\subseteq sub(\alpha)$ the set of sub-formulas whose complexity is 00, and $\kappa\geq 1$ the branching factor.
-Then, the maximum number of rows of the truth table for $\alpha$ in $\mathcal{M}$ is given by
-Definition 5.1.
-$ub_{\mathcal{M}}(\alpha)=\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}\times|\mathcal{V}|^{|atoms(\alpha)|}$ (1)
-In 5.1, we assume that the branching factor is the maximum degree of non-determinism. More formally,
-Definition 5.2.
-Let $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ be some nmatrix.
-The branching factor of $\mathcal{M}$ is the highest $\kappa$ such that, for every n-ary ~$\tilde{\circ}\in\mathcal{O}$, every formula $\alpha_{1},\ldots,\alpha_{n}$ and every valuation $v\in Val(\mathcal{M})$, |~$|\tilde{\circ}(v(\alpha_{1}),\ldots,v(\alpha_{n}))|\leq\kappa$.
-For example, in the case of $\mathcal{M}_{IPL}$, we have that disjunction and conjunction have branching factor $1$, while implication and negation have branching factor $2$.
-Therefore, in this case, $\kappa=2$.
-Remark 5.3.
-If we consider the valuations in the nmatrix as branches in trees, we can say that $|\mathcal{V}|^{|atoms(\alpha)|}$ gives the number of trees, while $\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}$ gives the (maximum) number of branches of each tree.
-We can also obtain a lower bound by assuming $\kappa=1$, which is the branching factor of deterministic matrices. This gives us
-Definition 5.4.
-$lb_{\mathcal{M}}(\alpha)=|\mathcal{V}|^{|atoms(\alpha)|}$ (2)
-From (5.1) and (5.4), it follows that
-$ub_{\mathcal{M}}(\alpha)=\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}\times lb_{\mathcal{M}}(\alpha)$ (3)
-$p$ $p$
-$\mathcal{V}_{1}$ $\mathcal{V}_{1}$
-$\vdots$ $\vdots$
-$\mathcal{V}_{n}$ $\mathcal{V}_{n}$
-Table 10: Base case.
-Now we shall prove that these are faithful lower and upper bounds for the number of rows.
-Theorem 5.5.
-Let $\mathcal{M}$ be a nmatrix and $rows(\alpha)$ the set of every row of the truth table for some formula $\alpha$. Then
-$lb_{\mathcal{M}}(\alpha)\leq|rows(\alpha)|\leq ub_{\mathcal{M}}(\alpha)$
-Proof.
-Let $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ be the nmatrix for the language of $\alpha$, where $\mathcal{V}=\{\mathcal{V}_{1},\ldots,\mathcal{V}_{n}\}$.
-Induction on the complexity of $\alpha$.
-Base case If $\alpha=p$, the truth table is depicted in Table 10, which has $|\mathcal{V}|=n$ rows.
-Now, $|sub(\alpha)|=|atoms(\alpha)|=1$.
-For this reason, $lb_{\mathcal{M}}(\alpha)=|\mathcal{V}|^{1}=|\mathcal{V}|$ and $ub(\alpha)=\kappa^{0}\times|\mathcal{V}|^{1}=|\mathcal{V}|$.
-Therefore, $lb_{\mathcal{M}}(p)=|rows(p)|=ub_{\mathcal{M}}(p)$.
-IH If $c(\alpha)<c(\alpha^{\prime})$, then the theorem holds for $\alpha$.
-$\alpha=\neg\beta$
-Let $rows(\beta)=\{v_{0},\ldots,v_{n}\}$ be the list of rows of the nmatrix for $\beta$.
-Because $atoms(\beta)=atoms(\neg\beta)$, we conclude that $lb_{\mathcal{M}}(\neg\beta)=lb_{\mathcal{M}}(\beta)$.
-Now, observe that $|sub(\neg\beta)|=|sub(\beta)|+1$.
-From this, we get $ub_{\mathcal{M}}(\neg\beta)=\kappa^{|sub(\beta)|-|atoms(\neg\beta)|+1}\times lb_{\mathcal{M}}(\neg\beta)=ub_{\mathcal{M}}(\beta)\times\kappa$.
-By the definition of branching factor, for every $v_{n}\in rows(\beta)$, |¬~$|\tilde{\neg}(v_{n}(\beta))|\leq\kappa$.
-Therefore, $|rows(\neg\beta)|\leq|rows(\beta)|\times\kappa$.
-Now, by (IH), $lb_{\mathcal{M}}(\beta)\leq|rows(\beta)|\leq ub_{\mathcal{M}}(\beta)$.
-Therefore, $lb_{\mathcal{M}}(\neg\beta)\leq|rows(\neg\beta)|\leq ub_{\mathcal{M}}(\neg\beta)$.
-[…]
-Let $rows(\beta)=\{v_{0},\ldots,v_{n}\}$ be the nmatrix of $\beta$ and let $rows(\gamma)=\{w_{0},\ldots,w_{n}\}$ be the nmatrix of $\gamma$. By (IH), we get the following
-$lb_{\mathcal{M}}(\beta)\leq|rows(\beta)|\leq ub_{\mathcal{M}}(\gamma)$ (4)
-$lb_{\mathcal{M}}(\gamma)\leq|rows(\gamma)|\leq ub_{\mathcal{M}}(\gamma)$ (5)
-Using 4 and 5 and some laws of algebra, we get
-$lb_{\mathcal{M}}(\beta)\times lb_{\mathcal{M}}(\gamma)\leq rows(\beta)\times rows(\gamma)\leq ub_{\mathcal{M}}(\beta)\times ub_{\mathcal{M}}(\gamma)$
-Now, to construct the truth table for $\alpha$, we combine $rows(\beta)$ with $rows(\gamma)$.
-In the worst case, $sub(\beta)\cap sub(\gamma)=\emptyset$.
-In this case, we append, to each row of $rows(\beta)$, a row of $rows(\gamma)$.
-Then, each row creates $\kappa-1$ fresh rows.
-Therefore,
-$|rows(\alpha)|\leq|rows(\beta)|\times|rows(\gamma)|\times\kappa$ (6)
-Moreover,
-$|sub(\alpha)|$ $\leq|sub(\beta)|+|sub(\gamma)|+1$ (7)
-$|atoms(\alpha)|$ $\leq|atoms(\beta)|+|atoms(\gamma)|$ (8)
-Therefore,
-$|\mathcal{V}|^{|atoms(\alpha)|}$ $\leq|\mathcal{V}|^{|atoms(\beta)|+|atoms(\gamma)|}$
-$=|\mathcal{V}|^{|atoms(\beta)|}\times|\mathcal{V}|^{|atoms(\gamma)|}$
-$=lb_{\mathcal{M}}(\beta)\times lb_{\mathcal{M}}(\gamma)$
-and
-$\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}\times|\mathcal{V}|^{|atoms(\alpha)|}$
-$\leq\kappa^{(|sub(\beta)|+|sub(\gamma)|+1)-(|atoms(\beta)|+|atoms(\gamma)|)}\times|\mathcal{V}|^{|atoms(\alpha)|}$
-$=\kappa^{(|sub(\beta)|-|atoms(\beta)|)+(|sub(\gamma)|-|atoms(\gamma)|)+1}\times\left(|\mathcal{V}|^{|atoms(\beta)|}\times|\mathcal{V}|^{|atoms(\gamma)|}\right)$
-$=\left(\kappa^{(|sub(\beta)|-|atoms(\beta)|)}\times|\mathcal{V}|^{|atoms(\beta)|}\right)\times\left(\kappa^{(|sub(\gamma)|-|atoms(\gamma)|)}\times|\mathcal{V}|^{|atoms(\gamma)|}\right)\times\kappa$
-$=ub_{\mathcal{M}}(\beta)\times ub_{\mathcal{M}}(\gamma)\times\kappa$
-Hence,
-$lb_{\mathcal{M}}(\beta\#\gamma)\leq|rows(\beta\#\gamma)|\leq ub_{\mathcal{M}}(\beta\#\gamma)$
-5.1 Algorithms
-Given a formula $\varphi$, to decide if $\varphi$ is valid we have to proceed in two steps: (1) construct the initial matrix by applying the definitions of the multi-operations, and, then, (2) refine the matrix by removing every spurious row.
-In the previous section, we have proposed a way to measure the space complexity of (1).
-Now, we will find the time complexity of (2).
-5.1.1 Pseudo-codes
-In what follows, given an ordered pair $\langle a;b\rangle$,
-$\pi_{1}(\langle a;b\rangle)=a$
-and
-$\pi_{2}(\langle a;b\rangle)=b$
-Moreover, $\mathcal{L}$ denotes the set of every intuitionistic formula and $\mathcal{V}=\{\textbf{F},\textbf{U},\textbf{T}\}$.
-Algorithm 1 Compatibility procedure.
-This algorithm receives two valuations, $v$ and $w$, and check if, for every evaluated formula $\alpha$, if $v(\alpha)=\textbf{T}$ then $w(\alpha)=\textbf{T}$ ($v$ is compatible with $w$).
-We assume that every valuation list is equally sorted.
-procedure isCompatible($v,w:list(\mathcal{L}\times\mathcal{V})$)
-     $i\leftarrow 0$
-     while $i<length(v)$ do
-         $a\leftarrow\pi_{2}(v[i])$triangleright Get the value in $v$ on index $i$.
-         if $a=$ T then
-              $b\leftarrow\pi_{2}(w[i])$triangleright Get the value in $w$ on the same index.
-              if $b=$ T then
-                  continue
-              else
-                  return $false$
-              end if
-         else
-              continue
-         end if
-         $i\leftarrow i+1$
-     end while
-     return $true$
-end procedure
-Algorithm 2 Search algorithm.
-This algorithm receives a formula $A\in\mathcal{L}$, a (partial) valuation $v:\mathcal{L}\rightarrow\mathcal{V}$ and a matrix.
-It runs over the entire matrix looking for a valuation $w$ that validates $v$, that is, a $w$ such that $v$ is compatible with and such that $w(A)=\textbf{F}$.
-procedure validate($A:\mathcal{L},v:list(\mathcal{L}\times\mathcal{V}),matrix:list(list(\mathcal{L}\times\mathcal{V}))$)
-     $i\leftarrow 0$
-     while $i<length(matrix)$ do
-         $w\leftarrow matrix[i]$triangleright Get a valuation function.
-         $j\leftarrow 0$
-         while $j<length(w)$ do
-              $A^{\prime}\leftarrow\pi_{1}(w[j])$triangleright Get a valuated formula.
-              if $A^{\prime}=A$ then
-                  $wA\leftarrow\pi_{2}(w[j])$triangleright Get $w(A)$.
-                  if $wA=$ F then
-                       $compatible\leftarrow isCompatible(v,w)$
-                       if $compatible$ then
-                           return $true$triangleright In this case, $w$ validates $v$.
-                       else
-                           continue
-                       end if
-                  else
-                       continue
-                  end if
-              else
-                  continue
-              end if
-              $j\leftarrow j+1$
-         end while
-         $i\leftarrow i+1$
-     end while
-     return $false$triangleright In this case, no $w$ validates $v$.
-end procedure
-Algorithm 3 Refinement algorithm. In this pseudo-code, we assume that $remove$ is a function that receives a matrix and a row, returning the matrix minus the row.
-procedure refineMatrix($matrix:list(list(\mathcal{L}\times\mathcal{V}))$)
-     $i\leftarrow 0$
-     $newmatrix\leftarrow matrix$
-     while $i<length(matrix)$ do
-         $v\leftarrow matrix[i]$triangleright Get a valuation function.
-         $j\leftarrow 0$
-         while $j<length(v)$ do
-              $A\leftarrow\pi_{1}(v[j])$triangleright Get a valuated formula.
-              $vA\leftarrow\pi_{2}(v[j])$triangleright Get $v(A)$.
-              if $vA=$ U then
-                  $isValid\leftarrow validate(A,v,newmatrix)$
-                  if $isValid$ then
-                       continue $\triangleright$ In this case, we keep the matrix unaltered.
-                  else
-                       $newmatrix\leftarrow remove(v,newmatrix)$
-                       break
-                  end if
-              else
-                  continue
-              end if
-              $j\leftarrow j+1$
-         end while
-         $i\leftarrow i+1$
-     end while
-     return $newmatrix$triangleright Return the refined matrix.
-end procedure
-Algorithm 4 In the best case, the refinement algorithm does not remove any row.
-However, if it does drag at least one row, we have to rerun it, given that it could have been the case that the removed row was used to validate some other rows.
-For this reason, in the worst case, the refinement algorithm removes every row and one row at once.
-This auxiliar procedure should address this possibility.
-procedure refineAux($matrix:list(list(\mathcal{L}\times\mathcal{V}))$)
-     $i\leftarrow 0$
-     $nrows\leftarrow length(matrix)$
-     while $i<nrows$ do
-         $matrix\leftarrow refineMatrix(matrix)$
-         $i\leftarrow i+1$
-     end while
-     return $matrix$
-end procedure
-Theorem 5.6.
-Given as input a matrix with $n$ rows and $m$ columns, the time complexity of refineAux is $O(n^{3}*m^{3})$.
-Proof.
-Given a list $v$ with length equal to $m$, in the worst case, the loop in $isCompatible$ runs $m$ times.
-Therefore, the complexity of this function is $O(m)$.
-Now, while, in the worst case, the external loop in $validate$ runs $n$ times, where $n$ is the number of rows in the matrix, the internal loop runs $m$ times.
-Because this functions calls $n*m$ times the function $isCompatible$, then the complexity of $validate$ is $O(n*m)*O(m)=O(n*m^{2})$.
-In the case of $refineMatrix$, the external loop runs $n$ times and the internal loop runs $m$ times.
-Therefore, it calls $validate$ $n*m$ times, which gives us $O(n*m)*O(n*m^{2})=O(n^{2}*m^{3})$.
-Finally, the function $refineAux$ calls $refineMatrix$ $n$ times.
-Therefore, the complexity of $refineAux$ is $O(n)*O(n^{2}*m^{3})=O(n^{3}*m^{3})$.
-∎
-5.2 Complexity of $\mathcal{R}(\mathcal{M}_{IPL})$
-In the valuation of a formula in $\mathcal{M}_{IPL}$, we consider only $\mathbf{F}$ and $\mathbf{T}$ in the atomic cases. Given that, in this case, $\kappa=2$, we can simplify the upper bound to:
-$ub_{\mathcal{M}_{IPL}}(\alpha)$ $=2^{(|sub(\alpha)|-|atoms(\alpha)|)}\times 2^{|atoms(\alpha)|}$ (9)
-$=2^{|sub(\alpha)|}$ (10)
-Now, given a formula $\varphi\in For(\Omega)$, assume that $|sub(\varphi)|=m$ and that $ub_{\mathcal{M}_{IPL}}(\varphi)=n$. Then
-$n^{3}$ $=(2^{m})^{3}=8^{m}$
-Therefore, given a formula $\varphi$, the time complexity to compute the truth table (in the worst case) in $\mathcal{R}(\mathcal{M}_{IPL})$ is given by $O(8^{m}*m^{3})$, which is in EXPTIME.
-6 Related work and concluding remarks
-In this paper, we proposed a new decision procedure for intuitionistic propositional logic IPL based on a 3-valued restricted Nmatrix semantics.
-Starting from the nice decision procedure for S4 obtained by Grätz in [4] based on a 3-valued RNmatrix, and taking advantage of the conservative translation of Gödel from IPL to S4, it is immediate to obtain an algorithm to decide validity in IPL.
-By means of an intermediate translation mapping from the language of IPL into the language of S4 we define a 3-valued Nmatrix which is the basis for an RNmatrix semantics for IPL.
-Moreover, by adapting Grätz’s algorithm for S4, a very simple algorithm for deleting spurious rows in the non-deterministic truth-table generated by the Nmatrix of IPL for a given formula is also provided.
-The sound and completeness proofs for the RNmatrix and for the obtained truth-tables algorithm for IPL, are presented with full technical details.
-These proofs were adapted from new ones for S4 we obtained in Subsection 3.2.
-On the one hand, we reduced the original Nmatrix for S4 presented by Grätz.
-On the other, we expanded it by adding the truth-tables for conjunction and disjunction.
-This required the extension and modification of the proofs originally given in [4].
-It is worth noting that the Nmatrix for IPL we introduced in Definition 4.10 has a clear and intuitive interpretation in terms of constructibility, by observing the definition of the canonical valuation associated to a $\varphi$-saturated set in IPL given in Definition 4.19, as well as the observations made in Remark 4.20.
-The idea of using finite non-deterministic truth-tables with restricted valuations for characterizing intuitionistic logic IPL is not new.
-Andréa Loparic already obtained in 1977, announced in [7] and finally published in [8], a semantical characterization of Johannson-Kolmogorov minimal logic and intuitionistic propositional logic IPL in terms of a $2$-valued, non truth-functional valuation semantics which induces a decision procedure for these logics.
-Similar to Grätz method (and to the adaptation to IPL proposed here), Loparic’s approach is based on a criterion for distinguish between that she calls semivaluations and valuations, the latter being the useful ones.
-This would correspond to level valuations, since they are exactly the semivaluations associate to $\varphi$-saturated sets, and the ones which characterizes the logic.
-The associated algorithm, as in the cases of S4 and IPL presented in this paper, generates a non-deterministic truth-table for a given formula in which, by using some formal criteria, some rows are deleted, and so the formula is tautology of the given logic if and only if it gets the value 1 in all the undeleted rows.
-More recently, Alejandro Solares-Rojas proposed a 3-valued Nmatrix semantics for IPL based on intuitionistic Kripke models and the novel notion of depth-bounded approximations (see [11, Chapter 5]).
-Within this semantic approach to IPL, a 3-valued Nmatrix – whose truth-values are intuitionistic truth (1), falsity (0) and indeterminacy ($u$) – gives the meaning of the connectives in terms of information that a given agent holds.
-The valuations over the Nmatrix intend to model actual information states, which are ordered by a certain refinement relation.
-Such refinement relation aims to model the way knowledge of the agent evolves when new information, provided by reliable external sources, is acquired.
-The semantics for IPL is given by intuitionistic Kripke frames expanded by a family of valuations indexed by the worlds and where the refinement order corresponds to the accessibility relation of the frame.
-This interesting 3-valued non-deterministic semantics for IPL departs from a different perspective than ours.
-Besides the 3-valued Nmatrices proposed in [11, Chapter 5] and the one we give in Definition 4.10 are quite different, it is not immediate whether the semantics proposed by Solares-Rojas could be seen as given by an RNmatrix.
-This is the subject of future investigation.
-It is interesting to notice that, by considering finite and decidable RNmatrix semantics (as the ones for S4 and IPL analyzed here) as a generalization of the standard notion of finite-valued (deterministic) truth-tables, the results presented here for S4 and IPL are a way to overcome the limiting results obtained by Gödel and Dugundji for IPL and S4, respectively, mentioned in Section 1.
-As a topic of future research we intend to explore the meaning of the partial valuations corresponding to the rows of the non-deterministic truth-tables in IPL and S4.
-In particular, we plan to investigate their possible connection with the mini canonical models considered by Hughes and Cresswell in [5, Chapter 8] to prove the finite model property for some modal systems, including S4.
-Finally, the extension of this methodology to other logics such as S5 and Prawitz’s ecumenical systems (see [10]) will also be investigated.
-Acknowledgments: The first and second author thanks the support of the São Paulo Research Foundation (FAPESP, Brazil) through the PhD scholarship grant #21/01025-3 and the Thematic Project RatioLog #20/16353-3.
+It follows from Theorems 8.12, 8.18 and 9.22. ∎
+The latter result shows that the procedure of constructing (non-deterministic) truth-tables in Definition 9.9 determines a decision method for propositional intuitionistic logic $\mathbf{IPL}$.
+10 Some considerations on complexity
+Unlike deterministic truth-tables, the size of non-deterministic truth-tables cannot be determined solely by the number of atoms.
+This is because, in principle, each subformula may (or may not) create new rows.
+Thus, we must also consider the number and shape of every evaluated subformula.
+A workaround for this problem is to establish size limits using boundaries.
+More formally, we can define a general upper bound for non-deterministic matrices by identifying their branching factor as follows.
+Definition 10.1 (Branching factor).
+Let $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ be some Nmatrix over a signature $\Theta$.
+The branching factor of $\mathcal{M}$ is the lowest $\kappa$ such that, for every n-ary $\in\Theta_{n}$ and every $a_{1},\ldots,a_{n}\in\mathcal{V}$, $|\mathcal{O}(\#)(a_{1},\ldots,a_{n})|\leq\kappa$.
+Definition 10.2 (Upper bound).
+Let $sub(\alpha)$ be the set of all subformulas of $\alpha$, $atoms(\alpha)$ the set of subformulas with complexity equal to 0, $\kappa\geq 1$ the branching factor, and $\theta\leq|\mathcal{V}|$ the number of truth-values used for atomic valuations.
+Then, the maximum number of rows of the initial matrix for $\alpha$ in $\mathcal{R}(\mathcal{M})$ is given by
+$ub_{\mathcal{M}}(\alpha)=\frac{\kappa^{|sub(\alpha)|}}{\kappa^{|atoms(\alpha)|}}\times\theta^{|atoms(\alpha)|}$
+Example 10.3.
+In $\mathcal{R}(\mathcal{M}_{IPL})$, $\kappa=2$ and $\theta=2$ . Hence,
+$ub_{\mathcal{M}_{IPL}}=2^{|sub(\alpha)|}$
+A fortiori, we can infer that the upper bound for the Nmatrix is also an upper bound for the restricted Nmatrix.
+Therefore, we can expect no more than $2^{|sub(\alpha)|}$ rows in a truth-table created for $\alpha$ with $\mathcal{R}(\mathcal{M}_{IPL})$.
+The truth-table may be shorter than $2^{|sub(\alpha)|}$, but not less than $2^{|atoms(\alpha)|}$, given that there are at least $2^{|atoms(\alpha)|}$ (trivially) valid partial valuations in the truth-table.
+This exponential growth in the number of subformulas may be problematic; a formula with only one variable can be arbitrarily large and produce immense tables.
+However, truth-tables are meaningful.
+By constructing a truth-table, one generates significant information about a given proposition or argument.
+A future work direction would be improving the average-case runtime by refining the algorithm with state-of-the-art strategies.
+11 Related works
+The idea of using finite non-deterministic semantics for characterizing intuitionistic logic and modal logic is not new.
+This section will mention some approaches to this topic proposed in the literature.
+We do not intend to present a detailed description and comparison between all these approaches here since they are beyond the scope of the present paper.
+(1) Andrea Loparic already obtained in 1977, announced in [23] and finally published in [24], a semantical characterization of Johannson-Kolmogorov minimal logic and intuitionistic propositional logic $\mathbf{IPL}$ in terms of a $2$-valued, non-truth-functional valuation semantics which induces a decision procedure for these logics.
+Similar to Grätz’s method (and to the adaptation to $\mathbf{IPL}$ proposed here), Loparic’s approach is based on a criterion for distinguishing between what she called semivaluations and valuations, the latter being the useful ones.
+This would correspond to level valuations, since such valuations are associated with $\varphi$-saturated sets, and they are the ones that characterize the logic.
+The associated algorithm generates a non-deterministic truth-table for a given formula, as in the cases of S4 and $\mathbf{IPL}$ presented in this paper.
+Then, using formal criteria based on implicative formulas, it deletes certain rows.
+Thus, a formula is a tautology of the given logic if and only if it receives the value 1 in all remaining rows.
+(2) J.
+Kearns proposed in [18] a semantics for (first-order) intuitionistic logic based on the notion of justifications.
+A 3-valued Nmatrix, whose domain is not formed by truth-values but by values of justification (++ for ‘justified’, $o$ for ‘weakly unjustified’, and - for ‘strongly unjustifiable’), is considered for interpreting the connectives of $\mathbf{IPL}$.
+A sentence is weakly unjustified if it is not justified by the currently justified sentences but might become justified later.
+In turn, strongly unjustifiable sentences are those already ruled out by the currently justified sentences.
+Given two valuations $v$ and $v^{\prime}$ over the Nmatrix, $v\preceq v^{\prime}$ ($v$ is contained in $v^{\prime}$) if $v^{\prime}$ preserves the values ++ (and so, the values -) assigned by $v$ to each formula.
+Hence, if $v^{\prime}(\alpha)=o$ then $v(\alpha)=o$, for any $\alpha$.
+The intuitive idea is that $v^{\prime}$ is a ‘possible future’ of $v$ (hence $v^{\prime}$ has more ‘knowledge’ than $v$, provided that the weak epistemic status $o$ may be corrected to the stronger ++ or - in the future).
+Kearns constructed a hierarchy of level valuations (as he would later do for modal logics) starting with level 0, which is formed by all valuations on the Nmatrix.
+The $(n+1)$th-level valuations are the $n$th-level valuations satisfying certain restrictions defined in terms of the sets $X_{n}^{v}=\{v^{\prime}\mid v^{\prime}$ is an $n$th-level valuation such that $v\preceq v^{\prime}\}$, for any $n$th-level valuation $v$.
+A basic feature is that, if $v$ is an $(n+1)$th-level valuation and $v^{\prime}(\alpha)\neq+$ for every $v^{\prime}\in X_{n}^{v}$ then $v(\alpha)=$-.
+That is, if $\alpha$ is not justified in any possible future of $v$, then $\alpha$ must be strongly unjustifiable at present $v$.
+In a second paper [20] on the subject, published two years after [18] and seven months after [19], Kearns improved the interpretation of the justification semantics and simplified its formal presentation by using trees instead of levels.
+The contained relation $\preceq$ is crucial to define the trees whose nodes are valuations.
+It should be stressed that Kearns’s proposal constitutes a 3-valued RNmatrix semantics for $\mathbf{IPL}$, although it differs significantly from our approach.
+In particular, Kearns’s RNmatrix does not seem to induce a decision procedure for $\mathbf{IPL}$.
+(3) Concerning modal logics, in [21], right after the publication of [12], O.
+Lahav and Y.
+Zohar proposed a decision procedure for systems K and KT based on 4- and 3-valued RNmatrices, respectively, by using a decision procedure different (but closely related) to Grätz’s approach.
+Indeed, for ${\bf L}\in\{{\bf K},{\bf KT}\}$ they considered a sequent calculus $\mathcal{G}_{\bf L}$ and sets $\mathbb{V}_{\bf L}^{\mathcal{F},n}$ of $n$th-level valuations relative to a set $\mathcal{F}$ of formulas (the domain of such valuations) and $n$ is the maximum number of applications of the necessitation rule validated by such valuations.
+Given a sequent $\Gamma\Rightarrow\varphi$, the expression $\vdash_{\mathcal{G}_{\bf L}}^{\mathcal{F},n}\Gamma\Rightarrow\varphi$ denotes that there exists a derivation in $\mathcal{G}_{\bf L}$ of $\Gamma\Rightarrow\varphi$ with a derivation containing exclusively formulas in $\mathcal{F}$ and such that the number of applications of the inference rule of necessitation $\Delta\Rightarrow\psi/\Box\Delta\Rightarrow\Box\psi$ in any of the branches of the derivation is limited by $n$.
+They proved that $\vdash_{\mathcal{G}_{\bf L}}^{\mathcal{F},n}\Gamma\Rightarrow\varphi$ if and only if $\varphi$ is a consequence of $\Gamma$ w.r.t.
+$n$th-level valuations in $\mathbb{V}_{\bf L}^{\mathcal{F},n}$.
+Moreover, they obtained a decision procedure for K by considering, for each sequent $\Gamma\Rightarrow\varphi$ to be tested, the set of level valuations $\mathbb{V}_{\bf L}^{\mathcal{F},n}$, where $\mathcal{F}$ is the set of subformulas of $\Gamma\cup\{\varphi\}$ and $n=4^{|\mathcal{F}|}$.
+For KT, it is enough considering $n=3^{|\mathcal{F}|}$, since the RNmatrix for KT is 3-valued.
+Moreover, that RNmatrix coincides with the one proposed in [12] for KT.
+It is worth noting that [21] considers the reduced version of the underlying Nmatrix, expanded to the full signature.
+This produces the Nmatrix $\mathcal{M}^{\prime}_{S4}$ for S4 of Table 2, but where $\Box^{\prime}(2)=\{1,2\}$.
+Despite these similarities, the decision procedure in [21] (as mentioned earlier) differs from that in [12].
+It consists of generating the set of $n$th-level valuations over $\mathcal{F}$ for appropriate $n$ and $\mathcal{F}$.
+(4) More recently, A.
+Solares-Rojas proposed a 3-valued Nmatrix semantics for $\mathbf{IPL}$ based on intuitionistic Kripke models and the novel notion of depth-bounded approximations (see [29, Chapter 5]).
+Within this semantic approach to $\mathbf{IPL}$, a 3-valued Nmatrix — whose truth-values are intuitionistic truth (1), falsity (0), and indeterminacy ($u$) — gives the meaning of the connectives in terms of information that a given agent holds.
+The valuations over the Nmatrix aim to model actual information states, ordered by a certain refinement relation $\sqsubseteq_{3}$.
+This relation models how the agent’s knowledge evolves when new information from reliable external sources is acquired.
+The semantics for $\mathbf{IPL}$ is defined via intuitionistic Kripke frames expanded with a family of valuations indexed by the worlds, where the refinement order corresponds to the frame’s accessibility relation.
+Some similarities between this approach and Kearns’s (mentioned in point (2)) may be observed.
+Conceptually, both approaches aim to model the evolution of states of knowledge: the information state of an agent in [29], and the knowledge of an idealized community in [18, 20].
+Regarding technical aspects, by identifying $1$, 0, and $u$ with ++, -, and $o$ respectively, the Nmatrix in [29, Chapter 5] coincides with that in [18], with one exception: while $o\vee o=\{o,+\}$ in Kearns’s work, $u\vee u=\{u\}$ (i.e., $o\vee o=\{o\}$) in Solares-Rojas’s approach.
+Additionally, the refinement relation $\sqsubseteq_{3}$ between valuations coincides with the contained relation $\preceq$ introduced in [18].
+The relationship between both approaches (in particular, the connections with the improved version [20], which is closer to Kripke and Beth’s semantics) deserves to be investigated.
+Regarding the connections with our approach, although the 3-valued Nmatrices proposed in [29, Chapter 5] and the one we gave in Definition 8.4 are quite different, it is not immediate whether the semantics proposed by Solares-Rojas could be seen as given by an RNmatrix.
+This is the subject of future investigation, as well as the relationship between our approach and the semantics of justification introduced by Kearns.
+12 Concluding remarks
+In this paper, we proposed a new decision procedure for intuitionistic propositional logic ($\mathbf{IPL}$) based on a 3-valued restricted Nmatrix semantics.
+This procedure was obtained by abstracting the composition of Grätz’s decision procedure for S4 [12] and Gödel’s faithful translation from $\mathbf{IPL}$ to S4.
+A 3-valued Nmatrix with an intuitive interpretation in epistemic terms was proposed for $\mathbf{IPL}$.
+By adapting Grätz’s algorithm for S4, a suitable notion of level valuations was obtained, which allowed us to define a straightforward algorithm for deleting spurious rows in the non-deterministic truth-tables generated by the Nmatrix for any formula.
+The soundness and completeness proofs for the RNmatrix and the resulting truth-table algorithm for $\mathbf{IPL}$ were presented in full technical detail.
+It is worth noting that the Nmatrix for $\mathbf{IPL}$ introduced in Definition 8.4 has a clear and intuitive interpretation in terms of knowledge (see comments after Definition 8.4) and constructibility.
+Indeed, recall that the valuation associated with a $\varphi$-saturated set $\Delta$ in $\mathbf{IPL}$ (given in Definition 8.14) acts as a 3-valued characteristic function of $\Delta$, and note the observations in Remark 8.15.
+By considering finite and decidable RNmatrix semantics as a generalization of standard finite-valued deterministic truth-tables, the results presented here provide a way to overcome the limitations obtained by Gödel and Dugundji for $\mathbf{IPL}$ and S4, respectively, as mentioned in Section 1.
+As a topic of future research, we intend to explore the meaning of the partial valuations corresponding to the rows of the non-deterministic truth-tables in $\mathbf{IPL}$ and S4.
+In particular, we are studying the possibility of generating finite Kripke models for $\mathbf{IPL}$, KT and S4 from the obtained truth-tables.
+This could be connected with finite approximability ([5, p.
+49 and Ch.
+11]) or with mini-canonical models ([13, Ch.
+8]).
+Finally, the extension of this methodology to other modal logics, as well as Prawitz’s ecumenical systems (see [27]), will also be investigated.
+Some steps in this direction were taken in [22].
+Acknowledgments: The first and second authors thank the support of the São Paulo Research Foundation (FAPESP, Brazil) through the PhD scholarship grant #21/01025-3 and the Thematic Project RatioLog #20/16353-3.
 The second author also acknowledges support from the National Council for Scientific and Technological Development (CNPq, Brazil), through the individual research grant #309830/2023-0.
+The third author thanks Fundação Carlos Chagas de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ, Brazil), through the research grant #E-26.210.971/2019.
 Appendix A
-The following truth-tables show that each axiom of the Hilbert calculus $\mathcal{H}_{IPL}$ for IPL is valid in the intuitionistic truth-tables.
-These tables were automatically generated by a Coq module for RNmatrices written by R.
+The following truth-tables show that each axiom of the usual Hilbert calculus for IPL is valid in the intuitionistic truth-tables.
+These tables were automatically generated by a Rocq module for RNmatrices written by R.
 Leme, see
@@
 T T T T T
-$p$ $q$ $p\lor q$ $p\to p\lor q$ $q\to p\lor q$
+$p$ $q$ $\lor q pp$$\to p\lor q$ $q\to p\lor q$
 F F F T T